A Complete Bibliography of The Journal of the American Statistical Association: 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.05

Title word cross-reference

χ^{2} [Ber49, Cro45]. t [Cro45, Wal49c, Wal49d].

-Test [Wal49c, Wal49d].

101st [Ano40w]. 102nd [Ano41f]. 103rd [Ano42b]. 104th [Ano45p]. 105th [Ano46f]. 106th [Ano47g]. 107th [Ano48b]. 108th [Ano49f]. 1929-38 [Ynt40b]. 1932 [Fri40b]. 1935-36 [Tuc42]. 1937 [Boe40]. 1938 [Phe40, Wor40a]. 1939 [Ano40w, Nor40, Mus40]. 1940 [Cro41a, Dun41d, Hay41b, Win44]. 1941 [Ano42b, Cra42, Phe41, Ros41, Wil41b]. 1942 [Ano43c, Ano43h, Arm42, Dur43, Hin43, RT43, Lum44]. 1943 [Dei43, Lot44, Sim43, Tho43, Wei44]. 1944 [Mol44]. 1945 [Mar46]. 1946 [LeC46, Mye46, Wil46a]. 1947 [Ano48b, Tho47]. 1948 [Kuz49].

25-Cent [Hin40]. 29 [AG46].

36 [Vai40, Vai40, Gil40a]. 39 [Den44a, Tip40].

44 [Mig44].

A. [Dur43]. Ability [Tuk47b]. Abnormal [Jam40]. Abraham [BF47a]. Accumulating [San41b]. Acheson [Dem45a]. Across [Boa41]. Act [Boe40]. Action [Wal45a]. Activities [Ano40e, Ano40f, Ano40g, vS40]. Activity [Sie44]. Acts [Vic41, WCH40, Wal41, WCH41]. Actual [Mye43]. Actuarial [AG46, Dod46, Hoh40, Mye47, Nie48]. Adequacy [Pal41a, Ree41b, Som41]. Adequate [Bur47f]. Adjacent [Lat42]. Adjusted [Whe46]. Adjusting [Pat46]. Adjustments [Lub47a]. Administration [Ric43]. Adults [Kle44]. Advancing [She46]. Aerodynamics [Smi43]. Affected [Gil42]. Affiliate [Edw42]. Age [LS40, Mye49b, Whe46, Nug40, Eps40]. Agencies [Cro44]. Aggregate [Vai40]. Aggregates [War45]. Aging [Pea40a]. Agricultural [Bat45a, Ebl44, Pet47, Rob44, SK42, Sne42, TT44, Wai42, Wor40a]. Agriculture [Coc45, DB45, Hol43, Jes45, Kin45, Vai40, Jes49, Tae47]. Aid [Cor47b, Som41]. Aids [HD43]. Aircraft [AG46]. Alex [Den44a]. Alexander [Wil49]. Alien [HR47]. Allen [Olm42, Ynt40c]. Allied [Vin40]. Alphabetical [Ano45a, Ano45b]. Alsberg [Nor40]. Aluminum [Gol46]. America [Fri40b]. American [Pro40a, Ano40d, Ano45c, Dav40, Fun41b, HM46, Ric44, Ste40b, Wal45c, Wil40b]. Amount [Sza46]. Amounts [Ynt40b]. Amplitudes [Jul41]. analyse [Sas40a]. Analyses [And48, Den44a]. Analysis [AG46, Alt40b, And47, Asc44, Bra41a, Bro42a, Coc43, Haf40, HM46, Kat47, Lur46, Mor40b, Mos41, Mow41, Nie45, OC49, Pat46, Pie44a, Pie44b, Rao47, Sch47a, Sch47b, Sch44a, Smi47, Tin40a, Tin46, Tuk47b, WM41, Wau43, Wil46b, Woo44b, Ynt40c, BF47a, Wil40a]. Analyzing [BO46, DM48, Kap42, Old49, Thu48]. Anderson [War45]. Angular [KC47]. Anniversary [Ano40x, Fun41b]. Annual [Ano40w, Ano40y, Ano41f, Ano42b, Ano45p, Ano46f, Ano47g, Ano48b, Ano49f, Hau42, McD40]. Anthony [Mar46]. AOQL [GB49]. Appendix [Ano44a]. Application [Ber44, CO49, Cow46, Eps48, Gir42a, Jul41, LA46, Mar45, Sim42, Wad45]. Applications [Dem45a, Tin46, Wal49b, WNP43, Woo42b, You46, Dem45a]. Applied [Fri40a, Kor40, McV47, Ric43]. Apportionment [Cor47b, Ull45]. Appraisal [Ano44a, DW41, MBC^{+}43, Per42]. Appraising [Pal41a]. Appreciation [Bur47a, Kuz49]. Approach [McC43, RB47]. Approaches [Bro41b]. Approximation [Ber46]. Area [Che49, Las48, McV47]. Areas [BO46, Bru44, Osb42, Ynt40b]. Arising [Gei49]. Army [McP45]. Arne [Mol44]. Arrays [Bak41, Eis46]. ASA [Ano45n]. Asia [Tae45b]. Aspects [Lan41, Nor40, Tae47]. Assay [Ber44, Ber49, Bli40]. Assayed [Bli44b]. Assays [KC47]. Assigned [Sch48b]. Assistance [Ynt40b]. Associate [Ano45d]. Association [Ano43m, Ano44l, Dav40, Ste40b, Wil40b, Ano40d, Ano45c, Coa40, Fun41b, HM46, Lub47b, Wal45c]. Associations [Con47, Vin40]. Assumed [Mye43]. Atlantic [Ano47g]. Attempt [PS49]. Attrition [Gra45a]. Audiences [Pol43]. Austin [Ric49]. Auto [CO49, OC49]. Auto- [CO49]. Auto-Regressive [OC49]. Availability [Dun41c]. Avalanche [Sau40]. Average [Lat42, Spu40]. Ayres [Bur47a, Wer40].

B [Phe41, AG46]. B-29 [AG46]. Back [Ano40a, Ano40b, Ano41a, Ano42a, Ano43a, Ano44b, Ano45e, Ano45f, Ano47a]. Balances [Len44]. Bank [GL47a, Joh41a, War48]. Banking [Sel40]. Barger [Fri44a]. Bartels [Haf40]. Based [Coc43, GW47, MW43b, Mye49b, San42a, San44b, San42b, Wil44b, Woo42b]. Basic [Cro41b, Mon40, Sne48]. Basis [Bec44b, HB45, Sim45, Ull45]. be [Bra40b, Wil40c]. Beck [Kle44]. Bellevue [Ano40w]. Beneficiaries [Aar44, Win44]. Beneficiary [Mye49b]. Benefits [Aar44]. Bennett [Nor40]. Berkson [Cam40, Fis43]. Bernoville [Sas40a]. Best [GW47, Wal49c]. Beta [Ban45]. Between [Bak42, Bra41a, Jul41, Sie41a, Sie41b, Sza45, Wal45a]. Bias [CF47]. Biases [Cor42]. Biltmore [Ano42b]. Binder [Nat40]. Binomial [Ber40b, MT49]. Bio [Ber44, Ber49]. Bio-Assay [Ber44, Ber49]. Biological [Bli40, KC47]. Biologically [Bli44b]. Biology [Coc45]. Biometrics [Ano45d, Ano45n]. Birth [GT41, LS40, SD49]. Blakey [Den44a]. Block [Old49, Rao47, THD43]. Blocks [Ree42]. Board [Gor44, NS48a, NS48b, NS48c, NS49]. Bond [Tos45]. Book [Abb40, Abb41, Abb42, Ack41, Ald41, Alt42, Alt44, Alt40a, And45, AS46, AM47, And49, And43, Ark44, Arn45, Arn46, Aro48, Aya43, Ayr42, Bak47, Bak49, Bal45, Ban45, Ban49a, Ban49b, Bar42, BL47, BF47a, Bar41, Bat45a, Bea44, Bel41, Ben42a, Ben42b, Ben43, Ber40a, Ber42a, Bla45a, Bla45b, Bla47, Bli41, Blo43, BV46, Bod41, Boe40, Bow49, Bra48, Bra41b, Bra41c, Bra43, Bra44, Bra46, Bre42, Bro42b, BT46, Bru43, Bue43a, Bue43b, Bur40a, Bur43, BN48, Bur40b, But43, Byr40, Cah49, Cam45, CM49a, Car41, Cas44, Cha42, Cha49, Cha41, Che40a, Chu46, CJ47, Chu49, CMB47, Chu47, Cio41, Cla43, Cli47, CD47a, Coa42, Coc42a, Coc44, CP48, Col45, Con45, Cor49, Cot47, Cov40, Cow43a, Cow44, Cow45, Cow41, Cox46, Cra45, Cra48]. Book [Cra43, Cra44, Cre43a, Cre43b, Cro42, Cru47, Cru49, Cur40a, Cur40b, Cur45, Cur46, Cur47, CD47b, CM49b, Cus43, Dal42, DD47, Dav44a, Dav44b, DWL48, DH49, Dea45, Dee46, Dem44, Dem45b, Dem45c, Dem45a, Dem48, Dem49, Den44a, Den44b, Dew42, Dod45, Dor40a, Dor40b, Dor40c, Dor42, Dou40, Dou41a, Dou41b, Dou41c, DuB44, Dud48, Dun42, Dun41a, Dun41b, Dur41, Dwy45a, Dwy48, Edg40, Eis41, Eis46, Eis47, Ell40, Ely46, End41, End42a, End42b, End43, Eng47a, Eps40, EM43, Eps47, Eps49, Esk44, Eud42, Eva49, Eze42, Eze43, Fab40, Fab41a, Fab41b, Fai42, Fel47, Fer47, FJ47, Fer45, FN47, FH49a, FS48a, Fie43, FY46, Fis46, Fis41, Fis42, Flo41, For41, For44, Fos44b, Fos44a, FM48, Fra41, Fre45, Fre47b, Fre47a, FG48, Fre49, FS48b]. Book [Fri40a, Fri40b, Fri41, Fri42, Fri43, Fri44a, Fri44b, Gag41, Gau46, Gee44, Gil41a, Gil43, Gil40a, Gil41b, Gir40, Gir42b, GK46, Gol41, Goo45, Gou44, Gou41, Gou46, Gra41a, Gra41b, Gra45b, Gra46, Gra44, Gra42, Gre41, Gre46, Gre47, Gre48a, Gre49, Gum47, Gut45, GW46, Gut47, GL47b, Hab40, Hae48, Hag43a, Hag41, Hag43b, Hag49, Han40, HL46, Har41, Har42a, Har42b, Har45, Har48, Har42c, Has43, Haw45, HW41, Hed41, Hed45, HW47, Hoh41, Hoo40, Hoo48, Hor49, Hot45, Hov43, Hoy41, Hur46, Hur47, Hut42, Hyd41, Ise42, Jam40, Jam41, Jas41, Jas42, Jen40, Joh40, Joh41b, Joh45, Joy42, JW48, Kac48, Kei45, Kel45a, Kem48, Ken49, Kin40, Kin42, Kir43, Kis43, Kle44, Kno46, Knu49, Kol41, Koo41, Koo47, Kre40, Kuz47a, Kuz47b, Lai45]. Book [Laz44, Lea44, LW47, Leo42, LS48, Lev49, Lew40, Lew45, Li46, Lit46, LT47, LS49, Lit49, Liv45a, Liv47, Loc42a, Loc42b, Lon42, Lor42, Lor43, Lou45, LS47, Luk48, Lum44, Lun41, Lun40, Mac42, Mag42, Mal41, Man45, Man40, Man42, Mar49, May40, May44, May42, McD40, McD41, McD42, McG44a, McI41a, McI42, MW43a, MS43, Men41, Mer43, Mer44, MB48, Mey49, Mig44, Mil43, Mil45, Min41, Mod48, Mol43, Mon40, MS46, Mor40b, MT46, Mos45a, Mos45b, Mos47, Mos49b, Mos49c, Mot49, Mud46, Mue48, Mur45, Mus43, Mus40, Mye44, Mye45a, Mye49a, NS45, Nag40, Nat40, Nei41, Ner40, Nes47, New42, Ney40a, Ney40b, Ney43, Nic43, Nic44, NP48, Nor40, Nor45a, Nug40, Ogl43, Old46, OO48, Olm42, Olm45, Øre48, Ott42]. Book [Pal41b, Pal42, Pal43b, Pal44a, Pal44b, Par42, Par40, Par43, Pea47, Pea49, Pec42, Pet45, Phe40, Phe42, Pit49, Pri43, Pro40a, Pro40b, Pro43, Pyl40, Ree41a, Ree44a, Ree44b, Ren43, Rey41, Rey42a, Rey42b, Rey44, Rid44, Rid47, Rig41, Rig44, Rob49, Rob40, Rob41b, Rob41c, Rom45, Rom49, RW48, Ros46, Sal45, Sam43, San41a, Sas40a, Sas40b, Sat41a, Sat41b, Sau40, Sca41, Sca42, SW46a, Sch46, Sch48a, Sch41a, Sch41b, Sch43, Sch44b, Sch42a, Sch40, Sch42b, Seb49, Sha44, Shi45, Sho43, Sib42, Sie46, Sma40, Smi40, Smi41a, Smi42, Smi45a, Smi46, Smi45b, Sna43, Sne47, Sni40, Sol49, Spe41, Spe44, Spu41, Spu46, SvS48, Spu49, Ste45, Ste41, Sti48, Str46, Swe41, Swe44, Tae44, Taf44, Tho46, Tho45, Tim43, Tin40b, Tin47, Tin48]. Book [Tip40, Tor45, Tou40, Tre41a, Tre41b, Tre41c, Tru44, TW46a, Vai40, Vai41, Vee40, Vil41, Wai47, Wal43a, Wal47a, Wal40, Wal43b, Wal44, Wal45b, Wal47b, Wal42, War45, War43, Wat41b, Wat41a, Wat40, Wel44, Wel41, Wer40, Wes42, Whe47, Wil42a, Wil40a, Wil41a, Wil41c, Wil43, Wil44a, Wil41d, Win44, Wis46, Wol40, Wol43, Wol44, Wol46, Wol49, Woo42a, Woo42b, Wor40a, Wor46a, Wor41, Wor44b, Woy41, Wyn44a, Wyn44b, Yen43, Yer48, Ynt40c, Ynt40a, Ynt42, Yod42, Yod43, Yod44, You43, Zim40, Zim43, Zub49, Gei48]. Books [Phe40, Ros46]. Borrowed [Rob41a]. Boston [Cog40]. Brazilian [Fal40]. Bressler [Vai40]. Brinton [Nag40]. Britain [Cru46, Mos49a]. British [Lan44]. Brownlee [TW46b]. Building [Tou40]. Bureau [Ano44a, MBC^{+}43, Hau47, SL49, Tep43]. Buren [Lot44]. Burgess [Byr40]. Burned [Pyl40]. Burned-Clay [Pyl40]. Business [Ano46a, Ano49f, Boa41, Bro41b, Bur40a, Bur47f, Che40b, Don46, Fri40b, GT41, HHG46, Mee45, Per42, Ree42, Ric43, Tuc45, War48, War45, Wer40, Swe44, Mor40b]. Business-Cycle [Fri40b]. By-Laws [Ano40c, Ano40d]. By-Product [Mor49].

C [Cra45, Cud40, NS45, Nag40, Nor40, Tuk49b, Wil40a]. Calculating [CD41, Dwy40, Sza46]. Calculation [Bli44b, CD41, Dwy40, Nor45b]. Calhoun [Mye46]. Callbacks [Ano49e, PS49]. Can [Bra40b]. Canadian [Coa40, DD43]. Capacity [Liv45c, Mee42, Sie42b]. Capita [Koz41]. Capital [Jam40, McD48, Rob41a, Tuc43b]. Capitalism [Kin40]. Capitalist [Mor40b]. Capt [Ric49]. Card [Ber41, BO46, DTG42, Knu42]. Cards [HM42, Les49, McP42, San41b]. Care [Kle44]. Careers [Mer44]. Carl [Kin40, Nor40, Bur46a]. Carlo [MU49]. Carroll [Arm42]. Case [Noe49]. Cases [Dor40a]. Casualty [Kor40]. Categories [Sch48b]. Causes [Hay41a, Per40]. Celebration [Ano40y]. Census [Bru41, Bru44, Che40b, Che49, DG41, Dun40a, Dun41c, Eck45b, HD43, Hau41, Hau42, Hol43, Ree42, SDH40, Tru41, Hau47, SL49]. Censuses [DS41b, Jes49]. Cent [Hen47, Hin40]. Centenary [Ano40x, Ano40y, Ste40b, Ste40a]. Central [Bro41a]. Centralized [Cro44]. Cents [Bak45]. Certain [Cor42, Cow47, Nie45]. Chaddock [Cro41a]. Chalfonte [Ano47g]. Change [Sza45]. Changes [Asc44, Mye45b]. Changing [Ynt40b]. Chapter [Ano40e, Ano40f, Ano40g, vS40, Cog40]. Charles [Boe40, Lum44, Sau40, Wil41b, Hin43]. Chart [Bli44a, Cro45, GB49, Mar45, Sch44a, Sch44c]. Charter [Ano40h, Ano40i]. Charts [KB43, Las48, Liv45b, Sch47a, Sch47b]. Cherington [Tho43]. Chi [Ber40b, Ber46, Bli44a, Nor45b, Sch47a, Sch47b]. Chi-Square [Ber40b, Ber46, Bli44a, Nor45b, Sch47a, Sch47b]. Chicago [Ano41f]. Choice [Hoe47, Wal49c, Wau43]. Choices [Sim45]. Chretien [Cra45]. Churchman [Tuk49b]. Chute [Pyl40]. Circular [Ano40j]. Citation [Ano47b]. Cities [TW43, Pro40a]. City [Ano47g, Ano48b, Old49]. Cityward [Abb40]. Clair [Kuz49]. Clarence [Fri40a]. Clark [Dur46]. Classes [Sog44, Tuc43b]. Classification [Dem42, FW40, Fra40, Kol40, Kol44, Sal42]. Classified [Cav43]. Clay [Pyl40]. Clerical [Cav43]. Cleveland [Ano46f, Ano49f]. Clinical [Kle44]. Clyde [Ric49]. Coding [Ber41]. Coefficient [Gir42a, Lew41b]. Coefficients [CD41, Dwy40, Knu42, Pla43, Pla44, Wau46]. Collection [Cro44, Eck45a]. College [Gil44]. Columbus [Win44]. Combined [All44, Bat44, Gor44, Rai42]. Comes [Nug40]. Comment [Sal42, Vic41]. Comments [BG41, Cam40]. Commercial [McV47, SH48]. Commission [Eng47b]. Committee [Ano43b, Ano44c, Fun41a, Hoy40, Mor40a, Mye43, Ste40c]. Commodity [Asc44, Kol44, Wor40b]. Commodore [Ano48b]. Common [Kap45, Sie43, Sza45]. Community [Fra42, Lum44]. Companies [Dir44]. Company [McD48]. Comparing [Bat45b, BO46]. Comparison [Mye43]. Compass [Las48]. Compensation [Win44]. Compiled [GL47a]. Complex [Nor45b]. Component [HB45]. Components [And47, And48]. Composition [Tuc43b, Woo44b]. Compulsory [Dod46, Mye47]. Computation [Wau46]. Computational [San42b]. Concept [Fer46, Sie44].

Concepts [Cow43b, Cro41b]. Concerning [Bat45b, BG41, Dun41c, Kap45, Sch44c]. Concluding [San44a]. Condition [San42a, San44b, Wil44b]. Conditions [Wal49b, Kle44]. Conference [McD40, Ynt40a]. Confidence [Noe49]. Congested [Bru44]. Congress [Ull45]. Conrad [Wil49]. Considerations [Ano49e, Pal41a]. Considered [Ber42b, Wil40c]. Consistency [Ban40]. Constitution [Ano40k]. Consume [Haa47]. Consumer [Gil40a, Goo47, Cov40, Gil40a]. Consumption [Gar46, Woo44a]. Contain [Ber41]. Containing [CO49]. Contents [Gag43]. Contingency [Bow48, Nor45b]. Contract [Par45]. Contribution [Kat47]. Control [Bro41b, Che49, Cro44, Lan41, Mar45, OK49, Ric43, Sch47a, Sch47b, Sch44a, Sch44c, Sim41, Wor45, Man40]. Controlled [Swe44]. Controlling [CF47]. Conversion [Dir44]. Coordinating [Dav43, Zei47]. Corporate [Ano40l, Rob41a]. Corporation [Eng47b, Her41]. Correction [And48, Ano43c, Ano46a, Ano47b, DB43, Pie43, Sch47a, Suk47a, Wal41]. Corrections [Gru48a]. Correlated [CO49]. Correlation [Bro42a, CD41, Cow43b, Dwy40, Dwy42, Dwy45b, Mow41, Pla43, Pla44, Wau46]. Correlations [Bak42]. Corrington [Mye46]. Cost [Ano44a, Ber47, Ber40c, BM44, EW41, MBC^{+}43, Mos41, Say41, Sie40, Tep43, TW43, Ulm46, WRS42]. Cost-Utility [Ber47]. Costs [Wil40c, Kle44]. Cottrell [Byr40]. Course [Sne48]. Courses [Wal45a]. Covariance [Bli40]. Cover [Osb42]. Cover-type [Osb42]. Creator [Kin40]. Credit [Cov40, Sni40, Jen40]. Criminal [Mer44]. Critical [Per42]. Criticism [Fis43]. Critique [Mye47]. Croner [Lun40]. Cube [Lan42]. Cubes [San41b]. Cudmore [Mar46]. Culture [Cra45]. Curran [May44]. Currency [Sha43]. Current [Bur45a, Bur48b, Bur48c, Eck45b, Par48, SL49]. Curve [Smi47, SA48]. Curves [Cow47, Koz41, Ras49]. Cut [San42b, Spu40, SA48]. CWA [Ynt40b]. Cycle [Fri40b]. Cycles [Ano43n, Fri40b, GT41, Mor40b, Wer40].

D [Fos44a, Sau40, War45, Ynt40c, Bli40]. D.C. [Ano45p]. Data [Ber41, Cro44, Daw45, DM48, DH46, Dun41c, Dwy40, ES43, EH44, Fra42, Hol43, Jon43, Kat47, Mor49, Mos48, Mye49b, Old49, Pat46, Ree42, Sch44a, Som41, Sza46, Thu48, Tin46, Wor44a, You46, Abb40]. David [Fri40a, Tuk49a]. Davis [RT43, Wor40a]. Dayton [Dor40a]. Dealer [DS46]. Death [Ano45o, SD49]. December [Ano40w, Ano42b, Ano48b]. Decentralization [Smi47]. Deciding [Wal45a]. Deep [THD43]. Defense [Den41, Mee42, Mye42, Tai41]. Deflation [Shi40]. Delano [Ano45o]. Delimitation [Bat42]. Demand [Cor45, Gir42a, Hen47, Koz41, Lew41a, Mos41]. Deming [Ban45, Man40, Sal42, Wor44a]. Democratic [Lub47a]. Demographic [Mye43, Mye49b, Tae45b]. Dental [Kle44]. Department [Lam47, Tae47]. Departure [BG41, Mad40]. Dependent [Wau43]. Deposits [Sha43]. Depression [FH49b]. Description [Ano45c]. Design [Bli40, Bra41a, DS46, Jes45, Wil46b]. Designed [Ban45, Ber41]. Designing [HHM48]. Designs [Bra40a, Cha49, Mad46, Rao47]. Destructive [Jac47].

Determination [HHM48, Las48, Pla43, Pla44]. Determined [Lat42]. Determining [Bor42, Knu43, Nor44]. Detroit [FP40]. Developing [Mil48]. Development [Gei49, Kin45]. Developments [Dwy42, Gre48b, Hin47, SK42, TT44]. Deviation [Coh49, GW47]. Deviations [Bak41, Pla43, Pla44]. Dewey [RT43]. Diabetes [Alt40b]. Difference [Che44b, Lat42, Sie41a, Sie41b]. Differences [Bak45, HM42, MW43b, Per40, Sie42a, Wil40c]. Differencing [LA46]. Difficulties [Bra40a]. Dimension [Mil46]. Direct [Las48]. Directory [Kel45c, Ste40a]. Discrete [Pie43]. Discriminant [Bat45b, DS42]. Discussion [Tuc45]. Disorders [Dor40a]. Dispersion [Hir43]. Displacement [Tae44]. Disproportionate [Pat46]. Distribution [All44, Bat44, Bli44a, Gor44, Hir43, Lew41b, Pie43, Rob44, Sch47c, Sha43, Tuc42]. Distributions [Coh49, Cra45, Cro45, DS41a, Rai42, Sch47c]. Do [Sol41]. Dodd [Sim43]. Donald [Kel45a]. Doolittle [Cow43b]. Dorothy [Kle44]. Doten [Arm42]. Douglas [Cra45]. Dr. [BG41, Dur46, Fis43]. Drafting [vH42]. Drugs [Bli44b]. Dugan [Den44a]. Duge [Sas40a]. Duncan [Dem45a]. Durum [Nor40]. Dynamics [Mye41a].

Earners [Cav43]. Earnings [GL47a, Leb48, Per40, Rob41a, Sie40, Ynt40b, Wel44]. Eastern [Tae45b]. Economic [Cav43, Cov40, Cru46, DH46, Kat47, Kin40, Koo45, Lub47a, Nic42, Par48, RB47, Tin40a, Tin46, Tip40, Ulm46, War45, Wat49, Woo44a]. Economics [Bur40a, McD40, Tou40]. Economists [Ynt40c]. Edge [Les49, Thu48]. Edge-Marking [Thu48]. Edition [Tuk47b]. Editorial [Kel45b]. Education [Fri40a, Kim44]. Edward [Sim43]. Edwards [Ban45, Man40, Wor44a]. Effect [ME42]. Effects [Fin48, Hin40, Vic41, WCH40, Wal41, WCH41]. Efficiencies [HH42, Hen44]. Efficiency [Ber47, Lav46, THD43]. Eighth [Tip40]. Elasticity [Gir42a, Lew41b, Mil46]. Eldon [Vai40]. Eleanor [Mer44]. Election [Min45]. Element [Cra45]. Elementary [Dem45a]. Elements [Kap42]. Eliminating [Ano49e]. Emergence [Bak45]. Emmet [Cro41a]. Empirical [RB47]. Employees [Leb48, Rob44]. Employment [Daw45, Eck41, Eck45b, Fri44a, Hen47, Hin40, JW45, Ree41b, SW46b, Tae45a, Tol45, Vai40, Wal45d, Wil44a]. Engineering [Sim42]. Engines [AG46]. England [Lum44]. Enterprise [Vai40, Bur40a]. Enumerations [Wad45]. Enumerative [Pal43a]. Equal [Bak41, Kir48]. Equations [Bre49]. Equilibrium [Smi45b]. Erkenntnismittel [Alt40a]. Ernest [Byr40]. Error [CO49, Eva42, Kap45]. Errors [AK49, Cod48, DTG42, Osb42]. Erscheinungen [Fur46, Gei48]. Estadistica [Tuk47a]. Estimate [Bro41a, GW47, Tos45]. Estimates [Daw45, Dod46, Eck45b, Gil40a, Mye47, Nie48, SL49, Woo42b]. Estimating [Bec44a, Coh49, Dav42, Gru48a, Gru48b, HR47, McV47, SD49]. Estimation [Koo45]. Eugene [Tae44]. Europe [Tae44]. European [Nor40]. Evaluation [Aro44, Bre49]. Events [Sol41]. Evidence [AK49, Ber42b, Kel43]. Examination [EH44]. Examinations [Lew44]. Excess [May44]. Excessive [Sel40]. Executive [Don46]. Executives [Boa41]. Expanding [Daw45]. Expansion [Wil44a]. Expected [Com43]. Expenditure [Wor43]. Expenditures [Cav43, Gil40a]. Expense [Mil48]. Experience [BW46, Ber43, Mye43, NS48a, NS48b, NS48c, NS49]. Experiences [Win44]. Experiment [Bra41a, DB45, Sne42]. Experimental [Bra40a, Tuk49b, Wor44a, Bat45a, Cha49]. Experimentation [TW46b]. Experiments [HHM48]. Exponential [Sto41]. Exports [Smi40]. Extent [Per40, SD49]. Extraction [Lan42]. Extreme [Eps48].

F [Ano47k, Dur47, Fri40a, Kle44, Tuk49a, Wal40]. Factorial [Bli40, Tuk47b]. Factors [Lat42, Pal43a, Pat46, San42a, San44b, Wil40c, Wil46b, Wil44b]. Factory [Lum44]. Facts [Dor40a, Ree44c]. Failure [Byr40]. Fair [Boe40, Bor42]. Falkner [Hay41b]. Families [Cav43]. Family [DB45, Pol43, Woo41a, Woo44b, Wor43]. Fanno [Jam40]. Farm [Hen47, McC43, Tae45a, Vai40]. Farmers [BM44, DW41, Tos45]. Farms [BM44, Hen44]. Farnsworth [Nor40]. Fatal [Ric48]. Feature [Sne49]. Features [Tru41]. Fecundity [Whe46]. Federal [Cro44, Eng47b, GL47a, Jes49, Joh41a, Leo45, RHTH45, Vic41, WCH40, Wal41, WCH41]. Fellows [Ano40m, Ano45g, Min45]. Fertility [FP40, Woo49]. Field [Bak45, Bev45, Eck45a, McP45, Som41]. Fields [Nic42]. Fifth [McD40]. Figures [Koz41, San42a, San44b, Wil44b]. Finance [Nug40]. Finances [Goo47]. Financial [Eps40, Gue40, Mus40, Nic42]. Financing [Dir44]. find [Vor41]. Finite [Cor47a, Cor44, Hir43]. First [Ano40v]. Fisher [Mol44, Tho47, Ano47j, Ber43, Fri47]. Fishing [ME42]. Fitting [Cow47, Jon43, Ras49, SA48, Sto41]. Fixing [Wai42]. Flexibility [Mil46]. Fluctuations [War48, Ynt40b]. Food [Bec44a, Mig44, Nor40, Mig44]. Foods [Mar45]. Force [BW46, DH46, Eck45a]. Forecasting [Gar46, Hoe47, Mor49, Wil42b, War45]. Foreclosure [Kap42]. Foreign [Tho48]. Form [OC49]. Formula [Che44b, Sie45]. Formulas [Cor47b, Hoe47, Pie43, Sie41a, Sie41b]. Formulation [Wal45d]. Forrest [LeC46]. Foundations [Kin40]. Founder [Wil40b]. Four [Ric44]. Fox [Dur43]. Fracture [Eps48]. Francis [Sni40]. Frank [Sau40]. Franklin [Ano45o]. Frederick [Wil46a]. Frequencies [San42b]. Frequency [Ric48, Sch47c, Eis46]. Freshmen [Gil44]. Fritz [Alt40a, Lun40]. Front [Ano40n, Ano40o, Ano40p, Ano40q, Ano40r, Ano40s, Ano41b, Ano41c, Ano41d, Ano41e, Ano43d, Ano43e, Ano43f, Ano43g, Ano44d, Ano44e, Ano44f, Ano44g, Ano45h, Ano45i, Ano45j, Ano45k, Ano45l, Ano46b, Ano46c, Ano46d, Ano46e, Ano47c, Ano47d, Ano47e, Ano47f, Ano48a, Ano49a, Ano49b, Ano49c, Ano49d, She46]. Full [JW45, ME42, SW46b, Tae45a, Tol45, Wal45d]. Full-Employment [SW46b]. Function [Ber44, Che44a, DD43, DS42, Sto41]. Functional [Bec44b]. Functions [Bak42, Ban45, Bat45b, EW41, Jon46, Mos41]. Fundamentals [Dem45a]. Funds [Her41]. Furlan [Gei48]. Further [Ano49e, Cam40, Sie41b, Wil44b]. Future [Dun40a].

G [Ano47b, Dem45a, Den44a, Vai40, Ynt40c]. Gains [Vic41, WCH40, Wal41, WCH41]. Gamma [Ban45]. General [Bat42, Che49, Rao47, Shr41, Wal49b, War45, Smi45b]. General-Equilibrium [Smi45b]. Generalized [Sie45]. Genetics [Gei49]. Geoffrey [Olm42]. Geographic [Ano45m]. Geographical [Fal40, Har47]. Geometric [Sie42a, Sza46]. George [Ano45p, Lot44]. German [Fur46, Gei48, Mus40]. Germany [Swe44]. Get [PS49]. Gill [Mye46]. Glover [Cra42]. Glueck [Mer44, Mer44]. Godfrey [Tuk47b]. Golden [Sau40]. Gompertz [Bre49, Sto41]. Goulden [Wil40a]. Government [Con47, Rie42, Leo45]. Governments [Sch45]. Gown [OK49]. Graded [Bli44b]. Graduation [Gre48b]. Graham [Sau40]. Grant [Cor47b, Eps40]. Grant-in-Aid [Cor47b]. Graphic [Spu40, Nag40]. Graphs [Spi48]. Gray [Fri40a]. Great [Mos49a]. Greece [JBKD47]. Greetings [Coa40, Fal40, Vin40]. Grids [Smi41b]. Gross [Rob41a]. Group [GW47]. Grouped [Jon41, Pie43]. Grouped-Distribution [Pie43]. Grouping [DB43]. Groups [DS42, Hen44]. Growth [Cow47, Pro40a]. Guide [Bur45a, Bur48b, Bur48c, Cru46]. Gusztav [Ros41].

H [Bat45a, Fri44a, Lot44, NS45, Olm42, Sch40, Tuk47b, Wil40a]. Haddon [Ano47g]. Halfway [Dew40]. Hall [Ano47g]. Hamburger [Swe44]. Hamilton [Pyl40]. Handled [Bra40b]. Harmoniegesetz [Fur46, Gei48]. Harmony [Fur46, Gei48]. Harold [Fri44a]. Harry [Bat45a]. Hart [Den44a]. Health [Dod46, Mye47, Sch44a]. Heilperin [Par40]. Helen [Nor40]. Henderson [Kel45a]. Henry [Wei44]. Herbert [Tou40]. High [Alg41]. Historical [Fer46, Mor40b, Sol41]. History [Ste48]. Hoffman [Wil46a]. Hog [And47, And48]. Holbrook [Nor40]. Hollerith [Ben47]. Holyoke [Gil44]. Homer [Pro40a]. Homes [PS49]. Honorary [Ano40t]. Hopkins [Vai40, Vai40]. Horace [Dei43]. Hosiery [Hin40]. Hospital [Fra40, Fra42]. Hotel [Ano40w, Ano41f, Ano42b, Ano45p, Ano46f, Ano48b, Ano49f]. Hourly [GL47a, Per40, Rob44, Sie40]. Household [Kis49]. Housing [Bru41, Cot43, Den41, ES43]. Hoyt [Pro40a]. Huhle [Alt40a]. Hultgren [Wil44a]. Human [Cor42, Tuk47b]. Hundredth [Fun41b]. Hyperbolic [Smi41b]. Hypotheses [Gir42a].

Ideal [Sie45]. II [Fri40b, All44, Jes45, Moo44, NS48b, RB47, Wil44a]. III [Gor44, NS48b]. Immer [LeC46]. Impact [Lev42, Mee42]. Implications [Dir44, Eva47, Mos41]. Importance [Alg41, Fra42]. Inadequacies [Jes49]. Income [Eck41, Gil42, Rob41a, Tuc42, Tuc43b, Woo44b, Ynt40a]. Incomes [Den44a]. Incomplete [Rao47]. Increased [CE46]. Index [Ano44a, Bur45a, Bur45b, Bur46b, Bur46c, Bur46d, Bur46e, Bur47c, Bur47d, Bur47e, Bur47b, Bur48a, Bur48b, Bur48c, CJL45, DW41, Fer46, Pet47, Shi42, Sie41a, Sie41b, Sie42a, Sie45, Tep43, Ulm46, Wai44, MBC^{+}43]. Index-Number [Sie41a, Sie41b, Sie42a, Sie45]. Indexes [Cow42, ES42, GL47a, HB45, Joh41a, McI41b, Say41, Tai41, WRS42].

Individual [Pol43]. Individuals [FW40, Sch48b]. Inducão [Tuk47a]. Industrial [FW40, Gar44, Kin40, Lum44, Mee42, Moo44, Smi47, Sog44, Fre47a, TW46a, TW46b]. Industries [Fri44a, Kol40]. Industry [Cat45, Edw41, EW41, Hin40, Mee45, Sim41, Tuc45]. Inequality [Sie43]. Inference [Dem42, Sal42, Tuk49b]. Inflation [Shi40]. Influence [Smi40]. Information [Ano40-34, Ano41k, Ano42h, Ano43o, Ano44m, Ano45t, Ano46i, Ano47l, Ano47m, Ano48d, Ano49g, Ban40, Wal49d]. Ingots [Jul41]. Initiation [Sim41, Sas40a]. Inquiries [HH42]. Inquiry [Hay41a]. Inspection [DG41, Lav46]. Institute [Ric44, Fal40, Nor40]. Instruments [Gru48a, Gru48b]. Insurance [Aar44, Ber40c, Cod48, Dod46, Gue40, Hoh40, Kor40, Len44, Lew48, Mil48, Mor49, Mye47, Mye49b, Nie48, Whi43, WW44]. Integral [Wal40]. Integrating [RHTH45]. Intelligence [RB47]. Inter [Ric44]. Interactions [Fin48]. Intercity [Wil40c]. Interdependence [Knu40, Fur46, Gei48]. Interdependenz [Fur46, Gei48]. Interest [Har47, Tuc43a]. Interior [Lam47]. Internal [Cro44, JW45, Shr43]. International [All44, Bat44, Fos44a, Gor44, Jam40, Len44, Smi45b, Par40]. Interpolation [Gre48b]. Interpretation [DS41b, Wai44]. Interpreting [Lew44]. Interregional [Wil40c]. Interviews [Ban40]. Interwar [GT41]. Inventories [Bro48, DS46, Par48]. Inverse [Bin41]. Inversions [Ros42]. investigation [Fur46, Gei48]. Investment [Mee42]. Investments [McD48]. Involved [Pal41a]. Irving [Ano47j, Fri47, Tho47]. Items [Sch48b]. IV [NS48c].

J [Dem45a, Fri40b, Tuk47a]. Jacob [Smi45b]. James [Dem45a, Den44a, May44, Ric49, Cra42]. Jane [Abb40]. JASA [Sie46]. Jersey [Ano47g]. Jesuits [Sol41]. John [Ano43c, Ano43h, Sch40, Vai40, Wil44a]. Joint [Ano45n]. Joseph [Jen40, Mor40b, Wor40a]. Josiah [Wil41b]. Jr [Vai40]. Jr. [Byr40]. Judge [Bat45b].

Kenneth [May44]. Kenyon [Mus40]. Kingston [Tuk47a]. Kinsey [Wal49a]. Known [Wal49c, Wal49d]. Knoxville [Boe40]. Konjunktur [Nat40]. Krisen [Phe40]. Kulischer [Tae44]. Kurtosis [Kap45]. Kuznets [Cra44].

L [Ano45a, Den44a, Jen40, NS45, Nor40, Smi45b, Swe44, Wil46a, Wil44a, Win44]. L. [Gei48]. Labor [Ano44a, BW46, Dav42, DB45, DH46, Eck45a, Hin47, Lev42, Liv45c, MD43, MBC^{+}43, Mye41a, Mye42, Sie40, Tep43, Tol45, Wat40, Mon40]. Lane [Ric49]. Languages [Dod49]. Large [Dee48, Suk47a, Suk47b, Sza46]. Large-Scale [Suk47b]. Laspeyres [Sie41a]. Latin [Lur46]. Latitude [TW43]. Law [Fur46, Gei48]. Laws [Ano40c, Ano40d, Wor43]. League [Nic42]. Least [CO49, Jon43, Sto41]. Lemuel [Wil40b]. Leo [Sas40a]. Leonard [Byr40, Wer40, Bur47a]. Level [Cav43, Coc45, Jul41, Sza45].

Levels [Hen47, Kir48]. Lewis [Sim43, Wei44, Boe40]. Librarian [Vor41]. Life [AG46, Ber40c, Cod48, Gra45a, Gue40, Lew48, McD48]. like [Vor41]. Likelihood [Ber49]. likes [Lov41]. Limits [Noe49]. Linear [Bak41, Ber49, Gir42a, Jon46]. Lines [Pro42]. Liquidity [Dir44]. List [Ano40u]. Listening [Zei47]. Listing [Ano45a, Ano45b, Ano45m]. Literature [Bur45a, Bur48b, Bur48c]. Living [Ano44a, BM44, MBC^{+}43, Say41, Tep43, TW43, Ulm46, Wil40c, WRS42, Woo41a]. Loan [Har47]. Local [Daw45, Sch45, SL49]. Logarithmic [DS41a]. Logistic [Ber44, Ras49, SA48]. Logits [Ber46]. Lorenz [Smi47]. Losses [Vic41, WCH40, Wal41, WCH41]. Lost [Wal49d]. Love [Bat45a].

M [Ano45b, Nor40, Nug40, Tae44, Sch48b]. M. [Cud40]. Machine [Lan42]. Machines [Ben47, CD41, Dwy40, LA46]. Maclaurin [Lum44]. MacLean [Cud40]. Madow [Ano47b, BG41]. Magnitude [Ric48]. Mail [CF47]. Main [Fin48]. Maintenance [Boe40]. Makeham [Bre49]. Making [Dod46, Mye47]. Male [Mye41b]. Management [Bev45, Eck45a, McP45, Par45]. Manual [War45]. Manuel [Dur43]. Manufacturers [DD43]. Manufacturing [Alg41, Dir44, Sie40]. Many [BO46]. Marco [Jam40]. Margaret [Eps40]. Marginal [Haa47]. Margins [Bro42a]. Mark [Dee48]. Market [Hin47, Wat40, Wil42b]. Marketing [ES43, Mar45, Pyl40]. Markets [And47, And48, McD48]. Marking [Les49, Thu48]. Marriage [Whe46, Byr40]. Massachusetts [Jen40]. Master [Jes45, Kin45]. Material [Lot43]. Materials [Gor44, Moo44]. Mathematical [Ban45, Gei49, McP42, Wal40, Ynt40c]. Matrix [Bin41]. Matter [Ano40a, Ano40b, Ano40n, Ano40o, Ano40p, Ano40q, Ano40r, Ano40s, Ano41a, Ano41b, Ano41c, Ano41d, Ano41e, Ano42a, Ano43a, Ano43d, Ano43e, Ano43f, Ano43g, Ano44b, Ano44d, Ano44e, Ano44f, Ano44g, Ano45e, Ano45f, Ano45h, Ano45i, Ano45j, Ano45k, Ano45l, Ano46b, Ano46c, Ano46d, Ano46e, Ano47a, Ano47c, Ano47d, Ano47e, Ano47f, Ano48a, Ano49a, Ano49b, Ano49c, Ano49d]. Matters [Sch44c]. Maximum [Ber49]. May [Nor40]. McDougall [Wil44a]. Mean [Bro41a, Che44b, Coh49, Com43, Sch40, Sza46]. Meaning [ES42, Joy41]. Means [Che49, Ras49, Sie42a]. Measure [Ber47, NS48c, Shi40]. Measurement [BW46, DB45, Gar44, Liv45c, McC43, Pol43, Sie42b]. Measurements [Gil44, Jon41, Zei47]. Measures [Hir43, Woo49]. Measuring [Ber40c, Bro48, Dod49, Gar46, Gil42, Gru48a, Gru48b, Gue40, Haa47, HH44, Spu40, Wil40c]. Mechanical [Pla43, Pla44]. Mechanization [vH42]. Median [Wal49b]. Medical [Fol42, Lew44, Mil48]. Meeting [Ano40w, Ano40x, Ano40y, Ano41f, Ano42b, Ano45p, Ano46f, Ano47g, Ano48b, Ano49f, Ano49f]. Meetings [Ano40v]. Members [Ano40l, Ano40m, Ano40t, Ano45a, Ano45b, Ano45d, Ano45n]. Membership [Ano40u, Fun41b, HM46, Kel45c, Ste40a]. Mendelian [Gei49]. Mental [Dor40a]. Merits [OC49]. Method [Bin41, BO46, Che49, Com43, Cow43b, DM48, Dod46, Dwy45b, HH44, Jon43, Kap42, Knu43, Lan42, Man40, Mar45, MU49, Mye47, Sch44a, SD49, Smi43, Spu40, SA48, Sto41, Thu48, Wal45a, Woo44b]. Methodology [Bur45a, Bur45b, Bur46b, Bur46c, Bur46d, Bur46e, Bur47c, Bur47d, Bur47e, Bur47b, Bur48a, Bur48b, Bur48c, Phe40]. Methods [Ano46a, Aro44, Bat45a, Cow47, EH44, Haa47, HHG46, Kor40, McV47, Rao47, Sim41, Sim42, WNP43, Wil46b, Woo41a, Tuk49a, Wil40a]. metrical [Fur46, Gei48]. metrische [Fur46, Gei48]. Michael [Par40]. Microclimatic [Wil46b]. Migration [Abb40, FH49b, HH44, JW45, Mye42, Shr43]. Military [Lan41]. Minimum [Ber49, Hin40]. Mining [Fri44a]. Minnesota [Den44a]. Minutes [Ano40v, Ano49f]. Misconception [Sch44c]. Missing [Lur46]. Mitchell [Kuz49]. Mobilization [Wat49]. Modern [CD41, Dwy40, Kin40, Ste48]. Moment [DB43]. Moments [Pie43]. Monetary [Par40]. Monte [MU49]. Montgomery [War45]. Months [Lat42]. Moon [ME42]. Moore [Abb40, Olm42]. Morbidity [Fra40, Fra42]. Mortality [Alt40b, BF47b, Cod48, Gue40, Lew48]. Mosak [Smi45b]. Mother [LS40]. Mount [Gil44]. Movement [Lum44, Whi43, WW44]. Movements [Wil42b]. Moving [Cow42, Spu40]. Mr. [WW43]. Multiple [Bak45, Ben47]. Multiple-Punching [Ben47]. Multiplicative [Shi42]. Multivariate [Tin46]. Myers [Lum44]. Myth [Edw42].

N [Tuk49a, Sch48b]. National [Den41, Gil42, Mye43, Tai41, Wal45d, Ynt40a, Cra44]. Nations [Nic42]. Nature [Hay41a, Lew41a]. Nazi [Swe44]. Needs [Ree44c]. Neglected [Jon46]. Neifeld [Nug40]. Neighborhoods [Pro40a]. Neil [Dor40a]. Neill [Hin43]. Net [HH44, Mye41b, Woo49]. News [Ano40-30, Ano40-31, Ano40-32, Ano40-33]. Newspaperman [Lov41]. Next [Dav40]. No [Bur45a, Bur45b, Bur46b, Bur46c, Bur46d, Bur46e, Bur47c, Bur47d, Bur47e, Bur47b, Bur48a, Bur48b, Bur48c, Sim45]. Nomenclature [NS48c]. Nominating [Ste40c]. Nominations [Ano43b, Ano44c, Fun41a]. Nomograph [Ras49]. Non [HH46, Lot43, Noe49, Rob44]. Non-Agricultural [Rob44]. Non-Material [Lot43]. Non-Parametric [Noe49]. Non-Response [HH46]. Nonfarm [Leb48]. Normal [Coh49, Hen47, Jam40]. Normality [BG41, Mad40]. Norman [Ano43c, Ano43h]. Note [AK49, Ber40b, BG41, Bro41a, Fer46, Fis43, Kel45b, Mad40, Pie44a, Pla44, Ros46, Sie43, Wor46b]. Notes [Ano40-30, Ano40-31, Ano40-32, Ano40-33, Ban45, Sie41b, WW43]. Nugent [Cov40]. Number [Sie41a, Sie41b, Sie42a, Sie45]. Numbers [Coc43, Com43, DW41, Fer46, Gag43, McV47, Pat46, Pet47, Ulm46].

Obituary [Wil49]. Objective [Kis49]. Objectives [DH46]. Observation [Mow41]. Observations [Lur46, Tep43, WW43]. Obtain [Knu42]. Obtaining [Bin41, DM48, HM42]. Occupation [Edw41]. Occupational [Pal41a]. Officer [Pea40b]. Official [Lan44]. Ohio [Ano49f, Win44, Ano46f].

Old [Mye49b, Ric44, Eps40]. Old-Age [Mye49b, Eps40]. One [Lat46]. Operating [AG46]. Operations [McP42]. Opinion [Bev45, McP45, Rop40]. Opportunities [Cat45, Cox45, Leo45, Sch45, Kel45a]. Orchards [McV47]. Order [LS40, Ros42]. Ordering [Sch48b]. Ordnance [Sim42]. Organisation [Par40]. Organization [Sch44a]. Organizing [Bec44b]. Orthogonal [Knu42]. Oscar [Phe40]. Other [Bli44b, Cod48, Lot43, Mad46, Woo49]. Our [Sel40]. Outlook [Nor40]. Output [Fri44a]. Overestimation [Com43]. Ownership [Tuc43b].

P [Bur47a, Hin43, Nor40, Wer40]. Paasche [Sie41a]. Paid [BM44]. Paper [MT49]. Papier [Win44]. Parameters [Bre49]. Parametric [Noe49]. Parity [Whe46]. Part [NS48c]. Partial [Lew41b, Wau46]. Particular [Ber49]. Past [Dun40a, Lub47b]. Paul [Nat40, Tho43]. Pay [DW41, Wil44a]. Payments [Len44]. Peace [Ree44c]. Peach [McV47]. Pearl [Dun41d]. Pearmain [Phe41]. Penicillin [Bli44b]. Pennsylvania [Cot43, Hoy40]. Percentages [Coc43, Rob41a]. Percentiles [Eva42]. Periodicity [AK49, Kel43]. Personal [Nug40]. phenomena [Fur46, Gei48]. Philadelphia [Ano40w, Wat40]. Physical [Bro48, EH44, Gil44, Mil46]. Physics [Ban45]. Pie [Las48]. Pie-Charts [Las48]. Plan [Ano49e]. Planning [All44, Bat44, Lam47, Lan44, SW46b]. Plans [Dun40b, GB49, Lav46, Mil48, Nie48]. Plot [Suk47a, Suk47b]. Point [Dew40, Tai41]. Points [Wer40]. Poisson [Ber40b, Sch47c, Wad45]. Policies [Mus40]. Policy [DH46, Gol44, Tho48, Wal45d, Wor40a, Sie46]. Political [Coa40]. Polynomial [Jon43]. Polynomials [Knu42]. Poole [Mus40]. Pooling [Mos48]. Poor [Bro41a]. Population [Cor47a, ES43, Gra45a, GW47, HH42, Hau42, HR47, JBKD47, MD43, McC43, Shr41, SL49, SDH40, Tae44, Tru41, Tru45, Wal49d, Woo42b]. Populations [Cor42, Cor44, Pea40a, Ull45]. Possible [Mye49b]. Post [Cru46, Hay41b, Lam47, Lan44]. Post-War [Cru46, Lam47, Lan44]. Postwar [Daw45, Dir44, Liv45c]. Potatoes [Bat45b]. Potency [Bli44b]. Power [Fer46, Jac47, Shi40]. Practical [Bra40a, Smi41b, WNP43]. Prairie [San42a, San44b, Wil44b]. Precision [Gru48a, Gru48b]. Predicting [Byr40]. Preliminary [Woo42b]. Presentation [Nag40]. Presented [Lov41]. Presenting [BO46, Don46]. President [Cog40]. Presidents [Lub47b]. Presiding [Pea40b]. Pressure [McC43]. Prewar [Cra44]. Price [Hin47, McI41b, Pet47, Sza45, Tai41, Wai42, Boe40]. Prices [And47, And48, Asc44, BM44, DW41, Kei42, Mil46, Smi40, Tip40, Wor40b]. Principles [Bat42, Boa41]. Privatanstallda [Lun40]. Private [Rob44, Ynt40b]. Probabilities [DS41a, Wor46b]. Probability [Che44a, MT49, Wal40, Tuk49a]. Probits [Ber46]. Problem [Bat44, Cam40, HH46, Sel40, Suk47a, Suk47b]. Problems [Ano46a, Ban45, BW46, Bec44a, Ben47, Bor42, Bur47f, Cam48, Dem42, Dun41c, Eps48, Gei49, Gil40b, HHG46, Hoh40, Kim44, Lan41, McP45, Mil48, RHTH45, Sal42, Wad45, You46]. Procedure [Kis49, SDH40]. Procedures [Boa41, Cam48, Cra45, Ste48]. Proceedings [Ano40w, Ano41f, Ano42b, Ano45p, Ano46f, Ano47g, Ano48b, Ano49f]. Process [Mor40b]. Processing [DG41, EH44]. Produce [Daw45, Liv45c]. Product [Cra44, Gru48b, Mor49]. Production [Alg41, Bec44a, Bro41b, Bru44, CJL45, DD43, Dav43, Gar44, Joh41a, Jul41, Moo44, NS48a, NS48b, NS48c, NS49, Tip40, Vai40, Wor45]. Productive [Sie44]. Productivity [ES42, Eva47, You46, Fri44a]. Products [Gru48a, Wai42, Pyl40]. Professional [Coc45]. Professor [Ber43]. Profile [Spi48]. Profits [Bur47f, May44]. Program [Ano40x, Ano40y, Dee48, Fol42, Hau47, McG44b, Mye49b, SW46b, Tae45a, Tae47, Tai41, Tol45]. Programming [All44, Bat44, Gor44]. Programs [Aar44]. Progress [Gil40b]. Progressions [Nie45, Pie44a, Pie44b]. Projects [Ynt40b]. Propensity [Haa47]. Properties [Che44a]. Proportionate [Las48]. Proposed [Hau42, Sne48]. Prospect [Ano42c]. Providing [Bur47f]. Provinces [San42a, San44b, Wil44b]. Psychological [Kat47]. Psychology [Fri40a]. Public [Nie48, Rop40, Sch44a, Som41, Ynt40b]. Publications [Ano40z, Ano40-27, Ano40-28, Ano40-29, Ano41g, Ano41h, Ano41i, Ano41j, Ano42d, Ano42e, Ano42f, Ano42g, Ano43i, Ano43j, Ano43k, Ano43l, Ano44h, Ano44i, Ano44j, Ano44k, Ano45q, Ano45r, Ano45s, Ano46g, Ano46h, Ano47h, Ano47i, Old46]. Punch [Ber41, San41b]. Punched [BO46, HM42, Knu42, McP42]. Punching [Ben47, DTG42]. Purchases [Tos45]. Purchasing [Fer46, Shi40]. Putting [Boa41].

Quality [Man40, OK49, Ric43, Sch44c, Sim41, Wor45]. Quantities [Mil46]. Quantity [Wai44]. Quarrels [Ric48]. Quarterly [Bur45a, Bur48b, Bur48c]. Questionnaire [NS48c]. Questionnaires [CF47].

R [Ber43, Nug40, Sau40, Ynt40c]. Radio [Zei47]. Railway [Wil44a, Wil44a]. Random [Cor47a, Gag43, Osb42, Ros42, Sol41]. Ranges [GW47]. Rank [Sch48b]. Rate [HH44, LS40, Tuc43a, Woo49]. Rates [FP40, GT41, Har47, Mye41b, Mye43, Rob44, SD49, Whe46, Wil44a]. Ratio [Kir48, Wal49c]. Ration [Woo42b]. Rationing [Mig44]. Rautenstrauch [Bur40a]. Raw [Gor44]. Raymond [Vai40, Dun41d]. Realization [Vic41, WCH40, Wal41, WCH41]. Rearmament [Bro41b]. Received [Ano40z, Ano40-27, Ano40-28, Ano40-29, Ano41g, Ano41h, Ano41i, Ano41j, Ano42d, Ano42e, Ano42f, Ano42g, Ano43i, Ano43j, Ano43k, Ano43l, Ano44h, Ano44i, Ano44j, Ano44k, Ano45q, Ano45r, Ano45s, Ano46g, Ano46h, Ano47h, Ano47i, Old46]. Recording [Old49]. Records [Ban40]. Reduced [OC49]. Reference [Ber49]. Reflections [Lot43]. Regarding [Ano49e]. Region [Har47]. Registration [SD49]. Regression [Bak41, CO49, Dwy45b, Jon46, OC49, Wau43]. Regressions [San42b]. Regressive [OC49]. Regular [Ano40m]. Regulation [Fos44a]. Rejoinder [San44b, WCH41, Woo41b, Zim41]. Related [Aar44, Nic42]. Relation [Bra41a, Den41, DH46, Haf40, Hen47, Jul41, Mad46, Sch47a, Sch47b, TW43, Woo49]. Relations [Koo45]. Relationship [Sza45]. Relationships [CO49, Mye45b]. Relative [HH42, Hen44, Jac47]. Reliability [Cra45]. Remarks [BG41, Cog40, Pea40b, San44a]. Renegotiation [Par45]. Rents [Bor42]. Reply [Ber43, Wil44b, Woo41a, Woo41b, Wor44a]. Report [McD40, Wal49a, McD40]. Reporting [NS48a, NS48b, NS48c, NS49]. Reports [Ano43m, Ano44l, Boa41, EH44, Eng47b, FW40]. Representatives [Ull45]. Represented [Sch47c]. Reproduction [Mye41b, Whe46, Woo49]. Requirements [Dav42, Wat49]. Research [Bev45, Cox45, Gol44, McD40, McP45, Nor40, Ynt40a, Bat45a, Cha49, Phe40]. Reserve [Cam48, GL47a, Joh41a]. Reserves [Sel40, War48]. Resident [HR47]. Residential [Pro40a]. Resources [All44, Bat44, Gor44, Mye43, Wil44a]. Respondent [Kis49]. Response [Ano47j, Ano47k, Bli44b, HH46, Pal43a]. Results [Cra45]. Retro [KB43]. Retrospect [Mer44]. Returns [DG41]. Revenue [Vic41, WCH40, Wal41, WCH41]. Review [Abb40, Abb41, Abb42, Ack41, Ald41, Alt42, Alt44, Alt40a, And45, AS46, AM47, And49, And43, Ark44, Arn45, Arn46, Aro48, Aya43, Ayr42, Bak47, Bak49, Bal45, Ban45, Ban49a, Ban49b, Bar42, BL47, BF47a, Bar41, Bat45a, Bea44, Bel41, Ben42a, Ben42b, Ben43, Ber40a, Ber42a, Bla45a, Bla45b, Bla47, Bli41, Blo43, BV46, Bod41, Boe40, Bow49, Bra48, Bra41b, Bra41c, Bra43, Bra44, Bra46, Bre42, Bro42b, BT46, Bru43, Bue43a, Bue43b, Bur40a, Bur43, BN48, Bur40b, But43, Byr40, Cah49, Cam45, CM49a, Car41, Cas44, Cha42, Cha49, Cha41, Che40a, Chu46, CJ47, Chu49, CMB47, Chu47, Cio41, Cla43, Cli47, CD47a, Coa42, Coc42a, Coc44, CP48, Col45, Con45, Cor49, Cot47, Cov40, Cow43a, Cow44, Cow45, Cow41, Cox46, Cra45, Cra48]. Review [Cra43, Cra44, Cre43a, Cre43b, Cro42, Cru47, Cru49, Cur40a, Cur40b, Cur45, Cur46, Cur47, CD47b, CM49b, Cus43, Dal42, DD47, Dav44a, Dav44b, DWL48, DH49, Dea45, Dee46, Dem44, Dem45b, Dem45c, Dem45a, Dem48, Dem49, Den44a, Den44b, Dew42, Dod45, Dor40a, Dor40b, Dor40c, Dor42, Dou40, Dou41a, Dou41b, Dou41c, DuB44, Dud48, Dun42, Dun41a, Dun41b, Dur41, Dwy45a, Dwy48, Edg40, Eis41, Eis46, Eis47, Ell40, Ely46, End41, End42a, End42b, End43, Eng47a, Eps40, EM43, Eps47, Eps49, Esk44, Eud42, Eva49, Eze42, Eze43, Fab40, Fab41a, Fab41b, Fai42, Fel47, Fer47, FJ47, Fer45, FN47, FH49a, FS48a, Fie43, FY46, Fis46, Fis41, Fis42, Flo41, For41, For44, Fos44b, Fos44a, FM48, Fra41, Fre45, Fre47b, Fre47a, FG48, Fre49, FS48b]. Review [Fri40a, Fri40b, Fri41, Fri42, Fri43, Fri44a, Fri44b, Gag41, Gau46, Gee44, Gei48, Gil41a, Gil43, Gil40a, Gil41b, Gir40, Gir42b, GK46, Gol41, Goo45, Gou44, Gou41, Gou46, Gra41a, Gra41b, Gra45b, Gra46, Gra44, Gra42, Gre41, Gre46, Gre47, Gre48a, Gre49, Gum47, Gut45, GW46, Gut47, GL47b, Hab40, Hae48, Hag43a, Hag41, Hag43b, Hag49, Han40, HL46, Har41, Har42a, Har42b, Har45, Har48, Har42c, Has43, Haw45, HW41, Hed41, Hed45, HW47, Hoh41, Hoo40, Hoo48, Hor49, Hot45, Hov43, Hoy41, Hur46, Hur47, Hut42, Hyd41, Ise42, Jam40, Jam41, Jas41, Jas42, Jen40, Joh40, Joh41b, Joh45, Joy42, JW48, Kac48, Kei45, Kel45a, Kem48, Ken49, Kin40, Kin42, Kir43, Kis43, Kle44, Kno46, Knu49, Kol41, Koo41, Koo47, Kre40, Kuz47a, Kuz47b]. Review [Lai45, Laz44, Lea44, LW47, Leo42, LS48, Lev49, Lew40, Lew45, Li46, Lit46, LT47, LS49, Lit49, Liv45a, Liv47, Loc42a, Loc42b, Lon42, Lor42, Lor43, Lou45, LS47, Luk48, Lum44, Lun41, Lun40, Mac42, Mag42, Mal41, Man45, Man40, Man42, Mar49, May40, May44, May42, McD40, McD41, McD42, McG44a, McI41a, McI42, MW43a, MS43, Men41, Mer43, Mer44, MB48, Mey49, Mig44, Mil43, Mil45, Min41, Mod48, Mol43, Mon40, MS46, Mor40b, MT46, Mos45a, Mos45b, Mos47, Mos49b, Mos49c, Mot49, Mud46, Mue48, Mur45, Mus43, Mus40, Mye44, Mye45a, Mye49a, NS45, Nag40, Nat40, Nei41, Ner40, Nes47, New42, Ney40a, Ney40b, Ney43, Nic43, Nic44, NP48, Nor40, Nor45a, Nug40, Ogl43, Old46, OO48, Olm42, Olm45, Øre48]. Review [Ott42, Pal41b, Pal42, Pal43b, Pal44a, Pal44b, Par42, Par40, Par43, Pea47, Pea49, Pec42, Pet45, Phe40, Phe42, Pit49, Pri43, Pro40a, Pro40b, Pro43, Pyl40, Ree41a, Ree44a, Ree44b, Ren43, Rey41, Rey42a, Rey42b, Rey44, Rid44, Rid47, Rig41, Rig44, Rob49, Rob40, Rob41b, Rob41c, Rom45, Rom49, RW48, Ros46, Sal45, Sam43, San41a, Sas40a, Sas40b, Sat41a, Sat41b, Sau40, Sca41, Sca42, SW46a, Sch46, Sch48a, Sch41a, Sch41b, Sch43, Sch44b, Sch42a, Sch40, Sch42b, Seb49, Sha44, Shi45, Sho43, Sib42, Sie46, Sma40, Smi40, Smi41a, Smi42, Smi45a, Smi46, Smi45b, Sna43, Sne47, Sni40, Sol49, Spe41, Spe44, Spu41, Spu46, SvS48, Spu49, Ste45, Ste41, Sti48, Str46, Swe41, Swe44, Tae44, Taf44, Tho46, Tho45, Tim43, Tin40b, Tin47]. Review [Tin48, Tip40, Tor45, Tou40, Tre41a, Tre41b, Tre41c, Tru44, TW46a, TW46b, Tuk47b, Tuk47a, Tuk49b, Tuk49a, Vai40, Vai41, Vee40, Vil41, Wai47, Wal43a, Wal47a, Wal40, Wal43b, Wal44, Wal45b, Wal47b, Wal42, War45, War43, Wat41b, Wat41a, Wat40, Wel44, Wel41, Wer40, Wes42, Whe47, Wil42a, Wil40a, Wil41a, Wil41c, Wil43, Wil44a, Wil41d, Win44, Wis46, Wol40, Wol43, Wol44, Wol46, Wol49, Woo41a, Woo42a, Wor40a, Wor46a, Wor41, Wor44b, Wor44a, Woy41, Wyn44a, Wyn44b, Yen43, Yer48, Ynt40c, Ynt40a, Ynt42, Yod42, Yod43, Yod44, You43, Zim40, Zim43, Zub49]. Reviews [Phe40]. Revised [Eck45b]. Rhinehart [LeC46]. Rich [RT43]. Rietz [Wei44]. Rising [Dir44]. Risk [Kap42]. Robert [Cro41a]. Robinson [Tou40]. Rods [Gue40]. Roland [Hay41b]. Role [Wal45c]. Rolf [Cov40]. Roosevelt [Ano45o]. Root [Dwy45b, Lan42]. Rose [Win44]. Round [RHTH45]. Roy [Den44a]. Royal [Coa40]. Rupert [Lum44]. Rural [Cot43].

S [Byr40, Mon40, Wel44, Wor40a]. Safety [Bra40b]. Sam [Fri44a]. Sample [Ano46a, Bru44, Cor47a, DG41, DS46, FS42, HHG46, HH46, HHM48, Hau42, Jac47, Jes45, JBKD47, Kin45, Nor44, PS49, Wal49c]. Samples [Cor42, Cor44, DS41b, Kir48]. Sampling [Cam48, Che49, Coc42b, Cod48, Cow46, Dem45a, Dem45d, ES43, Gag43, Goo47, GB49, HB45, HH42, HD43, Hau41, Hen44, Hol43, Lav46, Mad46, McV47, Mos49a, OC49, Osb42, Pro42, Rop40, SK42, SDH40, Ste48, SH48, THD43, Wal45a, WNP43, Wil46b, Wor46b]. Sampling-Units [Coc42b]. Saturday [Ano42b]. Saving [Tuc43a]. Scale [Dee48, Suk47a, Suk47b].

Scales [Cot43, Woo44a]. Schalthebel [Nat40]. Schumpeter [Mor40b]. Schurr [Fri44a]. Science [Coa40, Vin40]. Scores [Bat45b]. Seamless [Hin40]. Seasonal [Cow42, Don46, Jon43, Jul41, Lat42, Shi42, Ynt40b, Nor40]. Seasonality [Spu40]. Second [DB43, Tuk47b]. Secrist [Dei43]. Section [Ano45d, Ano45n]. Securities [Vic41, WCH40, Wal41, WCH41]. Security [Edw42, Eps40, Wel44, Mon40]. Sedley [Mar46]. Seedlings [Bak45]. Selden [Dur46]. Selectees [Lew44]. Selection [Kis49]. Selective [EH44, Fol42, McG44b]. Sensing [Dee48]. Sensitivity [DM48]. Sequence [Sol41]. Sequential [Cow46, Wal45a, BF47a]. Series [AK49, BO46, Eck45b, Haf40, Kel43, Knu40, MW43b, Nie45, Pie44a, Pie44b, Tin40a, Wad45, WM41, Tos45, Olm42]. Service [EH44, McG44b, Fol42]. Settings [Las48]. Severity [CE46]. Share [Rob41a]. Shattuck [Wil40b]. Shaw [Vai40]. Sheldon [Mer44]. Sheppard [Wal40]. Shewhart [Man40]. Short [AK49, Kel43, San42b, Spu40, SA48]. Short-Cut [SA48]. Shortage [Knu43]. Should [Liv45b]. Sickness [Nie48]. Sign [DM46]. Significance [Bak45, BG41, Ber42b, Ber43, Fis43, Kir48, Knu43, MW43b, Mye41b, Olm42, Par48, Sch40, Sne42, WM41, Wal49b, Wil42b]. Signs [MW43b]. Silberling [Ano43c, Ano43h]. Simon [Cra44]. Simplified [Bli44b, Che44b, Cow47]. Simultaneous [Koo45]. Since [Gar44]. Single [Gra45a, GB49]. Situation [Par48]. Six [Ano40v]. Size [Har47, Kir48, Nor44, Suk47a, Suk47b, TW43, Vai40]. Sizes [Coc42b, HHM48, Wal49c]. Skewed [DS41a]. Skewness [Rai42]. Small [Cor47a, Jac47]. Smith [Dem45a, Sch40]. Snider [Jen40]. Snyder [Bur46a, Kin40]. Social [Aar44, Ebl44, Eps40, Hoh40, Lub47a, Mon40, Vin40, Wel44, Whi43, WW44, Fur46, Gei48]. Society [Coa40, Kin40]. Sociological [Cha49]. Sociologisk [Lun40]. Solution [Ber49]. Some [AK49, Aro44, BG41, Gei49, HD43, Hol43, Jes49, Kim44, Lot43, Lur46, Mos41, Mye49b, Pal41a, Smi41b, Tae47, Tin46, Wad45, Wal49b]. Sources [Her41]. Southern [Tae45b]. Soviet [Sch47d]. sozialen [Fur46, Gei48]. Special [Ull45]. Specific [FP40, Kle44]. Square [Ber40b, Ber46, Bli44a, Dwy45b, Nor45b, Sch47a, Sch47b]. Squares [CO49, Com43, Jon43, Lur46, Sto41]. Stability [Cov40]. Stabilization [MO43]. Staffs [Bev45]. Stamp [Wil41b]. Stand [Tai41]. Stand-Point [Tai41]. Standard [Bak41, Coh49, Eva42, GW47, Kol40, Kol44, Pla43, Pla44]. Standardized [Cro41b]. State [Daw45, Gil40b, Hoy40, Sch45, SL49]. States [Fri40b, Gil40a, Mon40, HR47, Kol40, Rob44, Sha43, Sog44, Tuc43b, Wel44]. Station [Sne42]. Statist [Wil40b]. Statistical [Alg41, Alt40b, Ano40d, Ano40-30, Ano40-31, Ano40-32, Ano40-33, Ano45c, Asc44, Bec44b, Bur45a, Bur45b, Bur46b, Bur46c, Bur46d, Bur46e, Bur47c, Bur47d, Bur47e, Bur47b, Bur48a, Bur48b, Bur48c, Coa40, Coc45, Cra45, Cro41b, Dav40, Dem42, DM46, Fal40, Fun41b, Hau47, HM46, Kel45a, Kim44, Koo45, Kor40, Lam47, Lan41, Lat46, Leo45, Les49, Lot43, McG44b, Mil48, Mor40b, Mos41, Nic42, NS48a, NS48b, NS48c, NS49, Ogb40, OK49, Phe40, Ric44, Ros46, Sch45, She46, Sie43, Sim41, Sim42, Sim45, Smi43, Sni40, Ste40b, Tuk49a, Wal45c, Wat49, Wil40a, Wil40b, Wor43, Wor45, vH42, Fri40b, Man40]. Statisticall [Bra40b]. Statistician [Rie42]. Statisticians [Cat45, Hay41a, dSR45]. Statistics [All44, Ban45, BO46, Boa41, Bur47f, Cod48, Con47, Dem45a, Dun40b, Ebl44, Eck41, Eck45a, Edw41, Fol42, Gil40b, Gol46, Hin47, Jac47, Joy41, Lew44, Lew48, Lov41, McD40, Mee45, MBC^{+}43, Mon40, MO43, Mye49b, Pal41a, Par45, Per42, Ree41b, RHTH45, Sch47d, Shr41, SK42, Sne48, Sne49, SW46b, Tae45b, Tae45a, Tae47, Tho48, Tol45, TT44, Tuc45, Vor41, Wal45d, Wal49a, Fur46, Gei48, Ano44a, Hoy40, Lan44, Mor40a, Tep43, Fre47a, Fri40a, McD40, NS45, TW46a]. Statistik [Fur46, Gei48, Alt40a]. Statistique [Sas40a]. Statler [Ano46f, Ano49f]. Status [SL49]. Steel [EW41, Jul41, Lew41a, Mos41]. Stevens [Ano41f]. Stocks [Sza45]. Story [Liv45b]. Straight [Rob44]. Straight-time [Rob44]. Stratford [Ano40w]. Stratification [HB45, THD43]. Stratified [Cor47a]. Stratified-Random [Cor47a]. Structure [Pro40a]. Stuart [Wil49]. Students [Cow46]. Studie [Lun40]. Studies [Mye49b, Nor40, Pal43a, Woo41a, Wat40, Ynt40a]. Study [Ano43n, Dor40a, Fri44a, Lum44, OC49, Sni40]. Subclass [Pat46]. Success [Byr40]. Sum [Aro44, Sch47c]. Summation [Nie45, Pie44a, Pie44b]. Sumner [Phe41]. Sunday [Ano48b]. Supply [Cor45, Hen47, Lev42, Liv45c, MD43, Mig44, Mye41a, Mye42]. Survey [Ano46a, FS42, Goo47, HHG46, Nor40, Tip40]. Surveys [HH46, Osb42, Suk47a, Suk47b]. Survival [HH44]. Survivors [Mye49b]. Svenska [Lun40]. Swedish [Abb40]. Symbols [Cro41b]. Symmetry [Bow48]. Symposium [All44, Bat44, Gor44]. System [Dod46, EH44, McG44b, Mye47]. Systematic [Mad46, Osb42].

T [Fri40a, Tho43]. Table [GB49, RHTH45]. Tables [Bow48, Gra45a, Gue40, LA46, Nor45b, Wal40]. Taxation [May44]. Tea [Fos44a]. Teaching [Cox45, Kim44, OK49]. Technique [Dwy42, Haf40, Knu42]. Tell [Liv45b]. Tendency [Bro41a]. Tennessee [Boe40]. Teoria [Tuk47a]. Teper [WW43]. Terms [Ber49, CO49]. Test [Bak45, Ber40b, Ber47, Bow48, DM46, Olm42, Ros42, WM41, Wal49c, Wal49d]. Testing [Cow46, Dee48, Fri40b, Mar45]. Tests [BG41, Ber42b, Ber43, CE46, Fis43, Jac47, Mad40, MW43b, Sch47a, Sch47b, Sim45, Sne42, Wal49b, Sch40]. Their [Eva47, Woo44a, Nor40, TW43]. Them [Sch40]. Theoretical [Ano49e, Mor40b, Mos41, WNP43]. Theories [Fri40b]. Theory [Coc42b, Dem45a, Eps48, Gir42a, Smi45b, Tuk49a, Ulm46, Tuk49b, War45]. these [Mye49b]. Thirring [Ros41]. Thomson [Tuk47b]. Thor [Wil44a]. Three [Jac47, Sol41, Ynt40a]. Time [AK49, Gol46, Haf40, Kel43, Mil46, MW43b, Nie45, Olm42, Pie44a, Pie44b, Ree44c, Tin40a, WM41, Rob44]. Time-Series [Haf40]. Timoshenko [Nor40]. Tinbergen [Fri40b]. Tippett [Gag43, NS45]. Town [OK49]. Tract [Bat42]. Tracts [Ree42]. Trade [Boe40, Con47, Eng47b, Joh41a, Nor40, Smi45b]. Traffic [Wil44a]. Training [Coc45, Cor45, Dem45d, dSR45]. Transfers [Jam40]. Transform [Ber49]. Transformation [KC47]. Transformations [OC49]. Transverse [Pro42]. Traverse [Pro42]. Treatment [Wor44a]. Trend [LS40]. Trends [Eva47, Jon43, Ogb40, Vai40, Eps40, Vai40]. Trials [Bak45]. Tribute [Dur47, Fri47, Lub47b]. Trout [ME42]. Truncated [Coh49]. Tumbler [BF47b]. Turning [Wer40]. Two [And47, And48, DS42, Kir48, Lub47b, Sch47c, Wal45a]. type [Osb42]. Types [Cow47, DH46].

U.S. [Ano44a, JW45, Leb48, MBC^{+}43]. Unbiased [GW47]. Underwritting [Lew48]. Unemployment [FS42, FH49b, Joy41, Mor49, Win44]. Unequal [Coc42b, Coc43, San42b]. Ungrouped [Dwy40]. Unions [Jen40, Sni40]. Unique [Sne49]. Unit [Mil46, Old49, Sie40]. United [Fri40b, Gil40a, Mon40, Wel44, HR47, Kol40, Rob44, Sha43, Sog44, Tuc43b]. Units [Coc42b, HH42, Hen44, NS48c]. Universe [Lot43]. University [Ano45p]. Unpaid [DB45]. Untersuchung [Fur46, Gei48]. untitled [Abb41, Abb42, Ack41, Ald41, Alt42, Alt44, And45, AS46, AM47, And49, And43, Ano48c, Ark44, Arn45, Arn46, Aro48, Aya43, Ayr42, Bak47, Bak49, Bal45, Ban49a, Ban49b, Bar42, BL47, Bar41, Bea44, Bel41, Ben42a, Ben42b, Ben43, Ber40a, Ber42a, Bla45a, Bla45b, Bla47, Bli41, Blo43, BV46, Bod41, Bow49, Bra48, Bra41b, Bra41c, Bra43, Bra44, Bra46, Bre42, Bro42b, BT46, Bru43, Bue43a, Bue43b, Bur43, BN48, Bur40b, But43, Cah49, Cam45, CM49a, Car41, Cas44, Cha42, Cha41, Che40a, Chu46, CJ47, Chu49, CMB47, Chu47, Cio41, Cla43, Cli47, CD47a, Coa42, Coc42a, Coc44, CP48, Col45, Con45, Cor49, Cot47, Cow43a, Cow44, Cow45, Cow41, Cox46, Cra48, Cra43, Cre43a, Cre43b, Cro42, Cru47, Cru49, Cur40a, Cur40b, Cur45, Cur46]. untitled [Cur47, CD47b, CM49b, Cus43, Dal42, DD47, Dav44a, Dav44b, DWL48, DH49, Dea45, Dee46, Dem44, Dem45b, Dem45c, Dem48, Dem49, Den44b, Dew42, Dod45, Dor40b, Dor40c, Dor42, Dou40, Dou41a, Dou41b, Dou41c, DuB44, Dud48, Dun42, Dun41a, Dun41b, Dur41, Dwy45a, Dwy48, Edg40, Eis41, Eis47, Ell40, Ely46, End41, End42a, End42b, End43, Eng47a, EM43, Eps47, Eps49, Esk44, Eud42, Eva49, Eze42, Eze43, Fab40, Fab41a, Fab41b, Fai42, Fel47, Fer47, FJ47, Fer45, FN47, FH49a, FS48a, Fie43, FY46, Fis46, Fis41, Fis42, Flo41, For41, For44, Fos44b, FM48, Fra41, Fre45, Fre47b, FG48, Fre49, FS48b, Fri41, Fri42, Fri43, Fri44b, Gag41, Gau46, Gee44, Gil41a, Gil43, Gil41b, Gir40, Gir42b, GK46, Gol41, Goo45, Gou44, Gou41, Gou46, Gra41a]. untitled [Gra41b, Gra45b, Gra46, Gra44, Gra42, Gre41, Gre46, Gre47, Gre48a, Gre49, Gum47, Gut45, GW46, Gut47, GL47b, Hab40, Hae48, Hag43a, Hag41, Hag43b, Hag49, Han40, HL46, Har41, Har42a, Har42b, Har45, Har48, Har42c, Has43, Haw45, HW41, Hed41, Hed45, HW47, Hoh41, Hoo40, Hoo48, Hor49, Hot45, Hov43, Hoy41, Hur46, Hur47, Hut42, Hyd41, Ise42, Jam41, Jas41, Jas42, Joh40, Joh41b, Joh45, Joy42, JW48, Kac48, Kei45, Kem48, Ken49, Kin42, Kir43, Kis43, Kno46, Knu49, Kol41, Koo41, Koo47, Kre40, Kuz47a, Kuz47b, Lai45, Laz44, Lea44, LW47, Leo42, LS48, Lev49, Lew40, Lew45, Li46, Lit46, LT47, LS49, Lit49, Liv45a, Liv47, Loc42a, Loc42b, Lon42, Lor42, Lor43, Lou45, LS47, Luk48, Lun41, Mac42, Mag42]. untitled [Mal41, Man45, Man42, Mar49, May40, May42, McD41, McD42, McG44a, McI41a, McI42, MW43a, MS43, Men41, Mer43, MB48, Mey49, Mil43, Mil45, Min41, Mod48, Mol43, MS46, MT46, Mos45a, Mos45b, Mos47, Mos49b, Mos49c, Mot49, Mud46, Mue48, Mur45, Mus43, Mye44, Mye45a, Mye49a, Nei41, Ner40, Nes47, New42, Ney40a, Ney40b, Ney43, Nic43, Nic44, NP48, Nor45a, Ogl43, OO48, Olm45, Øre48, Ott42, Pal41b, Pal42, Pal43b, Pal44a, Pal44b, Par42, Par43, Pea47, Pea49, Pec42, Pet45, Phe42, Pit49, Pri43, Pro40b, Pro43, Ree41a, Ree44a, Ree44b, Ren43, Rey41, Rey42a, Rey42b, Rey44, Rid44, Rid47, Rig41, Rig44, Rob49, Rob40, Rob41b, Rob41c, Rom45, Rom49, RW48, Sal45, Sam43, San41a, Sas40b, Sat41a, Sat41b, Sca41, Sca42]. untitled [SW46a, Sch46, Sch48a, Sch41a, Sch41b, Sch43, Sch44b, Sch42a, Sch42b, Seb49, Sha44, Shi45, Sho43, Sib42, Sma40, Smi41a, Smi42, Smi45a, Smi46, Sna43, Sne47, Sol49, Spe41, Spe44, Spu41, Spu46, SvS48, Spu49, Ste45, Ste41, Sti48, Str46, Swe41, Taf44, Tho46, Tho45, Tim43, Tin40b, Tin47, Tin48, Tor45, Tre41a, Tre41b, Tre41c, Tru44, Vai41, Vee40, Vil41, Wai47, Wal43a, Wal47a, Wal43b, Wal44, Wal45b, Wal47b, Wal42, War43, Wat41b, Wat41a, Wel41, Wes42, Whe47, Wil42a, Wil41a, Wil41c, Wil43, Wil41d, Wis46, Wol40, Wol43, Wol44, Wol46, Wol49, Woo42a, Wor46a, Wor41, Wor44b, Woy41, Wyn44a, Wyn44b, Yen43, Yer48, Ynt42, Yod42, Yod43, Yod44, You43, Zim40, Zim43, Zub49]. Upon [Mee42]. Urban [Mye45b, Old49, Ynt40b]. Use [And47, And48, Ban45, Bat45b, Bra40a, DS42, Dee48, Dwy45b, HH44, Hau41, Jon41, Kin45, KC47, Koz41, Mos49a, Pat46, Ros42, Sch40, Sch47d, Sne42, Wal45d, Wil44a]. Used [EH44]. Useful [Nie45, Pie44a, Pie44b]. Usefulness [MT49]. Uses [Che40b, DH46, ES43, Her41, Hol43, MT49, Ree42, Smi41b, Ste48, SH48, Woo44a]. Using [Wal49d]. Utility [Ber47, Smi43]. Utilization [Lev42, Sie42b, Nor40].

V [Fos44a, Gei48, Kel45a, Nor40, NS49]. Valid [Wal49b]. Validity [Mye41b]. Valuations [Cam48]. Value [Ebl44]. Values [Eps48]. Variability [Gru48a, Gru48b, Pal43a]. Variable [Wau43]. Variables [Bak42, Jon46]. Variance [And47, And48, Bra41a, Coc43, Haf40, Kir48, Sch47a, Sch47b, Wal49d]. Variances [Wal49c]. Variates [Pie43]. Variation [Don46, Ric48]. Variations [Cav43]. Various [HH42]. Verrijn [Wil49]. Versus [Pol43]. Very [Wal49b]. VI [NS49]. Victor [Dur46]. Viewed [Tai41]. Viewpoint [Alg41, Man40]. VII [Wal40]. Vital [Dun40b]. Vitamin [Bli40]. Vol [Dem45a]. Volume [Ano40-34, Ano41k, Ano42h, Ano43o, Ano44m, Ano45t, Ano46i, Ano47l, Ano47m, Ano48d, Ano49g, Fri40b, Wal40, Ynt40a]. Voluntary [Mil48]. Votaw [Fri40a]. Votes [Sch48b].

W [Arm42, Ban45, Boe40, Byr40, Lum44, Man40, Mon40, Olm42, Tou40, Tuk49b, Wal40, Wel44]. W. [Wor44a]. Wage [Cav43, Hin40, MO43, Mye45b, Rob44, Tae45a, Wil44a]. Wage-Rates [Wil44a]. Wages [Kei42]. Wald [BF47a]. Wallis [Olm42]. Walter [Bur40a, Man40, Ano47k, Dur47]. War [CJL45, Cra44, Dav43, Gil42, Lev42, MD43, NS48a, NS48b, NS48c, Par45, Ree44c, RB47, Shr43, Wil44a, Wor40b, Wor45, Cru46, Gol46, Lam47, Lan44, NS49]. War-Time [Gol46]. Wars [Moo44]. Wartime [Bec44a, Mye45b, TT44, WRS42]. Washington [Ano45p]. Waterman [Cra42]. Wealth [Ynt40a, Ynt40a]. Weather [San42a, San44b, Wil44b]. Wednesday [Ano40w]. Weekly [GL47a]. Weinfeld [Den44a]. Wesley [Kuz49]. Wheat [Nor40, San42a, San44b, Wil44b, Nor40]. Wheats [Nor40]. Where [Cav43]. which [Wal49b]. Whittlesey [Sau40]. Wholesale [Asc44, McI41b]. Wickizer [Fos44a]. Wilcox [Sni40]. Willard [Nag40]. Willare [Ano45p]. Willcox [Ano47k, Dur47]. William [Den44a, Ano47b, Ric49]. Wirtschaftspolitik [Alt40a]. Wisconsin [McC43, Mor40a]. within [Kis49]. without [PS49]. Work [Bec44b, Coc45, Eck45a, Leo45, Nic42, Sch45]. Workers [Cav43, Lum44, Kel45a]. Working [Nor40]. World [Lat46, Lub47a, RB47, Tip40, Wil44a, Moo44, Nor40, Tip40]. would [Vor41]. Woytinsky [Mon40, Wel44]. WPA [Ynt40b]. WPB [CJL45]. Written [Ber41].

XXV [Cra45].

Year [Tip40]. Years [Dav40, Ric44]. Yes [Sim45]. Yes-No [Sim45]. Yield [Suk47a, Suk47b]. York [Ano42b, Ano48b, GL47a, Joh41a, Sni40].

Z [Ano45b].

References

Aaronson:1944:BBU

[Aar44] Franklin M. Aaronson. Benefits and beneficiaries under social insurance and related programs. Journal of the American Statistical Association, 39(226):183–196, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

         org/stable/2279948.

Abbott:1940:BRB

[Abb40] W. Lewis Abbott. Book review: Cityward Migration: Swedish Data, by Jane Moore. Journal of the American Statistical Association, 35(209):207–208, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

         stable/2279684.

Abbott:1941:BRB

[Abb41] W. Lewis Abbott. Book review: [untitled]. Journal of the American Statistical Association, 36(216):551–552, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278972.

Abbott:1942:BRB

[Abb42] Charles Cortez Abbott. Book review: [untitled]. Journal of the American Statistical Association, 37(218):298–299, June 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279237.

Ackley:1941:BRB

[Ack41] Gardner Ackley. Book review: [untitled]. Journal of the American Statistical Association, 36(215):448–450, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279636.

Altman:1946:AAO

[AG46] Oscar L. Altman and Charles G. Goor. Actuarial analysis of the operating life of B-29 aircraft engines. Journal of the American Statistical Association, 41(234):190–203, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280486.

Alchian:1949:NSE

[AK49] Armen A. Alchian and Truman L. Kelley. Note on some errors in “the evidence of periodicity in short time series”. Journal of the American Statistical Association, 44(248):559–564, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279907.

Alderfer:1941:BRB

[Ald41] Evan B. Alderfer. Book review: [untitled]. Journal of the American Statistical Association, 36(216):565–566, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278985.

Alger:1941:ISV

[Alg41] P. L. Alger. The importance of the statistical viewpoint in high production manufacturing. Journal of the American Statistical Association, 36(213):50–52, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278812.

Allen:1944:IPD

[All44] R. G. D. Allen. International programming of the distribution of resources: a symposium: II statistics and combined planning. Journal of the American Statistical Association, 39(227):286–290, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280033.

Altschul:1940:BRB

[Alt40a] Eugen Altschul. Book review: Statistik als ein Erkenntnismittel der Wirtschaftspolitik, by Fritz Huhle. Journal of the American Statistical Association, 35(210):429–430, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278736.

Altshuler:1940:DMS

[Alt40b] Constance Wynn Altshuler. Diabetes mortality: a statistical analysis. Journal of the American Statistical Association, 35(210):341–346, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278723.

Altman:1942:BRB

[Alt42] Oscar L. Altman. Book review: [untitled]. Journal of the American Statistical Association, 37(218):299–300, June 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279238.

Altman:1944:BRB

[Alt44] Isidore Altman. Book review: [untitled]. Journal of the American Statistical Association, 39(226):262–263, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279967.

Anderson:1947:BRB

[AM47] T. W. Anderson and Oskar Morgenstern. Book review: [untitled]. Journal of the American Statistical Association, 42(237):187–190, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280204.

Andrews:1943:BRB

[And43] Myron E. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 38(223):374–375, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279797.

Anderson:1945:BRB

[And45] R. L. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 40(231):393–394, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965644.

Anderson:1947:UVC

[And47] R. L. Anderson. Use of variance components in the analysis of hog prices in two markets. Journal of the American Statistical Association, 42(240):612–634, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280018.

Anderson:1948:CUV

[And48] R. L. Anderson. [correction: Use of variance components in the analyses of hog prices in two markets]. Journal of the American Statistical Association, 43(242):242, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280370.

Anderson:1949:BRB

[And49] T. W. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 44(246):323–325, June 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280611.

Anonymous:1940:BMa

[Ano40a] Anonymous. Back matter. Journal of the American Statistical Association, 35(209):212, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279686.

Anonymous:1940:BMb

[Ano40b] Anonymous. Back matter. Journal of the American Statistical Association, 35(211):594, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279305.

Anonymous:1940:L

[Ano40c] Anonymous. By-laws. Journal of the American Statistical Association, 35(210):4–5, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278759.

Anonymous:1940:LAS

[Ano40d] Anonymous. By-laws of the American Statistical Association. Journal of the American Statistical Association, 35(209):305, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279477.

Anonymous:1940:CAb

[Ano40e] Anonymous. Chapter activities. Journal of the American Statistical Association, 35(210):409–422, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278730.

Anonymous:1940:CAc

[Ano40f] Anonymous. Chapter activities. Journal of the American Statistical Association, 35(211):547–554, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279280.

Anonymous:1940:CAd

[Ano40g] Anonymous. Chapter activities. Journal of the American Statistical Association, 35(212):697–706, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279259.

Anonymous:1940:Cb

[Ano40h] Anonymous. Charter. Journal of the American Statistical Association, 35(209):308, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279479.

Anonymous:1940:Cc

[Ano40i] Anonymous. Charter. Journal of the American Statistical Association, 35(210):2, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278757.

Anonymous:1940:Ca

[Ano40j] Anonymous. Circular. Journal of the American Statistical Association, 35(209):303–304, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279476.

Anonymous:1940:Cd

[Ano40k] Anonymous. Constitution. Journal of the American Statistical Association, 35(210):2–4, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278758.

Anonymous:1940:CM

[Ano40l] Anonymous. Corporate members. Journal of the American Statistical Association, 35(210):6, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278761.

Anonymous:1940:FRM

[Ano40m] Anonymous. Fellows and regular members. Journal of the American Statistical Association, 35(210):7–83, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278762.

Anonymous:1940:FMa

[Ano40n] Anonymous. Front matter. Journal of the American Statistical Association, 35(209):??, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279649.

Anonymous:1940:FMb

[Ano40o] Anonymous. Front matter. Journal of the American Statistical Association, 35(209):??, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279460.

Anonymous:1940:FMc

[Ano40p] Anonymous. Front matter. Journal of the American Statistical Association, 35(210):??, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278719.

Anonymous:1940:FMd

[Ano40q] Anonymous. Front matter. Journal of the American Statistical Association, 35(210):??, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278755.

Anonymous:1940:FMe

[Ano40r] Anonymous. Front matter. Journal of the American Statistical Association, 35(211):??, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279270.

Anonymous:1940:FMf

[Ano40s] Anonymous. Front matter. Journal of the American Statistical Association, 35(212):??, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279248.

Anonymous:1940:HM

[Ano40t] Anonymous. Honorary members. Journal of the American Statistical Association, 35(210):6, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278760.

Anonymous:1940:ML

[Ano40u] Anonymous. Membership list, 1840. Journal of the American Statistical Association, 35(209):305–308, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279478.

Anonymous:1940:MFS

[Ano40v] Anonymous. Minutes of the first six meetings. Journal of the American Statistical Association, 35(209):298–302, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279475.

Anonymous:1940:PAM

[Ano40w] Anonymous. Proceedings: 101st Annual Meeting: Bellevue–Stratford Hotel, Philadelphia Wednesday, December 27, 1939. Journal of the American Statistical Association, 35(209):147–180, March 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279664.

Anonymous:1940:PCA

[Ano40x] Anonymous. Program of the centenary anniversary meeting. Journal of the American Statistical Association, 35(209):216, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279462.

Anonymous:1940:PCC

[Ano40y] Anonymous. Program of the centenary celebration at the annual meeting. Journal of the American Statistical Association, 35(209): 241, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279468.

Anonymous:1940:PRa

[Ano40z] Anonymous. Publications received. Journal of the American Statistical Association, 35(209):209–211, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279685.

Anonymous:1940:PRb

[Ano40-27] Anonymous. Publications received. Journal of the American Statistical Association, 35(210):451–454, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278754.

Anonymous:1940:PRc

[Ano40-28] Anonymous. Publications received. Journal of the American Statistical Association, 35(211):590–593, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279304.

Anonymous:1940:PRd

[Ano40-29] Anonymous. Publications received. Journal of the American Statistical Association, 35(212):718–720, December 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279269.

Anonymous:1940:SNNa

[Ano40-30] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 35(209):116–137, March 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279662.

Anonymous:1940:SNNb

[Ano40-31] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 35(210):385–408, June 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278729.

Anonymous:1940:SNNc

[Ano40-32] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 35(211):518–546, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279279.

Anonymous:1940:SNNd

[Ano40-33] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 35(212):674–696, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279258.

Anonymous:1940:VI

[Ano40-34] Anonymous. Volume information. Journal of the American Statistical Association, 35(209):vii–xv, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279648.

Anonymous:1941:BM

[Ano41a] Anonymous. Back matter. Journal of the American Statistical Association, 36(215):464, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279647.

Anonymous:1941:FMa

[Ano41b] Anonymous. Front matter. Journal of the American Statistical Association, 36(213):??, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278806.

Anonymous:1941:FMb

[Ano41c] Anonymous. Front matter. Journal of the American Statistical Association, 36(214):??, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279734.

Anonymous:1941:FMc

[Ano41d] Anonymous. Front matter. Journal of the American Statistical Association, 36(215):??, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279606.

Anonymous:1941:FMd

[Ano41e] Anonymous. Front matter. Journal of the American Statistical Association, 36(216):463–464, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278952.

Anonymous:1941:PAM

[Ano41f] Anonymous. Proceedings: 102nd Annual Meeting: Stevens Hotel, Chicago. Journal of the American Statistical Association, 36(213):123–144, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278823.

Anonymous:1941:PRa

[Ano41g] Anonymous. Publications received. Journal of the American Statistical Association, 36(213):162–165, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278839.

Anonymous:1941:PRb

[Ano41h] Anonymous. Publications received. Journal of the American Statistical Association, 36(214):325–328, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279775.

Anonymous:1941:PRc

[Ano41i] Anonymous. Publications received. Journal of the American Statistical Association, 36(215):460–463, September 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279646.

Anonymous:1941:PRd

[Ano41j] Anonymous. Publications received. Journal of the American Statistical Association, 36(216):580–582, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278998.

Anonymous:1941:VI

[Ano41k] Anonymous. Volume information. Journal of the American Statistical Association, 36(213):iv–x, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278805.

Anonymous:1942:BM

[Ano42a] Anonymous. Back matter. Journal of the American Statistical Association, 37(217):172, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279459.

Anonymous:1942:PAM

[Ano42b] Anonymous. Proceedings: 103rd Annual Meeting: Biltmore Hotel, New York: Saturday, December 27, 1941. Journal of the American Statistical Association, 37(217):117–144, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279439.

Anonymous:1942:P

[Ano42c] Anonymous. The prospect for 1942. Journal of the American Statistical Association, 37(217):12, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279423.

Anonymous:1942:PRa

[Ano42d] Anonymous. Publications received. Journal of the American Statistical Association, 37(217):170–171, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279458.

Anonymous:1942:PRb

[Ano42e] Anonymous. Publications received. Journal of the American Statistical Association, 37(218):310–312, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279247.

Anonymous:1942:PRc

[Ano42f] Anonymous. Publications received. Journal of the American Statistical Association, 37(219):411–414, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279024.

Anonymous:1942:PRd

[Ano42g] Anonymous. Publications received. Journal of the American Statistical Association, 37(220):561–562, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279055.

Anonymous:1942:VI

[Ano42h] Anonymous. Volume information. Journal of the American Statistical Association, 37(217):iii–viii, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279421.

Anonymous:1943:BM

[Ano43a] Anonymous. Back matter. Journal of the American Statistical Association, 38(223):??, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279805.

Anonymous:1943:CN

[Ano43b] Anonymous. Committee on nominations. Journal of the American Statistical Association, 38(223):366, September 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279790.

Anonymous:1943:CNJ

[Ano43c] Anonymous. A correction: Norman John Silberling, 1892–1942. Journal of the American Statistical Association, 38(222):249, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279549.

Anonymous:1943:FMa

[Ano43d] Anonymous. Front matter. Journal of the American Statistical Association, 38(221):??, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279307.

Anonymous:1943:FMb

[Ano43e] Anonymous. Front matter. Journal of the American Statistical Association, 38(222):??, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279533.

Anonymous:1943:FMc

[Ano43f] Anonymous. Front matter. Journal of the American Statistical Association, 38(223):??, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279776.

Anonymous:1943:FMd

[Ano43g] Anonymous. Front matter. Journal of the American Statistical Association, 38(224):??, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279338.

Anonymous:1943:NJS

[Ano43h] Anonymous. Norman John Silberling, 1892–1942. Journal of the American Statistical Association, 38(221):112, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279323.

Anonymous:1943:PRa

[Ano43i] Anonymous. Publications received. Journal of the American Statistical Association, 38(221):140–142, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279337.

Anonymous:1943:PRb

[Ano43j] Anonymous. Publications received. Journal of the American Statistical Association, 38(222):269–270, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279565.

Anonymous:1943:PRc

[Ano43k] Anonymous. Publications received. Journal of the American Statistical Association, 38(223):384–386, September 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279804.

Anonymous:1943:PRd

[Ano43l] Anonymous. Publications received. Journal of the American Statistical Association, 38(224):493–494, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279365.

Anonymous:1943:RA

[Ano43m] Anonymous. Reports of the association. Journal of the American Statistical Association, 38(221):113–124, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279324.

Anonymous:1943:SC

[Ano43n] Anonymous. The study of cycles. Journal of the American Statistical Association, 38(222):248–249, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279548.

Anonymous:1943:VI

[Ano43o] Anonymous. Volume information. Journal of the American Statistical Association, 38(221):iii–viii, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279306.

Anonymous:1944:AUB

[Ano44a] Anonymous. Appraisal of the U.S. Bureau of Labor Statistics cost of living index: Appendix. Journal of the American Statistical Association, 39(225):57–95, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280324.

Anonymous:1944:BM

[Ano44b] Anonymous. Back matter. Journal of the American Statistical Association, 39(227):??, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280061.

Anonymous:1944:CN

[Ano44c] Anonymous. Committee on nominations. Journal of the American Statistical Association, 39(227):390, September 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280045.

Anonymous:1944:FMa

[Ano44d] Anonymous. Front matter. Journal of the American Statistical Association, 39(225):??, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280316.

Anonymous:1944:FMb

[Ano44e] Anonymous. Front matter. Journal of the American Statistical Association, 39(226):??, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279943.

Anonymous:1944:FMc

[Ano44f] Anonymous. Front matter. Journal of the American Statistical Association, 39(227):??, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280030.

Anonymous:1944:FMd

[Ano44g] Anonymous. Front matter. Journal of the American Statistical Association, 39(228):??, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280618.

Anonymous:1944:PRa

[Ano44h] Anonymous. Publications received. Journal of the American Statistical Association, 39(225):134, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280344.

Anonymous:1944:PRb

[Ano44i] Anonymous. Publications received. Journal of the American Statistical Association, 39(226):269–270, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279972.

Anonymous:1944:PRc

[Ano44j] Anonymous. Publications received. Journal of the American Statistical Association, 39(227):409–410, September 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280060.

Anonymous:1944:PRd

[Ano44k] Anonymous. Publications received. Journal of the American Statistical Association, 39(228):544, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280649.

Anonymous:1944:RA

[Ano44l] Anonymous. Reports of the association. Journal of the American Statistical Association, 39(225):102–114, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280329.

Anonymous:1944:VI

[Ano44m] Anonymous. Volume information. Journal of the American Statistical Association, 39(225):iii–viii, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280315.

Anonymous:1945:ALMa

[Ano45a] Anonymous. Alphabetical listing of members: [A–L]. Journal of the American Statistical Association, 40(232):559–599, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279999.

Anonymous:1945:ALMb

[Ano45b] Anonymous. Alphabetical listing of members: [M–Z]. Journal of the American Statistical Association, 40(232):599–632, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280000.

Anonymous:1945:ASA

[Ano45c] Anonymous. The American Statistical Association: A description. Journal of the American Statistical Association, 40(232):555–557, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279998.

Anonymous:1945:AMB

[Ano45d] Anonymous. Associate members of the Biometrics Section. Journal of the American Statistical Association, 40(232):658–662, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280004.

Anonymous:1945:BMa

[Ano45e] Anonymous. Back matter. Journal of the American Statistical Association, 40(229):144, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280481.

Anonymous:1945:BMb

[Ano45f] Anonymous. Back matter. Journal of the American Statistical Association, 40(232):??, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280230.

Anonymous:1945:F

[Ano45g] Anonymous. Fellows. Journal of the American Statistical Association, 40(232):633, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280001.

Anonymous:1945:FMa

[Ano45h] Anonymous. Front matter. Journal of the American Statistical Association, 40(229):??, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280447.

Anonymous:1945:FMb

[Ano45i] Anonymous. Front matter. Journal of the American Statistical Association, 40(230):??, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280118.

Anonymous:1945:FMc

[Ano45j] Anonymous. Front matter. Journal of the American Statistical Association, 40(231):??, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965622.

Anonymous:1945:FMd

[Ano45k] Anonymous. Front matter. Journal of the American Statistical Association, 40(232):??, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280213.

Anonymous:1945:FMe

[Ano45l] Anonymous. Front matter. Journal of the American Statistical Association, 40(232):iii–iv, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279996.

Anonymous:1945:GL

[Ano45m] Anonymous. Geographic listing. Journal of the American Statistical Association, 40(232):634–654, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280002.

Anonymous:1945:JMA

[Ano45n] Anonymous. Joint members of the ASA and Biometrics Section. Journal of the American Statistical Association, 40(232):655–657, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280003.

Anonymous:1945:DFD

[Ano45o] Anonymous. On the death of Franklin Delano Roosevelt. Journal of the American Statistical Association, 40(230):260, June 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280140.

Anonymous:1945:PAM

[Ano45p] Anonymous. Proceedings: 104th Annual Meeting — The George Washington University and Willare Hotel Washington, D.C. Journal of the American Statistical Association, 40(229):102–122, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280463.

Anonymous:1945:PRa

[Ano45q] Anonymous. Publications received. Journal of the American Statistical Association, 40(229):142–143, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280480.

Anonymous:1945:PRb

[Ano45r] Anonymous. Publications received. Journal of the American Statistical Association, 40(230):275–276, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280149.

Anonymous:1945:PRc

[Ano45s] Anonymous. Publications received. Journal of the American Statistical Association, 40(231):415–416, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965655.

Anonymous:1945:VI

[Ano45t] Anonymous. Volume information. Journal of the American Statistical Association, 40(229):iii–viii, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280446.

Anonymous:1946:CPM

[Ano46a] Anonymous. Correction: Problems and methods of the sample survey of business. Journal of the American Statistical Association, 41(236):529, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280573.

Anonymous:1946:FMa

[Ano46b] Anonymous. Front matter. Journal of the American Statistical Association, 41(233):??, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280151.

Anonymous:1946:FMb

[Ano46c] Anonymous. Front matter. Journal of the American Statistical Association, 41(234):??, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280482.

Anonymous:1946:FMc

[Ano46d] Anonymous. Front matter. Journal of the American Statistical Association, 41(235):??, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280251.

Anonymous:1946:FMd

[Ano46e] Anonymous. Front matter. Journal of the American Statistical Association, 41(236):??, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280565.

Anonymous:1946:PAM

[Ano46f] Anonymous. Proceedings: 105th Annual Meeting — Statler Hotel, Cleveland, Ohio. Journal of the American Statistical Association, 41(233):77–110, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280159.

Anonymous:1946:PRa

[Ano46g] Anonymous. Publications received. Journal of the American Statistical Association, 41(233):141–143, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280172.

Anonymous:1946:PRb

[Ano46h] Anonymous. Publications received. Journal of the American Statistical Association, 41(235):413–414, September 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280272.

Anonymous:1946:VI

[Ano46i] Anonymous. Volume information. Journal of the American Statistical Association, 41(233):iii–viii, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280150.

Anonymous:1947:BM

[Ano47a] Anonymous. Back matter. Journal of the American Statistical Association, 42(239):??, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280564.

Anonymous:1947:CCW

[Ano47b] Anonymous. Correction to the citation of William G. Madow. Journal of the American Statistical Association, 42(239):460, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280550.

Anonymous:1947:FMa

[Ano47c] Anonymous. Front matter. Journal of the American Statistical Association, 42(237):??, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280175.

Anonymous:1947:FMb

[Ano47d] Anonymous. Front matter. Journal of the American Statistical Association, 42(238):??, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280650.

Anonymous:1947:FMc

[Ano47e] Anonymous. Front matter. Journal of the American Statistical Association, 42(239):??, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280540.

Anonymous:1947:FMd

[Ano47f] Anonymous. Front matter. Journal of the American Statistical Association, 42(240):??, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280006.

Anonymous:1947:PAM

[Ano47g] Anonymous. Proceedings: 106th Annual Meeting: Chalfonte– Haddon Hall, Atlantic City, New Jersey. Journal of the American Statistical Association, 42(237):134–164, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280194.

Anonymous:1947:PRa

[Ano47h] Anonymous. Publications received. Journal of the American Statistical Association, 42(237):209–210, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280212.

Anonymous:1947:PRb

[Ano47i] Anonymous. Publications received. Journal of the American Statistical Association, 42(239):489–490, September 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280562.

Anonymous:1947:RIF

[Ano47j] Anonymous. Response of irving Fisher. Journal of the American Statistical Association, 42(237):4–5, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280178.

Anonymous:1947:RWF

[Ano47k] Anonymous. Response of Walter F. Willcox. Journal of the American Statistical Association, 42(237):7–10, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280180.

Anonymous:1947:VIa

[Ano47l] Anonymous. Volume information. Journal of the American Statistical Association, 42(237):??, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280174.

Anonymous:1947:VIb

[Ano47m] Anonymous. Volume information. Journal of the American Statistical Association, 42(240):682–686, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280005.

Anonymous:1948:FM

[Ano48a] Anonymous. Front matter. Journal of the American Statistical Association, 43(241):??, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280062.

Anonymous:1948:PAM

[Ano48b] Anonymous. Proceedings: 107th Annual Meeting — Hotel Commodore, New York City: Sunday, December 23, 1947. Journal of the American Statistical Association, 43(242):299–317, June 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280375.

Anonymous:1948:U

[Ano48c] Anonymous. [untitled]. Journal of the American Statistical Association, 43(241):172, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280085.

Anonymous:1948:VI

[Ano48d] Anonymous. Volume information. Journal of the American Statistical Association, 43(244):617–623, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280702.

Anonymous:1949:FMa

[Ano49a] Anonymous. Front matter. Journal of the American Statistical Association, 44(245):??, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280345.

Anonymous:1949:FMb

[Ano49b] Anonymous. Front matter. Journal of the American Statistical Association, 44(246):??, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280593.

Anonymous:1949:FMc

[Ano49c] Anonymous. Front matter. Journal of the American Statistical Association, 44(247):??, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280231.

Anonymous:1949:FMd

[Ano49d] Anonymous. Front matter. Journal of the American Statistical Association, 44(248):??, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279899.

Anonymous:1949:FTC

[Ano49e] Anonymous. Further theoretical considerations regarding the plan for eliminating callbacks. Journal of the American Statistical Association, 44(245):17–31, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280348.

Anonymous:1949:PAM

[Ano49f] Anonymous. Proceedings: 108th Annual Meeting: Hotel Statler, Cleveland, Ohio: Minutes of the Annual Business Meeting. Journal of the American Statistical Association, 44(246):297–310, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280603.

Anonymous:1949:VI

[Ano49g] Anonymous. Volume information. Journal of the American Statistical Association, 44(248):583–590, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279898.

Arkin:1944:BRB

[Ark44] Herbert Arkin. Book review: [untitled]. Journal of the American Statistical Association, 39(226):260–261, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279965.

Armstrong:1942:CWD

[Arm42] Floyd E. Armstrong. Carroll W. Doten, 1871–1942. Journal of the American Statistical Association, 37(220):543, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279040.

Arnold:1945:BRB

[Arn45] Kenneth J. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 40(231):382–384, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965636.

Arnold:1946:BRB

[Arn46] Kenneth J. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 41(236):593–595, December 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280580.

Aroian:1944:SME

[Aro44] Leo A. Aroian. Some methods for the evaluation of a sum. Journal of the American Statistical Association, 39(228):511–515, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280628.

Aroian:1948:BRB

[Aro48] Leo Aroian. Book review: [untitled]. Journal of the American Statistical Association, 43(242):339–340, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280383.

Anderson:1946:BRB

[AS46] R. L. Anderson and Zenon Szatrowski. Book review: [untitled]. Journal of the American Statistical Association, 41(235):398–402, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280265.

Ascher:1944:SAR

[Asc44] Leonard Ascher. A statistical analysis of recent changes in commodity prices at wholesale. Journal of the American Statistical Association, 39(225):35–41, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280321.

Ayars:1943:BRB

[Aya43] Robert D. Ayars. Book review: [untitled]. Journal of the American Statistical Association, 38(224):480–481, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279354.

Ayres:1942:BRB

[Ayr42] Milan V. Ayres. Book review: [untitled]. Journal of the American Statistical Association, 37(219):408–409, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279022.

Baker:1941:LRW

[Bak41] G. A. Baker. Linear regression when the standard deviations of arrays are not all equal. Journal of the American Statistical Association, 36(216):500–506, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278957.

Baker:1942:CBF

[Bak42] G. A. Baker. Correlations between functions of variables. Journal of the American Statistical Association, 37(220):537–539, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279038.

Baker:1945:TSD

[Bak45] G. A. Baker. Test of the significance of the differences of per cents of emergence of seedlings in multiple field trials. Journal of the American Statistical Association, 40(229):93–97, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280460.

Baker:1947:BRB

[Bak47] G. A. Baker. Book review: [untitled]. Journal of the American Statistical Association, 42(238):341–342, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280669.

Baker:1949:BRB

[Bak49] G. A. Baker. Book review: [untitled]. Journal of the American Statistical Association, 44(246):329, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280614.

Balderston:1945:BRB

[Bal45] C. Canby Balderston. Book review: [untitled]. Journal of the American Statistical Association, 40(231):389–390, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965641.

Bancroft:1940:CIR

[Ban40] Gertrude Bancroft. Consistency of information from records and interviews. Journal of the American Statistical Association, 35 (210):377–381, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278727.

Bancroft:1945:BRB

[Ban45] T. A. Bancroft. Book review: The Gamma and Beta Functions: Notes and Problems Designed for Use in Mathematical Statistics and Mathematical Physics, by W. Edwards Deming. Journal of the American Statistical Association, 40(232):527–529, December 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280223.

Bancroft:1949:BRBa

[Ban49a] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 44(245):144–145, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280363.

Bancroft:1949:BRBb

[Ban49b] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 44(247):458–459, September 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280249.

Barnett:1941:BRB

[Bar41] Paul Barnett. Book review: [untitled]. Journal of the American Statistical Association, 36(216):571–572, December 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278990.

Barger:1942:BRB

[Bar42] Harold Barger. Book review: [untitled]. Journal of the American Statistical Association, 37(218):297–298, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279236.

Batschelet:1942:GPT

[Bat42] C. E. Batschelet. General principles of tract delimitation. Journal of the American Statistical Association, 37(218):245–246, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279220.

Batt:1944:IPD

[Bat44] William L. Batt. International programming of the distribution of resources: a symposium: I. The problem of combined planning. Journal of the American Statistical Association, 39(227):281–285, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280032.

Baten:1945:BRB

[Bat45a] W. D. Baten. Book review: Experimental Methods in Agricultural Research, by Harry H. Love. Journal of the American Statistical Association, 40(232):531–532, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280225.

Baten:1945:UDF

[Bat45b] William Dowell Baten. The use of discriminant functions in comparing judges’ scores concerning potatoes. Journal of the American Statistical Association, 40(230):223–228, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280128.

Bean:1944:BRB

[Bea44] Louis H. Bean. Book review: [untitled]. Journal of the American Statistical Association, 39(228):541–542, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280646.

Becker:1944:PEF

[Bec44a] Joseph A. Becker. Problems in estimating food production in wartime. Journal of the American Statistical Association, 39(225):

30–34, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280320.

Becknell:1944:OSW

[Bec44b] Harvey E. Becknell. Organizing statistical work on a functional basis. Journal of the American Statistical Association, 39(227):297–302, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280035.

Bell:1941:BRB

[Bel41] Spurgeon Bell. Book review: [untitled]. Journal of the American Statistical Association, 36(216):559–561, December 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278980.

Benedict:1942:BRB

[Ben42a] Murray R. Benedict. Book review: [untitled]. Journal of the American Statistical Association, 37(218):289–290, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279229.

Bennett:1942:BRB

[Ben42b] M. K. Bennett. Book review: [untitled]. Journal of the American Statistical Association, 37(220):547–548, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279043.

Bennett:1943:BRB

[Ben43] M. K. Bennett. Book review: [untitled]. Journal of the American Statistical Association, 38(222):264–265, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279560.

Benjamin:1947:PMP

[Ben47] Kurt Benjamin. Problems of multiple-punching with Hollerith machines. Journal of the American Statistical Association, 42(237):46– 71, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280188.

Berkson:1940:BRB

[Ber40a] Joseph Berkson. Book review: [untitled]. Journal of the American Statistical Association, 35(211):556–558, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279283.

Berkson:1940:NCS

[Ber40b] Joseph Berkson. A note on the chi-square test, the Poisson and the binomial. Journal of the American Statistical Association, 35(210):362– 367, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278725.

Berry:1940:MCL

[Ber40c] Thomas S. Berry. On measuring the cost of life insurance. Journal of the American Statistical Association, 35(212):631–643, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279252.

Berkson:1941:PCD

[Ber41] Joseph Berkson. A punch card designed to contain written data and coding. Journal of the American Statistical Association, 36 (216):535–538, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278963.

Berkson:1942:BRB

[Ber42a] Joseph Berkson. Book review: [untitled]. Journal of the American Statistical Association, 37(217):145–147, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279440.

Berkson:1942:TSC

[Ber42b] Joseph Berkson. Tests of significance considered as evidence. Journal of the American Statistical Association, 37(219):325–335, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279000.

Berkson:1943:ETS

[Ber43] Joseph Berkson. Experience with tests of significance: a reply to Professor R. A. Fisher. Journal of the American Statistical Association, 38(222):242–246, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279546.

Berkson:1944:ALF

[Ber44] Joseph Berkson. Application of the logistic function to bio-assay. Journal of the American Statistical Association, 39(227):357–365, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280041.

Berkson:1946:ACS

[Ber46] Joseph Berkson. Approximation of chi-square by “probits” and by “logits”. Journal of the American Statistical Association, 41(233):70– 74, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280157.

Berkson:1947:CUM

[Ber47] Joseph Berkson. “cost-utility” as a measure of the efficiency of a test. Journal of the American Statistical Association, 42(238):246– 255, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280654.

Berkson:1949:MML

[Ber49] Joseph Berkson. Minimum χ^{2} and maximum likelihood solution in terms of a linear transform, with particular reference to bio-assay. Journal of the American Statistical Association, 44(246):273–278, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280599.

Bevis:1945:MFS

[Bev45] Joseph C. Bevis. Management of field staffs in the opinion research field. Journal of the American Statistical Association, 40(230):245– 246, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280135.

Barnard:1947:BRB

[BF47a] G. A. Barnard and Robert Ferber. Book review: Sequential Analysis, by Abraham Wald. Journal of the American Statistical Association, 42(240):658–665, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280027.

Brown:1947:TM

[BF47b] George W. Brown and Merrill M. Flood. Tumbler mortality. Journal of the American Statistical Association, 42(240):562–574, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280012.

Berkson:1941:CDM

[BG41] Joseph Berkson and R. C. Geary. Comments on Dr. Madow’s “note on tests of departure from normality” with some remarks concerning tests of significance. Journal of the American Statistical Association, 36(216):539–543, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278964. See [Mad40].

Bingham:1941:NMO

[Bin41] M. D. Bingham. A new method for obtaining the inverse matrix. Journal of the American Statistical Association, 36(216):530–534, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278962.

Barna:1947:BRB

[BL47] T. Barna and H. G. Lewis. Book review: [untitled]. Journal of the American Statistical Association, 42(237):165–170, March 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280195.

Blackburn:1945:BRB

[Bla45a] Raymond F. Blackburn. Book review: [untitled]. Journal of the American Statistical Association, 40(229):139–140, March 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280478.

Blair:1945:BRB

[Bla45b] John M. Blair. Book review: [untitled]. Journal of the American Statistical Association, 40(229):125–127, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280465.

Blankenship:1947:BRB

[Bla47] Albert B. Blankenship. Book review: [untitled]. Journal of the American Statistical Association, 42(240):666–667, December 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280028.

Bliss:1940:FDC

[Bli40] C. I. Bliss. Factorial design and covariance in the biological assay of vitamin D. Journal of the American Statistical Association, 35 (211):498–506, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279276.

Bliss:1941:BRB

[Bli41] Charles A. Bliss. Book review: [untitled]. Journal of the American Statistical Association, 36(216):563–564, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278983.

Bliss:1944:CCS

[Bli44a] C. I. Bliss. A chart of the chi-square distribution. Journal of the American Statistical Association, 39(226):246–248, June 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279956.

Bliss:1944:SCP

[Bli44b] C. I. Bliss. A simplified calculation of the potency of penicillin and other drugs assayed biologically with a graded response. Journal of the American Statistical Association, 39(228):479–487, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280624.

Bloch:1943:BRB

[Blo43] Henry Simon Bloch. Book review: [untitled]. Journal of the American Statistical Association, 38(224):477–478, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279351.

Black:1944:CLF

[BM44] John D. Black and Althea MacDonald. Cost of living on farms and prices paid by farmers. Journal of the American Statistical Association, 39(227):377–386, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280043.

Burt:1948:BRB

[BN48] Cyril Burt and Horace W. Norton. Book review: [untitled]. Journal of the American Statistical Association, 43(242):328–332, June 1948.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280380.

Black:1946:PCM

[BO46] Bertram J. Black and Edward B. Olds. A punched card method for presenting, analyzing, and comparing many series of statistics for areas. Journal of the American Statistical Association, 41 (235):347–355, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280258.

Boatwright:1941:PPP

[Boa41] John W. Boatwright. Principles and procedures for putting across business statistics reports to executives. Journal of the American Statistical Association, 36(214):219–224, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279741.

Boddy:1941:BRB

[Bod41] Francis M. Boddy. Book review: [untitled]. Journal of the American Statistical Association, 36(216):550–551, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278971.

Boer:1940:BRB

[Boe40] Arend E. Boer. Book review: Price Maintenance in Knoxville, Tennessee, under the Tennessee Fair Trade Act of 1937, by Charles W. Lewis. Journal of the American Statistical Association, 35(210):437, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278743.

Borders:1942:PDF

[Bor42] Karl Borders. The problems of determining fair rents. Journal of the American Statistical Association, 37(217):34–40, March 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279426.

Bowker:1948:TSC

[Bow48] Albert H. Bowker. A test for symmetry in contingency tables. Journal of the American Statistical Association, 43(244):572–574, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280710.

Bowker:1949:BRB

[Bow49] Albert H. Bowker. Book review: [untitled]. Journal of the American Statistical Association, 44(245):132–133, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280357.

Brandt:1940:PDM

[Bra40a] A. E. Brandt. Practical difficulties met in the use of experimental designs. Journal of the American Statistical Association, 35(209):101– 106, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279660.

Brayer:1940:CSH

[Bra40b] Edward F. Brayer. Can safety be handled statisticall? Journal of the American Statistical Association, 35(210):382–384, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278728.

Brandt:1941:RBD

[Bra41a] A. E. Brandt. The relation between the design of an experiment and the analysis of variance. Journal of the American Statistical Association, 36(214):283–292, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279749.

Bratt:1941:BRBa

[Bra41b] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 36(214):315–316, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279767.

Bratt:1941:BRBb

[Bra41c] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 36(216):572–573, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278991.

Bratt:1943:BRB

[Bra43] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 38(223):371–372, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279794.

Bratt:1944:BRB

[Bra44] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 39(228):529–531, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280638.

Bratt:1946:BRB

[Bra46] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 41(236):599–602, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280582.

Brandt:1948:BRB

[Bra48] A. E. Brandt. Book review: [untitled]. Journal of the American Statistical Association, 43(242):323–325, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280378.

Brearley:1942:BRB

[Bre42] H. C. Brearley. Book review: [untitled]. Journal of the American Statistical Association, 37(220):546–547, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279042.

Brennan:1949:EPG

[Bre49] J. F. Brennan. Evaluation of parameters in the Gompertz and Makeham equations. Journal of the American Statistical Association, 44(245):116–121, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280354.

Brookner:1941:NMP

[Bro41a] R. J. Brookner. A note on the mean as a poor estimate of central tendency. Journal of the American Statistical Association, 36 (215):410–412, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279617.

Brown:1941:BAR

[Bro41b] Theodore H. Brown. Business approaches to rearmament production control. Journal of the American Statistical Association, 36(213):27– 35, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278809.

Broster:1942:CAM

[Bro42a] E. J. Broster. Correlation analysis by margins. Journal of the American Statistical Association, 37(219):359–366, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279004.

Brown:1942:BRB

[Bro42b] Bishop Brown. Book review: [untitled]. Journal of the American Statistical Association, 37(220):548–549, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279044.

Brown:1948:MPI

[Bro48] Weir M. Brown. Measuring physical inventories. Journal of the American Statistical Association, 43(243):377–390, September 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280275.

Brunsman:1941:HC

[Bru41] Howard G. Brunsman. The housing census of 1940. Journal of the American Statistical Association, 36(215):393–400, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279615.

Brumbaugh:1943:BRB

[Bru43] Martin A. Brumbaugh. Book review: [untitled]. Journal of the American Statistical Association, 38(222):266–267, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279562.

Brunsman:1944:SCC

[Bru44] Howard G. Brunsman. The sample census of congested production areas. Journal of the American Statistical Association, 39 (227):303–310, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280036.

Brown:1946:BRB

[BT46] George W. Brown and Alan E. Treloar. Book review: [untitled]. Journal of the American Statistical Association, 41(236):595–599, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280581.

Buehler:1943:BRBa

[Bue43a] Alfred G. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 38(221):129–130, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279328.

Buehler:1943:BRBb

[Bue43b] Alfred G. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 38(224):479–480, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279353.

Burgess:1940:BRB

[Bur40a] Robert W. Burgess. Book review: The Economics of Business Enterprise, by Walter Rautenstrauch. Journal of the American Statistical Association, 35(210):439–440, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278745.

Burtt:1940:BRB

[Bur40b] Harold E. Burtt. Book review: [untitled]. Journal of the American Statistical Association, 35(212):713–715, December 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279265.

Burke:1943:BRB

[Bur43] Laverne K. Burke. Book review: [untitled]. Journal of the American Statistical Association, 38(224):489–490, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279362.

Buros:1945:SMIa

[Bur45a] Oscar Krisen Buros. Statistical methodology index, no. 1: a quarterly guide to current literature. Journal of the American Statistical Association, 40(231):417–423, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965656.

Buros:1945:SMIb 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280229.

Burgess:1946:CS

[Bur46a] W. Randolph Burgess. Carl Snyder. Journal of the American Statistical Association, 41(234):244–246, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280494.

Buros:1946:SMIa

[Bur46b] Oscar Krisen Buros. Statistical methodology index, no. 3. Journal of the American Statistical Association, 41(233):144–154, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280173.

Buros:1946:SMIb

[Bur46c] Oscar Krisen Buros. Statistical methodology index, no. 4. Journal of the American Statistical Association, 41(234):270–274, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280505.

Buros:1946:SMIc

[Bur46d] Oscar Krisen Buros. Statistical methodology index, no. 5. Journal of the American Statistical Association, 41(235):415–420, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280273.

Buros:1946:SMId

[Bur46e] Oscar Krisen Buros. Statistical methodology index, no. 6. Journal of the American Statistical Association, 41(236):625–635, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280592.

Burgess:1947:LPA

[Bur47a] W. Randolph Burgess. Leonard P. Ayres: An appreciation. Journal of the American Statistical Association, 42(237):128–133, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280193.

Buros:1947:SMId 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280029.

Buros:1947:SMIa

[Bur47c] Oscar Krisen Buros. Statistical methodology index, no. 7. Journal of the American Statistical Association, 42(237):203–208, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280211.

Buros:1947:SMIb

[Bur47d] Oscar Krisen Buros. Statistical methodology index. no. 8. Journal of the American Statistical Association, 42(238):353–355, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280673.

Buros:1947:SMIc

[Bur47e] Oscar Krisen Buros. Statistical methodology index, no. 9. Journal of the American Statistical Association, 42(239):491–496, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280563.

Burr:1947:PPA

[Bur47f] Susan S. Burr. Problems in providing adequate statistics on business profits. Journal of the American Statistical Association, 42 (239):432–438, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280546.

Buros:1948:SMIa

[Bur48a] Oscar Krisen Buros. Statistical methodology index, no. 11. Journal of the American Statistical Association, 43(241):174–199, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280087.

Buros:1948:SMIb

[Bur48b] Oscar Krisen Buros. Statistical methodology index, no. 12: a quarterly guide to current literature. Journal of the American Statistical Association, 43(242):356–360, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280390.

Buros:1948:SMIc

[Bur48c] Oscar Krisen Buros. Statistical methodology index, no. 13: a quarterly guide to current literature. Journal of the American Statistical Association, 43(243):500–508, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280288.

Butz:1943:BRB

[But43] E. L. Butz. Book review: [untitled]. Journal of the American Statistical Association, 38(222):265–266, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279561.

Blommers:1946:BRB

[BV46] Paul J. Blommers and P. E. Vernon. Book review: [untitled]. Journal of the American Statistical Association, 41(234):255–259, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280500.

Bancroft:1946:REP

[BW46] Gertrude Bancroft and Emmett H. Welch. Recent experience with problems of labor force measurement. Journal of the American Statistical Association, 41(235):303–312, September 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280254.

Byrnes:1940:BRB

[Byr40] Agnes M. H. Byrnes. Book review: Predicting Success or Failure in Marriage, by Ernest W. Burgess; Leonard S. Cottrell, Jr. Journal of the American Statistical Association, 35(210):448–449, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278752.

Cahen:1949:BRB

[Cah49] Alfred Cahen. Book review: [untitled]. Journal of the American Statistical Association, 44(247):444–447, September 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280241.

Camp:1940:FCB

[Cam40] Burton H. Camp. Further comments on Berkson’s problem. Journal of the American Statistical Association, 35(210):368–376, June 1940.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278726.

Camp:1945:BRB

[Cam45] Burton H. Camp. Book review: [untitled]. Journal of the American Statistical Association, 40(230):261–263, June 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280141.

Campbell:1948:PSP

[Cam48] George C. Campbell. Problems with sampling procedures for reserve valuations. Journal of the American Statistical Association, 43(243):413–427, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280278.

Carmichael:1941:BRB

[Car41] F. L. Carmichael. Book review: [untitled]. Journal of the American Statistical Association, 36(215):451–452, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279639.

Castendyck:1944:BRB

[Cas44] Elsa Castendyck. Book review: [untitled]. Journal of the American Statistical Association, 39(228):543, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280648.

Catlin:1945:OSI

[Cat45] John B. Catlin. Opportunities for statisticians in industry. Journal of the American Statistical Association, 40(229):75–79, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280457.

Cave:1943:VEW

[Cav43] Roy C. Cave. Variations in expenditures where families of wage earners and clerical workers are classified by economic level. Journal of the American Statistical Association, 38(224):445–452, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279344.

Comrie:1941:CCC

[CD41] L. J. Comrie and P. S. Dwyer. On the calculation of correlation coefficients with modern calculating machines. Journal of the American Statistical Association, 36(215):429–430, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279622.

Clifford:1947:BRBb

[CD47a] Paul C. Clifford and Bernard P. Dudding. Book review: [untitled]. Journal of the American Statistical Association, 42(238):332–335, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280667.

Curtiss:1947:BRBb

[CD47b] J. H. Curtiss and Wilfrid J. Dixon. Book review: [untitled]. Journal of the American Statistical Association, 42(238):324–330, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280664.

Churchman:1946:TIS

[CE46] C. West Churchman and Benjamin Epstein. Tests of increased severity. Journal of the American Statistical Association, 41(236):567–590, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280578.

Clausen:1947:CBM

[CF47] John A. Clausen and Robert N. Ford. Controlling bias in mail questionnaires. Journal of the American Statistical Association, 42 (240):497–511, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280007.

Charlesworth:1941:BRB

[Cha41] James C. Charlesworth. Book review: [untitled]. Journal of the American Statistical Association, 36(214):306–307, June 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279757.

Chapin:1942:BRB

[Cha42] F. Stuart Chapin. Book review: [untitled]. Journal of the American Statistical Association, 37(219):396–397, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279012.

Chapin:1949:BRB

[Cha49] F. Stuart Chapin. Book review: Experimental Designs in Sociological Research. Journal of the American Statistical Association, 44(247):460–462, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280250.

Cherington:1940:BRB

[Che40a] Paul T. Cherington. Book review: [untitled]. Journal of the American Statistical Association, 35(211):583–584, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279298.

Cherington:1940:UCB

[Che40b] Paul T. Cherington. The uses of the census of business. Journal of the American Statistical Association, 35(209):107–115, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279661.

Cheng:1944:NPF

[Che44a] Tseng-Tung Cheng. A new probability function and its properties. Journal of the American Statistical Association, 39(226):243–245, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279955.

Cheng:1944:SFM

[Che44b] Tseng-Tung Cheng. A simplified formula for mean difference. Journal of the American Statistical Association, 39(226):240–242, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279954.

Chevry:1949:CGC

[Che49] Gabriel Chevry. Control of a general census by means of an area sampling method. Journal of the American Statistical Association, 44(247):373–379, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280235.

Churchill:1946:BRB

[Chu46] Edmund Churchill. Book review: [untitled]. Journal of the American Statistical Association, 41(235):403–404, September 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280267.

Churchman:1947:BRBb

[Chu47] C. West Churchman. Book review: [untitled]. Journal of the American Statistical Association, 42(239):487–488, September 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280561.

Churchill:1949:BRB

[Chu49] Edmund Churchill. Book review: [untitled]. Journal of the American Statistical Association, 44(248):572–574, December 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279912.

Ciocco:1941:BRB

[Cio41] Antonio Ciocco. Book review: [untitled]. Journal of the American Statistical Association, 36(214):313–314, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279765.

Churchill:1947:BRB

[CJ47] Edmund Churchill and H. M. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 42(237):195–198, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280208.

Copeland:1945:WIW

[CJL45] Morris A. Copeland, Jerome Jacobson, and Herman Lasken. The WPB index of war production. Journal of the American Statistical Association, 40(230):145–159, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280119.

Clark:1943:BRB

[Cla43] Harold F. Clark. Book review: [untitled]. Journal of the American Statistical Association, 38(224):490–491, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279363.

Clifford:1947:BRBa

[Cli47] Paul C. Clifford. Book review: [untitled]. Journal of the American Statistical Association, 42(237):173–174, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280198.

Camp:1949:BRB

[CM49a] Burton H. Camp and Alexander M. Mood. Book review: [untitled]. Journal of the American Statistical Association, 44(246):313– 314, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280606.

Curtiss:1949:BRB

[CM49b] J. H. Curtiss and E. H. MacNiece. Book review: [untitled]. Journal of the American Statistical Association, 44(245):139–142, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280361.

Churchman:1947:BRBa

[CMB47] C. West Churchman, Joseph Manuele, and Roscoe Byers. Book review: [untitled]. Journal of the American Statistical Association, 42(238):312–315, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280661.

Cochrane:1949:ALS

[CO49] D. Cochrane and G. H. Orcutt. Application of least squares regression to relationships containing auto- correlated error terms. Journal of the American Statistical Association, 44(245):32–61, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280349.

Coats:1940:GRS

[Coa40] R. H. Coats. Greetings of the Royal Statistical Society and the Canadian Political Science Association. Journal of the American Statistical Association, 35(209):275–276, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279473.

Coats:1942:BRB

[Coa42] Robert H. Coats. Book review: [untitled]. Journal of the American Statistical Association, 37(217):166–168, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279456.

Cochran:1942:BRB

[Coc42a] W. G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 37(218):291–292, June 1942. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279231.

Cochran:1942:STW

[Coc42b] W. G. Cochran. Sampling theory when the sampling-units are of unequal sizes. Journal of the American Statistical Association, 37(218):199–212, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279214.

Cochran:1943:AVP

[Coc43] W. G. Cochran. Analysis of variance for percentages based on unequal numbers. Journal of the American Statistical Association, 38 (223):287–301, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279779.

Cochran:1944:BRB

[Coc44] W. G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 39(228):525–527, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280635.

Cochran:1945:TPL

[Coc45] W. G. Cochran. Training at the professional level for statistical work in agriculture and biology. Journal of the American Statistical Association, 40(230):160–166, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280120.

Cody:1948:SEM

[Cod48] Donald D. Cody. Sampling errors in mortality and other statistics in life insurance. Journal of the American Statistical Association, 43(243):442–450, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280280.

Cogswell:1940:RPB

[Cog40] Edmund S. Cogswell. Remarks of the President of the Boston Chapter. Journal of the American Statistical Association, 35(209):219–220, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279464.

Cohen:1949:EMS

[Coh49] A. C. Cohen, Jr. On estimating the mean and standard deviation of truncated normal distributions. Journal of the American Statistical Association, 44(248):518–525, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279903.

Colm:1945:BRB

[Col45] Gerhard Colm. Book review: [untitled]. Journal of the American Statistical Association, 40(229):128–129, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280466.

Comstock:1943:OMS

[Com43] R. E. Comstock. Overestimation of mean squares by the method of expected numbers. Journal of the American Statistical Association, 38(223):335–340, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279784.

Cone:1945:BRB

[Con45] Frederick M. Cone. Book review: [untitled]. Journal of the American Statistical Association, 40(230):268–270, June 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280145.

Constantine:1947:TAG

[Con47] Earl Constantine. Trade associations and government statistics. Journal of the American Statistical Association, 42(237):20–22, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280182.

Cornfield:1942:CBS

[Cor42] Jerome Cornfield. On certain biases in samples of human populations. Journal of the American Statistical Association, 37(217):63–68, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279430.

Cornfield:1944:SFP

[Cor44] Jerome Cornfield. On samples from finite populations. Journal of the American Statistical Association, 39(226):236–239, June 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279953.

Cornell:1945:TSD

[Cor45] Francis G. Cornell. Training to supply the demand. Journal of the American Statistical Association, 40(230):167–171, June 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280121.

Cornell:1947:SRS

[Cor47a] F. G. Cornell. A stratified-random sample of a small finite population. Journal of the American Statistical Association, 42(240):523–532, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280009.

Cornell:1947:GAA

[Cor47b] Francis G. Cornell. Grant-in-aid apportionment formulas. Journal of the American Statistical Association, 42(237):92–104, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280190.

Cornell:1949:BRB

[Cor49] F. G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 44(246):327–329, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280613.

Cottam:1943:HSR

[Cot43] Howard R. Cottam. Housing scales for rural Pennsylvania. Journal of the American Statistical Association, 38(224):406–416, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279340.

Cotterman:1947:BRB

[Cot47] C. W. Cotterman. Book review: [untitled]. Journal of the American Statistical Association, 42(237):184–185, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280202.

Cover:1940:BRB

[Cov40] John H. Cover. Book review: Consumer Credit and Economic Stability, by Rolf Nugent. Journal of the American Statistical Association, 35(210):442–443, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278747.

Cowles:1941:BRB

[Cow41] Alfred Cowles. Book review: [untitled]. Journal of the American Statistical Association, 36(216):555–556, December 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278976.

Cowden:1942:MSI

[Cow42] Dudley J. Cowden. Moving seasonal indexes. Journal of the American Statistical Association, 37(220):523–524, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279036.

Cowan:1943:BRB

[Cow43a] Donald R. G. Cowan. Book review: [untitled]. Journal of the American Statistical Association, 38(223):382–383, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279803.

Cowden:1943:CCD

[Cow43b] Dudley J. Cowden. Correlation concepts and the Doolittle method. Journal of the American Statistical Association, 38(223):327–334, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279783.

Cowden:1944:BRB

[Cow44] Dudley J. Cowden. Book review: [untitled]. Journal of the American Statistical Association, 39(225):124, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280336.

Cowden:1945:BRB

[Cow45] Dudley J. Cowden. Book review: [untitled]. Journal of the American Statistical Association, 40(231):378–379, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965633.

Cowden:1946:ASS

[Cow46] Dudley J. Cowden. An application of sequential sampling to testing students. Journal of the American Statistical Association, 41 (236):547–556, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280576.

Cowden:1947:SMF

[Cow47] Dudley J. Cowden. Simplified methods of fitting certain types of growth curves. Journal of the American Statistical Association, 42 (240):585–590, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280014.

Cox:1945:OTR

[Cox45] Gertrude M. Cox. Opportunities for teaching and research. Journal of the American Statistical Association, 40(229):71–74, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280456.

Cox:1946:BRB

[Cox46] Gertrude M. Cox. Book review: [untitled]. Journal of the American Statistical Association, 41(233):135–137, March 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280170.

Cohen:1948:BRB

[CP48] A. C. Cohen, Jr. and Paul Peach. Book review: [untitled]. Journal of the American Statistical Association, 43(242):350–352, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280387.

Craig:1942:JWG

[Cra42] Cecil C. Craig. James Waterman Glover, 1868–1941. Journal of the American Statistical Association, 37(217):116, March 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279438.

Crane:1943:BRB

[Cra43] Robert I. Crane. Book review: [untitled]. Journal of the American Statistical Association, 38(224):488–489, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279361.

Crawford:1944:BRB

[Cra44] John M. Crawford. Book review: National Product, War and Prewar, by Simon Kuznets. Journal of the American Statistical Association, 39(227):392–393, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280047.

Craig:1945:BRB

[Cra45] C. C. Craig. Book review: Culture Element Distributions: XXV, Reliability of Statistical Procedures and Results, by C. Douglas Chretien. Journal of the American Statistical Association, 40(232):526–527, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280222.

Craig:1948:BRB

[Cra48] Cecil C. Craig. Book review: [untitled]. Journal of the American Statistical Association, 43(242):352–353, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280388.

Creamer:1943:BRBa

[Cre43a] Daniel Creamer. Book review: [untitled]. Journal of the American Statistical Association, 38(224):475–476, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279349.

Creamer:1943:BRBb

[Cre43b] Daniel Creamer. Book review: [untitled]. Journal of the American Statistical Association, 38(224):481–482, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279355.

Croxton:1941:REC

[Cro41a] Frederick E. Croxton. Robert Emmet Chaddock, 1879–1940. Journal of the American Statistical Association, 36(213):116–119, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278820.

Croxton:1941:TSS

[Cro41b] Frederick E. Croxton. Toward standardized symbols for basic statistical concepts. Journal of the American Statistical Association, 36(215):426–428, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279621.

Cromwell:1942:BRB

[Cro42] Richard O. Cromwell. Book review: [untitled]. Journal of the American Statistical Association, 37(219):406–407, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279020.

Crowder:1944:CIC

[Cro44] Edward T. Crowder. Centralized internal control of data collection by Federal agencies. Journal of the American Statistical Association, 39(226):155–164, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279946.

Crow:1945:CD

[Cro45] James F. Crow. A chart of the χ^{2} and t distributions. Journal of the American Statistical Association, 40(231):376, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965631.

Crump:1946:BNP

[Cru46] Norman Crump. Britain’s new post-War economic guide. Journal of the American Statistical Association, 41(234):171–172, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280484.

Crump:1947:BRB

[Cru47] S. Lee Crump. Book review: [untitled]. Journal of the American Statistical Association, 42(240):638–639, December 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280021.

Crump:1949:BRB

[Cru49] S. Lee Crump. Book review: [untitled]. Journal of the American Statistical Association, 44(247):455–456, September 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280247.

Cudmore:1940:MCM

[Cud40] S. A. Cudmore. M. C. MacLean. Journal of the American Statistical Association, 35(211):554–555, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279282.

Curtis:1940:BRBa

[Cur40a] William R. Curtis. Book review: [untitled]. Journal of the American Statistical Association, 35(211):568–570, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279289.

Curtis:1940:BRBb

[Cur40b] William R. Curtis. Book review: [untitled]. Journal of the American Statistical Association, 35(211):570–572, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279290.

Curtiss:1945:BRB

[Cur45] J. H. Curtiss. Book review: [untitled]. Journal of the American Statistical Association, 40(231):397–399, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965647.

Curtiss:1946:BRB

[Cur46] J. H. Curtiss. Book review: [untitled]. Journal of the American Statistical Association, 41(233):133–135, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280169.

Curtiss:1947:BRBa

[Cur47] J. H. Curtiss. Book review: [untitled]. Journal of the American Statistical Association, 42(238):315–318, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280662.

Cusick:1943:BRB

[Cus43] James F. Cusick. Book review: [untitled]. Journal of the American Statistical Association, 38(222):256–257, June 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279554.

Daly:1942:BRB

[Dal42] Michael Daly. Book review: [untitled]. Journal of the American Statistical Association, 37(220):544–545, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279041.

Davis:1940:NYA

[Dav40] Joseph S. Davis. The next 100 years of the American statistical association. Journal of the American Statistical Association, 35(209):261– 272, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279471.

Davenport:1942:ELR

[Dav42] Donald H. Davenport. Estimating labor requirements. Journal of the American Statistical Association, 37(217):48–53, March 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279428.

Davis:1943:CPW

[Dav43] Shelby Cullom Davis. Coordinating production for war. Journal of the American Statistical Association, 38(224):417–424, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279341.

David:1944:BRB

[Dav44a] Paul T. David. Book review: [untitled]. Journal of the American Statistical Association, 39(228):534, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280642.

Davis:1944:BRB

[Dav44b] Joseph S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 39(225):115–116, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280330.

Dawson:1945:ELP

[Daw45] Harris P. Dawson, Jr. Expanding local postwar employment data to produce state estimates. Journal of the American Statistical Association, 40(230):197–204, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280125.

Davies:1943:SMC

[DB43] G. R. Davies and Nancy Bruner. A second moment correction for grouping. Journal of the American Statistical Association, 38(221): 63–68, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279313.

Ducoff:1945:EMU

[DB45] Louis J. Ducoff and Gertrude Bancroft. Experiment in the measurement of unpaid family labor in agriculture. Journal of the American Statistical Association, 40(230):205–213, June 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280126.

Daly:1943:PFC

[DD43] Patricia Daly and Paul H. Douglas. The production function for Canadian manufacturers. Journal of the American Statistical Association, 38(222):178–186, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279537.

Daly:1947:BRB

[DD47] Joseph F. Daly and Walter L. Deemer, Jr. Book review: [untitled]. Journal of the American Statistical Association, 42(238):318–324, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280663.

Dean:1945:BRB

[Dea45] Prentice N. Dean. Book review: [untitled]. Journal of the American Statistical Association, 40(229):130, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280468.

Deemer:1946:BRB

[Dee46] Walter L. Deemer, Jr. Book review: [untitled]. Journal of the American Statistical Association, 41(236):609–610, December 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280585.

Deemer:1948:UMS

[Dee48] Walter L. Deemer, Jr. The use of mark sensing in a large scale testing program. Journal of the American Statistical Association, 43(241): 40–52, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280065.

Deibler:1943:HS

[Dei43] F. S. Deibler. Horace Secrist, 1881–1943. Journal of the American Statistical Association, 38(223):365–366, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279789.

Deming:1942:CPS

[Dem42] W. Edwards Deming. On a classification of the problems of statistical inference. Journal of the American Statistical Association, 37(218):173–185, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279212.

Deming:1944:BRB

[Dem44] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 39(225):119–122, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280334.

Deming:1945:BRBc

[Dem45a] W. Edwards Deming. Book review: Fundamentals of the Theory of Statistics: Vol. 1, Elementary Statistics and Applications; Vol. 2, Sampling Statistics and Applications, by James G. Smith; Acheson J. Duncan. Journal of the American Statistical Association, 40(232):534–536, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280227.

Deming:1945:BRBa

[Dem45b] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 40(230):263–265, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280142.

Deming:1945:BRBb

[Dem45c] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 40(231):403–404, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965649.

Deming:1945:TS

[Dem45d] W. Edwards Deming. On training in sampling. Journal of the American Statistical Association, 40(231):307–316, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965625.

Deming:1948:BRB

[Dem48] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 43(242):346–350, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280386.

Deming:1949:BRB

[Dem49] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 44(248):580–582, December 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279916.

Dennis:1941:HRN

[Den41] Samuel J. Dennis. Housing in relation to national defense. Journal of the American Statistical Association, 36(213):36–44, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278810.

Denison:1944:BRB

[Den44a] Edward F. Denison. Book review: Analyses of Minnesota Incomes, 1938–39, by Roy G. Blakey; William Weinfeld; James E. Dugan; Alex L. Hart. Journal of the American Statistical Association, 39 (227):393–394, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280048.

Densen:1944:BRB

[Den44b] Paul M. Densen. Book review: [untitled]. Journal of the American Statistical Association, 39(228):524–525, December 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280634.

Dewey:1940:HP

[Dew40] Davis R. Dewey. The halfway point. Journal of the American Statistical Association, 35(209):236–240, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279467.

Dewey:1942:BRB

[Dew42] Ralph L. Dewey. Book review: [untitled]. Journal of the American Statistical Association, 37(219):407–408, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279021.

Deming:1941:SIP

[DG41] W. Edwards Deming and Leon Geoffrey. On sample inspection in the processing of census returns. Journal of the American Statistical Association, 36(215):351–360, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279609.

Ducoff:1946:OUT

[DH46] Louis J. Ducoff and Margaret Jarman Hagood. Objectives, uses and types of labor force data in relation to economic policy. Journal of the American Statistical Association, 41(235):293–302, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280253.

DeLury:1949:BRB

[DH49] D. B. DeLury and Boyd Harshbarger. Book review: [untitled]. Journal of the American Statistical Association, 44(246):317–320, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280609.

Dirks:1944:RLM

[Dir44] Frederick C. Dirks. The rising liquidity of manufacturing companies and its implications for financing postwar conversion. Journal of the American Statistical Association, 39(226):207–217, June 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279950.

Dixon:1946:SST

[DM46] W. J. Dixon and A. M. Mood. The statistical sign test. Journal of the American Statistical Association, 41(236):557–566, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280577.

Dixon:1948:MOA

[DM48] W. J. Dixon and A. M. Mood. A method for obtaining and analyzing sensitivity data. Journal of the American Statistical Association, 43(241):109–126, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280071.

Dodge:1945:BRB

[Dod45] H. F. Dodge. Book review: [untitled]. Journal of the American Statistical Association, 40(231):411–413, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965653.

Dodd:1946:MMA

[Dod46] Paul A. Dodd. A method of making actuarial estimates for a compulsory health insurance system. Journal of the American Statistical Association, 41(233):58–69, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280156.

Dodd:1949:ML

[Dod49] Stuart C. Dodd. On measuring languages. Journal of the American Statistical Association, 44(245):77–88, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280351.

Donnahoe:1946:PSV

[Don46] Alan S. Donnahoe. Presenting seasonal variation to the business executive. Journal of the American Statistical Association, 41(236):468–471, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280569.

Dorn:1940:BRBa

[Dor40a] Harold F. Dorn. Book review: New Facts on Mental Disorders, Study of 89,190 Cases, by Neil A. Dayton. Journal of the American Statistical Association, 35(210):449–450, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278753.

Dorn:1940:BRBb

[Dor40b] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 35(211):589–590, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279303.

Dorn:1940:BRBc

[Dor40c] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 35(212):716–717, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279268.

Dorn:1942:BRB

[Dor42] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 37(220):552, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279047.

Douglas:1940:BRB

[Dou40] Paul H. Douglas. Book review: [untitled]. Journal of the American Statistical Association, 35(212):707–709, December 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279260.

Douglas:1941:BRB

[Dou41a] Paul H. Douglas. Book review: [untitled]. Journal of the American Statistical Association, 36(215):443–444, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279632.

Douty:1941:BRBa

[Dou41b] H. M. Douty. Book review: [untitled]. Journal of the American Statistical Association, 36(213):155–156, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278834.

Douty:1941:BRBb

[Dou41c] H. M. Douty. Book review: [untitled]. Journal of the American Statistical Association, 36(215):451, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279638.

Davies:1941:PLS

[DS41a] G. R. Davies and R. H. Smith. Probabilities in logarithmic skewed distributions. Journal of the American Statistical Association, 36 (216):493–499, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278956.

Deming:1941:ICS

[DS41b] W. Edwards Deming and Frederick F. Stephan. On the interpretation of censuses as samples. Journal of the American Statistical Association, 36(213):45–49, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278811.

Day:1942:UDF

[DS42] Besse B. Day and Marion M. Sandomire. Use of the discriminant function for more than two groups. Journal of the American Statistical Association, 37(220):461–472, December 1942. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279030.

Deming:1946:DSD

[DS46] W. Edwards Deming and Willard Simmons. On the design of a sample for dealers’ inventories. Journal of the American Statistical Association, 41(233):16–33, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280153.

Rodrigues:1945:TS

[dSR45] Milton da Silva Rodrigues. On the training of statisticians. Journal of the American Statistical Association, 40(230):172–174, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280122.

Deming:1942:ECP

[DTG42] W. Edwards Deming, Benjamin J. Tepping, and Leon Geoffrey. Errors in card punching. Journal of the American Statistical Association, 37(220):525–536, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279037.

DuBrul:1944:BRB

[DuB44] Stephen M. DuBrul. Book review: [untitled]. Journal of the American Statistical Association, 39(226):266–267, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279970.

Dudley:1948:BRB

[Dud48] John W. Dudley, Jr. Book review: [untitled]. Journal of the American Statistical Association, 43(241):172–173, March 1948. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280086.

Dunn:1940:CPF

[Dun40a] Halbert L. Dunn. Census — past and future. Journal of the American Statistical Association, 35(209):242–251, March 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279469.

Dunn:1940:PVS

[Dun40b] Halbert L. Dunn. Plans for vital statistics in 1940. Journal of the American Statistical Association, 35(209):86–92, March 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279658.

Dunn:1941:BRBa

[Dun41a] Halbert L. Dunn. Book review: [untitled]. Journal of the American Statistical Association, 36(213):152–153, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278831.

Dunn:1941:BRBb

[Dun41b] Halbert L. Dunn. Book review: [untitled]. Journal of the American Statistical Association, 36(214):310–312, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279762.

Dunn:1941:PCA

[Dun41c] Halbert L. Dunn. Problems concerning the availability of census data. Journal of the American Statistical Association, 36(214):240– 247, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279744.

Dunn:1941:RP

[Dun41d] Halbert L. Dunn. Raymond Pearl, 1879–1940. Journal of the American Statistical Association, 36(213):120–121, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278821.

Duncan:1942:BRB

[Dun42] Kenneth Duncan. Book review: [untitled]. Journal of the American Statistical Association, 37(217):151, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279444.

Durand:1941:BRB

[Dur41] E. Dana Durand. Book review: [untitled]. Journal of the American Statistical Association, 36(216):548–549, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278969.

Durand:1943:MF

[Dur43] E. Dana Durand. A. Manuel Fox, 1889–1942. Journal of the American Statistical Association, 38(221):105–106, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279320.

Durand:1946:DVS

[Dur46] E. Dana Durand. Dr. Victor Selden Clark. Journal of the American Statistical Association, 41(235):390–392, September 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280261.

Durand:1947:TWF

[Dur47] E. Dana Durand. Tribute to Walter F. Willcox. Journal of the American Statistical Association, 42(237):5–7, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280179.

Dean:1941:AIN

[DW41] Joel Dean and Mary Hilton Wise. An appraisal of index numbers of prices farmers pay. Journal of the American Statistical Association, 36(214):210–218, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279740.

DeLury:1948:BRB

[DWL48] D. B. DeLury, A. E. R. Westman, and E. L. Lahman. Book review: [untitled]. Journal of the American Statistical Association, 43(242):318–321, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280376.

Dwyer:1940:CCC

[Dwy40] P. S. Dwyer. The calculation of correlation coefficients from ungrouped data with modern calculating machines. Journal of the American Statistical Association, 35(212):671–673, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279257.

Dwyer:1942:RDC

[Dwy42] Paul S. Dwyer. Recent developments in correlation technique. Journal of the American Statistical Association, 37(220):441–460, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279029.

Dwyer:1945:BRB

[Dwy45a] Paul S. Dwyer. Book review: [untitled]. Journal of the American Statistical Association, 40(231):409–410, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965652.

Dwyer:1945:SRM

[Dwy45b] Paul S. Dwyer. The square root method and its use in correlation and regression. Journal of the American Statistical Association, 40(232):493–503, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280218.

Dwyer:1948:BRB

[Dwy48] Paul S. Dwyer. Book review: [untitled]. Journal of the American Statistical Association, 43(241):136–137, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280074.

Ebling:1944:SVA

[Ebl44] Walter H. Ebling. The social value of agricultural statistics. Journal of the American Statistical Association, 39(225):21–29, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280319.

Eckler:1941:EIS

[Eck41] A. Ross Eckler. Employment and income statistics. Journal of the American Statistical Association, 36(215):381–386, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279613.

Eckler:1945:MFW

[Eck45a] A. Ross Eckler. Management of field work and collection of statistics of the labor force. Journal of the American Statistical Association, 40(230):249–250, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280137.

Eckler:1945:RCS

[Eck45b] A. Ross Eckler. The revised census series of current employment estimates. Journal of the American Statistical Association, 40(230):187– 196, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280124.

Edgerton:1940:BRB

[Edg40] Harold A. Edgerton. Book review: [untitled]. Journal of the American Statistical Association, 35(212):716, December 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279266.

Edwards:1941:OIS

[Edw41] Alba M. Edwards. Occupation and industry statistics. Journal of the American Statistical Association, 36(215):387–392, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279614.

Edwards:1942:MSA

[Edw42] George W. Edwards. The myth of the security affiliate. Journal of the American Statistical Association, 37(218):225–232, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279217.

Edwards:1944:MUP

[EH44] Thomas I. Edwards and Louis P. Hellman. Methods used in processing data from the physical examination reports of the selective service system. Journal of the American Statistical Association, 39(226):165– 182, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279947.

Eisenhart:1941:BRB

[Eis41] Churchill Eisenhart. Book review: [untitled]. Journal of the American Statistical Association, 36(216):568–569, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278987.

Eisenhart:1946:BRB

[Eis46] Churchill Eisenhart. Book review: Frequency Arrays. Journal of the American Statistical Association, 41(236):620–621, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280589.

Eisenhart:1947:BRB

[Eis47] Churchill Eisenhart. Book review: [untitled]. Journal of the American Statistical Association, 42(239):468–473, September 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280553.

Ellis:1940:BRB

[Ell40] Howard S. Ellis. Book review: [untitled]. Journal of the American Statistical Association, 35(211):575–577, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279294.

Ely:1946:BRB

[Ely46] Owen Ely. Book review: [untitled]. Journal of the American Statistical Association, 41(236):618–619, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280588.

Epstein:1943:BRB

[EM43] Ralph C. Epstein and John H. Myers. Book review: [untitled]. Journal of the American Statistical Association, 38(221):136–138, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279335.

Endler:1941:BRB

[End41] Oscar L. Endler. Book review: [untitled]. Journal of the American Statistical Association, 36(214):318–319, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279770.

Endler:1942:BRBa

[End42a] Oscar L. Endler. Book review: [untitled]. Journal of the American Statistical Association, 37(217):168–169, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279457.

Endler:1942:BRBb

[End42b] Oscar L. Endler. Book review: [untitled]. Journal of the American Statistical Association, 37(220):558–559, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279053.

Endler:1943:BRB

[End43] Oscar L. Endler. Book review: [untitled]. Journal of the American Statistical Association, 38(222):260–261, June 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279557.

Engelhart:1947:BRB

[Eng47a] Max D. Engelhart. Book review: [untitled]. Journal of the American Statistical Association, 42(237):179–180, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280200.

England:1947:FTC

[Eng47b] William H. England. The Federal Trade Commission’s corporation reports. Journal of the American Statistical Association, 42(237):22– 24, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280183.

Epstein:1940:BRB

[Eps40] Abraham Epstein. Book review: Old-Age Security, Social and Financial Trends, by Margaret Grant. Journal of the American Statistical Association, 35(210):447–448, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278751.

Epstein:1947:BRB

[Eps47] Benjamin Epstein. Book review: [untitled]. Journal of the American Statistical Association, 42(237):200–202, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280210.

Epstein:1948:ATE

[Eps48] Benjamin Epstein. Application of the theory of extreme values in fracture problems. Journal of the American Statistical Association, 43(243):403–412, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280277.

Epstein:1949:BRB

[Eps49] Benjamin Epstein. Book review: [untitled]. Journal of the American Statistical Association, 44(246):326–327, June 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280612.

Evans:1942:MPI

[ES42] W. Duane Evans and Irving H. Siegel. The meaning of productivity indexes. Journal of the American Statistical Association, 37(217):103–111, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279436.

Eckler:1943:MSU

[ES43] A. Ross Eckler and E. P. Staudt. Marketing and sampling uses of population and housing data. Journal of the American Statistical Association, 38(221):87–92, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279316.

Eskin:1944:BRB

[Esk44] Leonard Eskin. Book review: [untitled]. Journal of the American Statistical Association, 39(228):535–539, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280644.

Eudey:1942:BRB

[Eud42] Mark Eudey. Book review: [untitled]. Journal of the American Statistical Association, 37(220):557–558, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279052.

Evans:1942:SEP

[Eva42] W. Duane Evans. The standard error of percentiles. Journal of the American Statistical Association, 37(219):367–376, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279005.

Evans:1947:RPT

[Eva47] W. D. Evans. Recent productivity trends and their implications. Journal of the American Statistical Association, 42(238):211–223, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280651.

Evans:1949:BRB

[Eva49] Henry K. Evans. Book review: [untitled]. Journal of the American Statistical Association, 44(247):451–452, September 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280244.

Ezekiel:1941:CFS

[EW41] Mordecai Ezekiel and Kathryn H. Wylie. Cost functions for the steel industry. Journal of the American Statistical Association, 36(213): 91–99, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278817.

Ezekiel:1942:BRB

[Eze42] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 37(218):288–289, June 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279228.

Ezekiel:1943:BRB

[Eze43] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 38(224):484, December 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279357.

Fabricant:1940:BRB

[Fab40] Solomon Fabricant. Book review: [untitled]. Journal of the American Statistical Association, 35(211):579–581, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279296.

Fabricant:1941:BRBa

[Fab41a] Solomon Fabricant. Book review: [untitled]. Journal of the American Statistical Association, 36(215):445–446, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279634.

Fabricant:1941:BRBb

[Fab41b] Solomon Fabricant. Book review: [untitled]. Journal of the American Statistical Association, 36(216):562–563, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278982.

Fairchild:1942:BRB

[Fai42] Mildred Fairchild. Book review: [untitled]. Journal of the American Statistical Association, 37(219):395–396, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279011.

Falcao:1940:GBG

[Fal40] Ildefonso Falcao. Greetings of the Brazilian Geographical and Statistical Institute. Journal of the American Statistical Association, 35(209):220–223, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279465.

Feller:1947:BRB

[Fel47] W. Feller. Book review: [untitled]. Journal of the American Statistical Association, 42(239):467–468, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280552.

Fertig:1945:BRB

[Fer45] John W. Fertig. Book review: [untitled]. Journal of the American Statistical Association, 40(231):386–388, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965639.

Ferger:1946:HNP

[Fer46] Wirth F. Ferger. Historical note on the purchasing power concept, and index numbers. Journal of the American Statistical Association, 41(233):53–57, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280155.

Ferber:1947:BRB

[Fer47] Robert Ferber. Book review: [untitled]. Journal of the American Statistical Association, 42(237):171–173, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280197.

Freeman:1948:BRB

[FG48] H. A. Freeman and Eugene L. Grant. Book review: [untitled]. Journal of the American Statistical Association, 43(242):353–355, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280389.

Fertig:1949:BRB

[FH49a] John W. Fertig and A. Bradford Hill. Book review: [untitled]. Journal of the American Statistical Association, 44(245):148–150, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280366.

Freedman:1949:UMD

[FH49b] Ronald Freedman and Amos H. Hawley. Unemployment and migration in the Depression (1930–1935). Journal of the American Statistical Association, 44(246):260–272, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280598.

Fiedler:1943:BRB

[Fie43] Reginald H. Fiedler. Book review: [untitled]. Journal of the American Statistical Association, 38(224):486–487, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279359.

Finney:1948:MEI

[Fin48] D. J. Finney. Main effects and interactions. Journal of the American Statistical Association, 43(244):566–571, December 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280709.

Fisher:1941:BRB

[Fis41] Ernest M. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 36(215):452–454, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279640.

Fisher:1942:BRB

[Fis42] Ernest M. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 37(218):300–301, June 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279239.

Fisher:1943:NDB

[Fis43] R. A. Fisher. Note on Dr. Berkson’s criticism of tests of significance. Journal of the American Statistical Association, 38(221):103– 104, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279319.

Fischer:1946:BRB

[Fis46] Carl H. Fischer. Book review: [untitled]. Journal of the American Statistical Association, 41(236):591–593, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280579.

Ferris:1947:BRB

[FJ47] Charles D. Ferris and N. Lloyd Johnson. Book review: [untitled]. Journal of the American Statistical Association, 42(239):482–486, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280559.

Florence:1941:BRB

[Flo41] P. Sargant Florence. Book review: [untitled]. Journal of the American Statistical Association, 36(214):303–304, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279754.

Frankel:1948:BRB

[FM48] Lester R. Frankel and Philip J. McCarthy. Book review: [untitled]. Journal of the American Statistical Association, 43(241):157–161, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280081.

Fertig:1947:BRB

[FN47] John W. Fertig and Horace W. Norton. Book review: [untitled]. Journal of the American Statistical Association, 42(239):477–479, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280556.

Folk:1942:SSM

[Fol42] Oliver Harold Folk. Selective Service’s medical statistics program. Journal of the American Statistical Association, 37(220):425–429, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279026.

Ford:1941:BRB

[For41] Robert S. Ford. Book review: [untitled]. Journal of the American Statistical Association, 36(213):157–158, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278836.

Forsling:1944:BRB

[For44] C. L. Forsling. Book review: [untitled]. Journal of the American Statistical Association, 39(225):122–123, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280335.

Foster:1944:BRBb

[Fos44a] Carol H. Foster. Book review: Tea under International Regulation, by V. D. Wickizer. Journal of the American Statistical Association, 39(227):402–403, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280055.

Foster:1944:BRBa

[Fos44b] Carol H. Foster. Book review: [untitled]. Journal of the American Statistical Association, 39(225):117–118, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280332.

Fischer:1940:SFR

[FP40] Carl H. Fischer and Henry H. Pixley. Specific fertility rates for Detroit. Journal of the American Statistical Association, 35 (212):653–661, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279254.

Fraenkel:1940:CHM

[Fra40] Marta Fraenkel. Classification of hospital morbidity. Journal of the American Statistical Association, 35(211):489–497, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279275.

Franklin:1941:BRB

[Fra41] Harry L. Franklin. Book review: [untitled]. Journal of the American Statistical Association, 36(215):442, September 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279631.

Fraenkel:1942:IHM

[Fra42] Marta Fraenkel. The importance of hospital morbidity data for the community. Journal of the American Statistical Association, 37(218):213–219, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279215.

Freeman:1945:BRB

[Fre45] Harold A. Freeman. Book review: [untitled]. Journal of the American Statistical Association, 40(231):386, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965638.

Freeman:1947:BRBb

[Fre47a] H. A. Freeman. Book review: Industrial Statistics. Journal of the American Statistical Association, 42(238):348–352, June 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280672.

Freeman:1947:BRBa

[Fre47b] H. A. Freeman. Book review: [untitled]. Journal of the American Statistical Association, 42(238):330–331, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280665.

Freeman:1949:BRB

[Fre49] H. A. Freeman. Book review: [untitled]. Journal of the American Statistical Association, 44(248):574–575, December 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279913.

Fritz:1940:BRBa

[Fri40a] Martin F. Fritz. Book review: Statistics Applied to Education and Psychology, by Clarence T. Gray; David F. Votaw. Journal of the American Statistical Association, 35(209):184–185, March 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279668.

Fritz:1940:BRBb

[Fri40b] Wilbert G. Fritz. Book review: Statistical Testing of Business-Cycle Theories, Volume II: Business Cycles in the United States of America 1919–1932, by J. Tinbergen. Journal of the American Statistical Association, 35(209):186–187, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279671.

Fritz:1941:BRB

[Fri41] Martin F. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 36(214):307–308, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279758.

Fritz:1942:BRB

[Fri42] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 37(219):410, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279023.

Fritz:1943:BRB

[Fri43] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 38(223):375–376, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279798.

Fritz:1944:BRBa

[Fri44a] Wilbert G. Fritz. Book review: The Mining Industries, 1899–1939: A Study of Output, Employment and Productivity, by Harold Barger; Sam H. Schurr. Journal of the American Statistical Association, 39(227):399–401, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280053.

Fritz:1944:BRBb

[Fri44b] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 39(228):528–529, December 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280637.

Frisch:1947:TIF

[Fri47] Ragnar Frisch. Tribute to irving Fisher. Journal of the American Statistical Association, 42(237):2–4, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280177.

Frankel:1942:SSU

[FS42] Lester R. Frankel and J. Stevens Stock. On the sample survey of unemployment. Journal of the American Statistical Association, 37(217):77–80, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279432.

Festinger:1948:BRB

[FS48a] Leon Festinger and Frederick F. Stephan. Book review: [untitled]. Journal of the American Statistical Association, 43(242):335–339, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280382.

Friedman:1948:BRB

[FS48b] Milton Friedman and Max Sasuly. Book review: [untitled]. Journal of the American Statistical Association, 43(241):139–145, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280077.

Funkhouser:1941:CN

[Fun41a] R. L. Funkhouser. Committee on nominations. Journal of the American Statistical Association, 36(215):435, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279625.

Funkhouser:1941:MAS

[Fun41b] Richard L. Funkhouser. Membership of the American Statistical Association on its hundredth anniversary. Journal of the American Statistical Association, 36(215):329–342, September 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279607.

Furlan:1946:HSU

[Fur46] Luigi Vladimir Furlan. Das Harmoniegesetz der Statistik: Eine Untersuchung über die metrische Interdependenz der sozialen Erscheinungen. (German) [The Law of Harmony in statistics: an investigation of the metrical interdependence of social phenomena], volume 2 of Beihefte zum Assekuranzjahrbuch. Verlag für Recht und Gesellschaft A.-G., Basel, Switzerland, 1946. xii + 504 pp. LCCN HA29 .F85. See [Gei48].

Fels:1940:ICR

[FW40] Bruno Fels and P. K. Whelpton. An industrial classification for reports from individuals. Journal of the American Statistical Association, 35(209):74–85, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279657.

Finney:1946:BRB

[FY46] D. J. Finney and W. J. Youden. Book review: [untitled]. Journal of the American Statistical Association, 41(234):262–266, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280502.

Gage:1941:BRB

[Gag41] Robert Gage. Book review: [untitled]. Journal of the American Statistical Association, 36(215):438, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279628.

Gage:1943:CTR

[Gag43] Robert Gage. Contents of Tippett’s “Random Sampling Numbers”. Journal of the American Statistical Association, 38(222):223–227, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279542.

Garfield:1944:MIP

[Gar44] Frank R. Garfield. Measurement of industrial production since 1939. Journal of the American Statistical Association, 39(228):439–454, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280621.

Garfield:1946:MFC

[Gar46] Frank R. Garfield. Measuring and forecasting consumption. Journal of the American Statistical Association, 41(235):322–333, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280256.

Gause:1946:BRB

[Gau46] G. R. Gause. Book review: [untitled]. Journal of the American Statistical Association, 41(233):130–132, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280166.

Greb:1949:ASS

[GB49] Donald J. Greb and Julio N. Berrettoni. AOQL single sampling plans from a single chart and table. Journal of the American Statistical Association, 44(245):62–76, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280350.

Gee:1944:BRB

[Gee44] Wilson Gee. Book review: [untitled]. Journal of the American Statistical Association, 39(225):116–117, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280331.

Geiringer:1948:RVF

[Gei48] Hilda Geiringer. Review of L. V. Furlan’s book: Das Harmoniegesetz der Statistik: Eine Untersuchung über die metrische Interdependenz der sozialen Erscheinungen. (German) [The Law of Harmony in statistics: an investigation of the metrical interdependence of social phenomena]. Journal of the American Statistical Association, 43(242):325–328, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280379. See [Fur46].

Geiringer:1949:SMP

[Gei49] Hilda Geiringer. On some mathematical problems arising in the development of Mendelian genetics. Journal of the American Statistical Association, 44(248):526–547, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279904.

Gilboy:1940:BRB

[Gil40a] Elizabeth W. Gilboy. Book review: Consumer Expenditures in the United States, Estimates for 1935–36. Consumer Expenditures in the United States, Estimates for 1935–36. Journal of the American Statistical Association, 35(210):430–431, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278737.

Gillett:1940:PPS

[Gil40b] R. L. Gillett. Problems and progress in state statistics. Journal of the American Statistical Association, 35(210):335–340, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278722.

Gilbert:1941:BRB

[Gil41a] Donald W. Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 36(215):444–445, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279633.

Gilboy:1941:BRB

[Gil41b] Elizabeth W. Gilboy. Book review: [untitled]. Journal of the American Statistical Association, 36(214):304–305, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279755.

Gilbert:1942:MNI

[Gil42] Milton Gilbert. Measuring national income as affected by the war. Journal of the American Statistical Association, 37(218):186–198, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279213.

Gilbert:1943:BRB

[Gil43] Milton Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 38(224):473–475, December 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279348.

Gillim:1944:PMM

[Gil44] Marion Gillim. Physical measurements of Mount Holyoke College freshmen in 1918 and 1943. Journal of the American Statistical Association, 39(225):53–56, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280323.

Girshick:1940:BRB

[Gir40] M. A. Girshick. Book review: [untitled]. Journal of the American Statistical Association, 35(211):559–561, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279285.

Girshick:1942:ATL

[Gir42a] M. A. Girshick. The application of the theory of linear hypotheses to the coefficient of elasticity of demand. Journal of the American Statistical Association, 37(218):233–237, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279218.

Girshick:1942:BRB

[Gir42b] M. A. Girshick. Book review: [untitled]. Journal of the American Statistical Association, 37(218):292–293, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279232.

Goffman:1946:BRB

[GK46] Casper Goffman and Lloyd A. Knowler. Book review: [untitled]. Journal of the American Statistical Association, 41(234):247–249, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280495.

Garvy:1947:NIH

[GL47a] George Garvy and Robert E. Lewis. New indexes of hourly and weekly earnings compiled by the Federal Reserve Bank of New York. Journal of the American Statistical Association, 42(238):256–270, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280655.

Guttman:1947:BRBb

[GL47b] Louis Guttman and D. N. Lawley. Book review: [untitled]. Journal of the American Statistical Association, 42(240):651–658, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280026.

Goldsmith:1941:BRB

[Gol41] Raymond W. Goldsmith. Book review: [untitled]. Journal of the American Statistical Association, 36(216):577–578, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278996.

Goldenweiser:1944:RP

[Gol44] E. A. Goldenweiser. Research and policy. Journal of the American Statistical Association, 39(225):1–9, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280317.

Goldstein:1946:WTA

[Gol46] Mortimer D. Goldstein. War-time aluminum statistics. Journal of the American Statistical Association, 41(233):34–52, March 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280154.

Goode:1945:BRB

[Goo45] Richard Goode. Book review: [untitled]. Journal of the American Statistical Association, 40(229):131, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280469.

Goodman:1947:SSC

[Goo47] Roe Goodman. Sampling for the 1947 survey of consumer finances. Journal of the American Statistical Association, 42(239):439–448, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280547.

Gordon:1944:IPD

[Gor44] R. A. Gordon. International programming of the distribution of resources: a symposium: III the combined raw materials board. Journal of the American Statistical Association, 39(227):291–296, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280034.

Goulden:1941:BRB

[Gou41] C. H. Goulden. Book review: [untitled]. Journal of the American Statistical Association, 36(214):313, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279764.

Gould:1944:BRB

[Gou44] Joseph S. Gould. Book review: [untitled]. Journal of the American Statistical Association, 39(225):118–119, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280333.

Goulden:1946:BRB

[Gou46] C. H. Goulden. Book review: [untitled]. Journal of the American Statistical Association, 41(233):132–133, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280167.

Graham:1941:BRB

[Gra41a] George A. Graham. Book review: [untitled]. Journal of the American Statistical Association, 36(216):549–550, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278970.

Grant:1941:BRB

[Gra41b] Eugene L. Grant. Book review: [untitled]. Journal of the American Statistical Association, 36(216):574–575, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278993.

Gray:1942:BRB

[Gra42] E. R. Gray. Book review: [untitled]. Journal of the American Statistical Association, 37(218):302–303, June 1942. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279241.

Grauman:1944:BRB

[Gra44] Jacob Grauman. Book review: [untitled]. Journal of the American Statistical Association, 39(228):539–541, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280645.

Grabill:1945:ALT

[Gra45a] Wilson H. Grabill. Attrition life tables for the single population. Journal of the American Statistical Association, 40(231):364–375, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965630.

Grant:1945:BRB

[Gra45b] Eugene L. Grant. Book review: [untitled]. Journal of the American Statistical Association, 40(231):377–378, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965632.

Grant:1946:BRB

[Gra46] David A. Grant. Book review: [untitled]. Journal of the American Statistical Association, 41(235):397–398, September 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280264.

Greenway:1941:BRB

[Gre41] Harold F. Greenway. Book review: [untitled]. Journal of the American Statistical Association, 36(213):145–146, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278824.

Greville:1946:BRB

[Gre46] Thomas N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 41(233):133, March 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280168.

Greville:1947:BRB

[Gre47] T. N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 42(239):479–481, September 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280557.

Greville:1948:BRB

[Gre48a] T. N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 43(241):161–163, March 1948. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280082.

Greville:1948:RDG

[Gre48b] Thomas N. E. Greville. Recent developments in graduation and interpolation. Journal of the American Statistical Association, 43 (243):428–441, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280279.

Greville:1949:BRB

[Gre49] T. N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 44(246):311–312, June 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280604.

Grubbs:1948:CEP

[Gru48a] Frank E. Grubbs. Corrections: On estimating precision of measuring instruments and products variability. Journal of the American Statistical Association, 43(244):564, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280707.

Grubbs:1948:EPM

[Gru48b] Frank E. Grubbs. On estimating precision of measuring instruments and product variability. Journal of the American Statistical Association, 43(242):243–264, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280371.

Galbraith:1941:BRI

[GT41] Virginia L. Galbraith and Dorothy S. Thomas. Birth rates and the interwar business cycles. Journal of the American Statistical Association, 36(216):465–476, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278953.

Guertin:1940:MTF

[Gue40] Alfred N. Guertin. Mortality tables — financial measuring rods in life insurance. Journal of the American Statistical Association, 35(209): 55–64, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279655.

Gumbel:1947:BRB

[Gum47] E. J. Gumbel. Book review: [untitled]. Journal of the American Statistical Association, 42(239):473–475, September 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280554.

Guttman:1945:BRB

[Gut45] Louis Guttman. Book review: [untitled]. Journal of the American Statistical Association, 40(231):390–391, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965642.

Guttman:1947:BRBa

[Gut47] Louis Guttman. Book review: [untitled]. Journal of the American Statistical Association, 42(237):191–192, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280206.

Guttman:1946:BRB

[GW46] Louis Guttman and Helen M. Walker. Book review: [untitled]. Journal of the American Statistical Association, 41(236):602–606, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280583.

Grubbs:1947:BUE

[GW47] Frank E. Grubbs and Chalmers L. Weaver. The best unbiased estimate of population standard deviation based on group ranges. Journal of the American Statistical Association, 42(238):224–241, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280652.

Haavelmo:1947:MMM

[Haa47] Trygve Haavelmo. Methods of measuring the marginal propensity to consume. Journal of the American Statistical Association, 42(237):105–122, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280191.

Haberler:1940:BRB

[Hab40] Gottfried Haberler. Book review: [untitled]. Journal of the American Statistical Association, 35(211):575, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279293.

Haemer:1948:BRB

[Hae48] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 43(244):613–616, December 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280715.

Hafstad:1940:BTT

[Haf40] L. R. Hafstad. On the Bartels technique for time-series analysis, and its relation to the analysis of variance. Journal of the American Statistical Association, 35(210):347–361, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278724.

Hagood:1941:BRB

[Hag41] Margaret Jarman Hagood. Book review: [untitled]. Journal of the American Statistical Association, 36(214):309–310, June 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279761.

Hagen:1943:BRB

[Hag43a] Everett E. Hagen. Book review: [untitled]. Journal of the American Statistical Association, 38(223):368–369, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279792.

Hagood:1943:BRB

[Hag43b] Margaret Jarman Hagood. Book review: [untitled]. Journal of the American Statistical Association, 38(222):252–253, June 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279551.

Hagood:1949:BRB

[Hag49] Margaret Jarman Hagood. Book review: [untitled]. Journal of the American Statistical Association, 44(246):312–313, June 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280605.

Hansen:1940:BRB

[Han40] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 35(211):581–583, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279297.

Harmon:1941:BRB

[Har41] Gaius E. Harmon. Book review: [untitled]. Journal of the American Statistical Association, 36(214):312, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279763.

Harris:1942:BRB

[Har42a] Seymour E. Harris. Book review: [untitled]. Journal of the American Statistical Association, 37(217):161–164, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279453.

Hartkemeier:1942:BRB

[Har42b] Harry Pelle Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 37(220):554–555, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279049.

Hartshorne:1942:BRB

[Har42c] E. Y. Hartshorne. Book review: [untitled]. Journal of the American Statistical Association, 37(217):153–154, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279446.

Hartkemeier:1945:BRB

[Har45] Harry Pelle Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 40(231):388–389, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965640.

Hartwell:1947:IRL

[Har47] John M. Hartwell, Jr. Interest rates by loan size and geographical region. Journal of the American Statistical Association, 42(238):242– 245, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280653.

Hartkemeier:1948:BRB

[Har48] Harry Pelle Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 43(241):137–139, March 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280075.

Hastings:1943:BRB

[Has43] Constance Hastings. Book review: [untitled]. Journal of the American Statistical Association, 38(221):132–133, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279331.

Hauser:1941:USC

[Hau41] Philip M. Hauser. The use of sampling in the census. Journal of the American Statistical Association, 36(215):369–375, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279611.

Hauser:1942:PAS

[Hau42] Philip M. Hauser. Proposed annual sample census of population. Journal of the American Statistical Association, 37(217):81–88, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279433.

Hauser:1947:SPC

[Hau47] Philip M. Hauser. The statistical program of the Census Bureau. Journal of the American Statistical Association, 42(237):24–30, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280184.

Hawkins:1945:BRB

[Haw45] E. R. Hawkins. Book review: [untitled]. Journal of the American Statistical Association, 40(231):404–407, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965650.

Hayford:1941:INC

[Hay41a] F. Leslie Hayford. An inquiry into the nature and causes of statisticians. Journal of the American Statistical Association, 36(213):1–10, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278807.

Hayford:1941:RPF

[Hay41b] F. Leslie Hayford. Roland Post Falkner, 1866–1940. Journal of the American Statistical Association, 36(216):543–545, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278965.

Hagood:1945:CIB

[HB45] Margaret Jarman Hagood and Eleanor H. Bernert. Component indexes as a basis for stratification in sampling. Journal of the American Statistical Association, 40(231):330–341, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965627.

Hansen:1943:SCA

[HD43] Morris H. Hansen and W. Edwards Deming. On some census aids to sampling. Journal of the American Statistical Association, 38 (223):353–357, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279787.

Hedges:1941:BRB

[Hed41] J. Edward Hedges. Book review: [untitled]. Journal of the American Statistical Association, 36(216):564–565, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278984.

Hedges:1945:BRB

[Hed45] J. Edward Hedges. Book review: [untitled]. Journal of the American Statistical Association, 40(229):132–133, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280471.

Hendricks:1944:REG

[Hen44] Walter A. Hendricks. The relative efficiencies of groups of farms as sampling units. Journal of the American Statistical Association, 39(227):366–376, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280042.

Hendricks:1947:FEL

[Hen47] Walter A. Hendricks. Farm employment levels in relation to supply and demand as per cent of normal. Journal of the American Statistical Association, 42(238):271–281, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280656.

Hersey:1941:SUC

[Her41] Arthur B. Hersey. Sources and uses of corporation funds. Journal of the American Statistical Association, 36(214):265–274, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279747.

Hansen:1942:REV

[HH42] Morris H. Hansen and William N. Hurwitz. Relative efficiencies of various sampling units in population inquiries. Journal of the American Statistical Association, 37(217):89–94, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279434.

Hamilton:1944:USR

[HH44] C. Horace Hamilton and F. M. Henderson. Use of the survival rate method in measuring net migration. Journal of the American Statistical Association, 39(226):197–206, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279949.

Hansen:1946:PNR

[HH46] Morris H. Hansen and William N. Hurwitz. The problem of nonresponse in sample surveys. Journal of the American Statistical Association, 41(236):517–529, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280572.

Hansen:1946:PMS

[HHG46] Morris H. Hansen, William N. Hurwitz, and Margaret Gurney. Problems and methods of the sample survey of business. Journal of the American Statistical Association, 41(234):173–189, June 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280485.

Harris:1948:DSS

[HHM48] Marilyn Harris, D. G. Horvitz, and A. M. Mood. On the determination of sample sizes in designing experiments. Journal of the American Statistical Association, 43(243):391–402, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280276.

Hinrichs:1940:ECM

[Hin40] A. F. Hinrichs. Effects of the 25-cent minimum wage on employment in the seamless hosiery industry. Journal of the American Statistical Association, 35(209):13–23, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279651.

Hinrichs:1943:CPN

[Hin43] A. F. Hinrichs. Charles P. Neill, 1865–1942. Journal of the American Statistical Association, 38(221):110–111, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279322.

Hinrichs:1947:DLM

[Hin47] A. F. Hinrichs. Developments in labor market and price statistics in 1945 and 1946. Journal of the American Statistical Association, 42(237):31–37, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280185.

Hirschman:1943:MDF

[Hir43] Albert O. Hirschman. On measures of dispersion for a finite distribution. Journal of the American Statistical Association, 38 (223):346–352, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279786.

Hansen:1946:BRB

[HL46] Morris H. Hansen and John Lintner. Book review: [untitled]. Journal of the American Statistical Association, 41(236):612–618, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280587.

Hartkemeier:1942:ODP

[HM42] Harry Pelle Hartkemeier and Herman E. Miller. Obtaining differences from punched cards. Journal of the American Statistical Association, 37(218):285–287, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279227.

Hurwitz:1946:MAS

[HM46] Abner Hurwitz and Floyd C. Mann. The membership of the American Statistical Association — an analysis. Journal of the American Statistical Association, 41(234):155–170, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280483.

Hoel:1947:CFF

[Hoe47] Paul G. Hoel. On the choice of forecasting formulas. Journal of the American Statistical Association, 42(240):605–611, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280017.

Hohaus:1940:APS

[Hoh40] Reinhard A. Hohaus. Actuarial problems in social insurance. Journal of the American Statistical Association, 35(209):37–46, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279653.

Hohman:1941:BRB

[Hoh41] Elmo Paul Hohman. Book review: [untitled]. Journal of the American Statistical Association, 36(213):150–151, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278829.

Holmes:1943:SSU

[Hol43] Irvin Holmes. Some sampling uses of data from the census of agriculture. Journal of the American Statistical Association, 38(221):78–86, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279315.

Hoover:1940:BRB

[Hoo40] Edgar M. Hoover, Jr. Book review: [untitled]. Journal of the American Statistical Association, 35(211):573–575, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279292.

Hoover:1948:BRB

[Hoo48] Edgar M. Hoover. Book review: [untitled]. Journal of the American Statistical Association, 43(244):601–603, December 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280712.

Horton:1949:BRB

[Hor49] H. Burke Horton. Book review: [untitled]. Journal of the American Statistical Association, 44(246):331–332, June 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280616.

Hotelling:1945:BRB

[Hot45] Susanna P. Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 40(230):267–268, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280144.

Hovde:1943:BRB

[Hov43] Howard T. Hovde. Book review: [untitled]. Journal of the American Statistical Association, 38(223):376–377, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279799.

Hoyle:1940:PCS

[Hoy40] Harry Hoyle. The Pennsylvania Committee on State Statistics. Journal of the American Statistical Association, 35(212):662–666, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279255.

Hoyt:1941:BRB

[Hoy41] Homer Hoyt. Book review: [untitled]. Journal of the American Statistical Association, 36(216):552–553, December 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278973.

Hutchinson:1947:ERA

[HR47] E. P. Hutchinson and Ernest Rubin. Estimating the resident alien population of the United States. Journal of the American Statistical Association, 42(239):385–400, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280542.

Hurwicz:1946:BRB

[Hur46] Leonid Hurwicz. Book review: [untitled]. Journal of the American Statistical Association, 41(234):254–255, June 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280499.

Hurwicz:1947:BRB

[Hur47] Leonid Hurwicz. Book review: [untitled]. Journal of the American Statistical Association, 42(239):461–467, September 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280551.

Hutchinson:1942:BRB

[Hut42] E. P. Hutchinson. Book review: [untitled]. Journal of the American Statistical Association, 37(220):556–557, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279051.

Heck:1941:BRB

[HW41] Harold J. Heck and Alexander Wall. Book review: [untitled]. Journal of the American Statistical Association, 36(213):158–160, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278837.

Hoel:1947:BRB

[HW47] Paul G. Hoel and J. Wolfowitz. Book review: [untitled]. Journal of the American Statistical Association, 42(237):174–179, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280199.

Hyde:1941:BRB

[Hyd41] Dorsey W. Hyde, Jr. Book review: [untitled]. Journal of the American Statistical Association, 36(213):149–150, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278828.

Ise:1942:BRB

[Ise42] John Ise. Book review: [untitled]. Journal of the American Statistical Association, 37(218):307–308, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279245.

Jacobson:1947:RPT

[Jac47] Paul H. Jacobson. The relative power of three statistics for small sample destructive tests. Journal of the American Statistical Association, 42(240):575–584, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280013.

James:1940:BRB

[Jam40] Clifford L. James. Book review: Normal and Abnormal International Capital Transfers, by Marco Fanno. Journal of the American Statistical Association, 35(210):433–434, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278739.

James:1941:BRB

[Jam41] Clifford L. James. Book review: [untitled]. Journal of the American Statistical Association, 36(214):317–318, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279769.

Jasny:1941:BRB

[Jas41] N. Jasny. Book review: [untitled]. Journal of the American Statistical Association, 36(216):554–555, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278975.

Jasny:1942:BRB

[Jas42] N. Jasny. Book review: [untitled]. Journal of the American Statistical Association, 37(218):290–291, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279230.

Jessen:1947:PSG

[JBKD47] Raymond J. Jessen, Richard H. Blythe, Jr., Oscar Kempthorne, and W. Edwards Deming. On a population sample for Greece. Journal of the American Statistical Association, 42(239):357–384, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280541.

Jensen:1940:BRB

[Jen40] William B. Jensen. Book review: Credit Unions in Massachusetts, by Joseph L. Snider. Journal of the American Statistical Association, 35(209):204–205, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279682.

Jessen:1945:MSA

[Jes45] R. J. Jessen. The master sample of agriculture: II, design. Journal of the American Statistical Association, 40(229):46–56, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280453.

Jessen:1949:SIF

[Jes49] Raymond J. Jessen. Some inadequacies of the Federal Censuses of Agriculture. Journal of the American Statistical Association, 44(246):279–292, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280600.

Johnson:1940:BRB

[Joh40] Sherman E. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 35(211):572–573, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279291.

Johnson:1941:FRB

[Joh41a] Norris O. Johnson. Federal Reserve Bank of New York indexes of production and trade. Journal of the American Statistical Association, 36(215):423–425, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279620.

Johnson:1941:BRB

[Joh41b] Palmer O. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 36(213):153–154, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278832.

Johnson:1945:BRB

[Joh45] E. C. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 40(229):136–137, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280475.

Jones:1941:UGM

[Jon41] Howard L. Jones. The use of grouped measurements. Journal of the American Statistical Association, 36(216):525–529, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278961.

Jones:1943:FPT

[Jon43] Howard L. Jones. Fitting polynomial trends to seasonal data by the method of least squares. Journal of the American Statistical Association, 38(224):453–465, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279345.

Jones:1946:LRF

[Jon46] Howard L. Jones. Linear regression functions with neglected variables. Journal of the American Statistical Association, 41(235): 356–369, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280259.

Joy:1941:MUS

[Joy41] Aryness Joy. The meaning of unemployment statistics. Journal of the American Statistical Association, 36(214):167–174, June 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279735.

Joy:1942:BRB

[Joy42] Aryness Joy. Book review: [untitled]. Journal of the American Statistical Association, 37(218):304–305, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279242.

Juliber:1941:RBS

[Jul41] Gustave S. Juliber. Relation between seasonal amplitudes and the level of production — an application to the production of steel ingots. Journal of the American Statistical Association, 36(216):485–492, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278955.

Jaffe:1945:IMF

[JW45] A. J. Jaffe and Seymour L. Wolfbein. Internal migration and full employment in the U.S. Journal of the American Statistical Association, 40(231):351–363, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965629.

Juran:1948:BRB

[JW48] Joseph M. Juran and Mason E. Wescott. Book review: [untitled]. Journal of the American Statistical Association, 43(242):343–346, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280385.

Kac:1948:BRB

[Kac48] M. Kac. Book review: [untitled]. Journal of the American Statistical Association, 43(243):496–497, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280286.

Kaplan:1942:MAE

[Kap42] Mortimer Kaplan. A method of analyzing the elements of foreclosure risk. Journal of the American Statistical Association, 37(218):247– 255, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279221.

Kaplansky:1945:CEC

[Kap45] Irving Kaplansky. A common error concerning kurtosis. Journal of the American Statistical Association, 40(230):259, June 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280139.

Katona:1947:CPD

[Kat47] George Katona. Contribution of psychological data to economic analysis. Journal of the American Statistical Association, 42(239): 449–459, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280548.

Karsten:1943:RC

[KB43] Karl Karsten and Edith Brooks. “retro” charts. Journal of the American Statistical Association, 38(223):302–310, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279780.

Knudsen:1947:UAT

[KC47] Lila F. Knudsen and Jack M. Curtis. The use of the angular transformation in biological assays. Journal of the American Statistical Association, 42(238):282–296, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280657.

Keim:1942:PW

[Kei42] Walter G. Keim. Prices and wages. Journal of the American Statistical Association, 37(219):377–382, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279006.

Keith:1945:BRB

[Kei45] E. Gordon Keith. Book review: [untitled]. Journal of the American Statistical Association, 40(229):131–132, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280470.

Kelley:1943:EPS

[Kel43] Truman L. Kelley. The evidence for periodicity in short time series. Journal of the American Statistical Association, 38(223):319–326, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279782.

Kellogg:1945:BRB

[Kel45a] Lester S. Kellogg. Book review: Opportunities for Statistical Workers, by Donald E. V. Henderson. Journal of the American Statistical Association, 40(232):530–531, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280224.

Kellogg:1945:EN

[Kel45b] Lester S. Kellogg. [editorial note]. Journal of the American Statistical Association, 40(231):??, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965623.

Kellogg:1945:MD

[Kel45c] Lester S. Kellogg. The membership directory. Journal of the American Statistical Association, 40(232):553–554, December 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279997.

Kempthorne:1948:BRB

[Kem48] O. Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 43(243):489–492, September 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280283.

Kenney:1949:BRB

[Ken49] J. F. Kenney. Book review: [untitled]. Journal of the American Statistical Association, 44(245):147–148, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280365.

Kimelman:1944:SPS

[Kim44] Juan Kimelman. Some problems of statistical education and teaching. Journal of the American Statistical Association, 39(226):229– 235, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279952.

King:1940:BRB

[Kin40] Willford I. King. Book review: Capitalism the Creator, the Economic Foundations of Modern Industrial Society, by Carl Snyder. Journal of the American Statistical Association, 35(210):440–442, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278746.

King:1942:BRB

[Kin42] Harold J. King. Book review: [untitled]. Journal of the American Statistical Association, 37(217):164–165, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279454.

King:1945:MSA

[Kin45] A. J. King. The master sample of agriculture: I, development and use. Journal of the American Statistical Association, 40(229):38–45, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280452.

Kirkpatrick:1943:BRB

[Kir43] Forrest H. Kirkpatrick. Book review: [untitled]. Journal of the American Statistical Association, 38(223):377–379, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279800.

Kirchen:1948:LSV

[Kir48] C. J. Kirchen. Levels of significance for variance ratio of two samples of equal size. Journal of the American Statistical Association, 43(244):564–565, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280708.

Kiser:1943:BRB

[Kis43] Clyde V. Kiser. Book review: [untitled]. Journal of the American Statistical Association, 38(222):253–254, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279552.

Kish:1949:POR

[Kis49] Leslie Kish. A procedure for objective respondent selection within the household. Journal of the American Statistical Association, 44 (247):380–387, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280236.

Klein:1944:BRB

[Kle44] Henry Klein. Book review: Costs of Dental Care for Adults under Specific Clinical Conditions, by Dorothy F. Beck. Journal of the American Statistical Association, 39(227):405–406, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280057.

Knowler:1946:BRB

[Kno46] Lloyd A. Knowler. Book review: [untitled]. Journal of the American Statistical Association, 41(235):402–403, September 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280266.

Knudsen:1940:IS

[Knu40] Lila F. Knudsen. Interdependence in a series. Journal of the American Statistical Association, 35(211):507–514, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279277.

Knudsen:1942:PCT

[Knu42] Lila F. Knudsen. A punched card technique to obtain coefficients of orthogonal polynomials. Journal of the American Statistical Association, 37(220):496–506, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279033.

Knudsen:1943:MDS

[Knu43] Lila F. Knudsen. A method for determining the significance of a shortage. Journal of the American Statistical Association, 38 (224):466–470, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279346.

Knudsen:1949:BRB

[Knu49] Lila F. Knudsen. Book review: [untitled]. Journal of the American Statistical Association, 44(247):448–450, September 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280243.

Kolesnikoff:1940:SCI

[Kol40] Vladimir S. Kolesnikoff. Standard classification of industries in the United States. Journal of the American Statistical Association, 35(209):65–73, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279656.

Kolesnikoff:1941:BRB

[Kol41] V. S. Kolesnikoff. Book review: [untitled]. Journal of the American Statistical Association, 36(215):446–448, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279635.

Kolesnikoff:1944:SCC

[Kol44] Vladimir S. Kolesnikoff. Standard commodity classification. Journal of the American Statistical Association, 39(225):42–52, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280322.

Koopmans:1941:BRB

[Koo41] Tjalling Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 36(214):319–320, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279771.

Koopmans:1945:SES

[Koo45] Tjalling Koopmans. Statistical estimation of simultaneous economic relations. Journal of the American Statistical Association, 40(232):448–466, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280215.

Koopmans:1947:BRB

[Koo47] Tjalling C. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 42(237):170–171, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280196.

Kormes:1940:SMA

[Kor40] Mark Kormes. Statistical methods as applied to casualty insurance. Journal of the American Statistical Association, 35(209):47–54, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279654.

Kozlik:1941:UPC

[Koz41] Adolf Kozlik. The use of per capita figures for demand curves. Journal of the American Statistical Association, 36(215):417–422, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279619.

Kreps:1940:BRB

[Kre40] Theodore J. Kreps. Book review: [untitled]. Journal of the American Statistical Association, 35(212):709–711, December 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279262.

Kuznets:1947:BRBa

[Kuz47a] G. M. Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 42(239):481–482, September 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280558.

Kuznets:1947:BRBb

[Kuz47b] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 42(240):639–641, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280022.

Kuznets:1949:WCM

[Kuz49] Simon Kuznets. Wesley Clair Mitchell, 1874–1948: An appreciation. Journal of the American Statistical Association, 44(245):126– 131, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280356.

Laderman:1946:AMD

[LA46] Jack Laderman and Milton Abramowitz. Application of machines to differencing of tables. Journal of the American Statistical Association, 41(234):233–237, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280490.

Laidler:1945:BRB

[Lai45] Harry L. Laidler. Book review: [untitled]. Journal of the American Statistical Association, 40(229):134–135, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280473.

Lamb:1947:PWS

[Lam47] G. A. Lamb. Post-War statistical planning in the Department of the Interior. Journal of the American Statistical Association, 42(237):37– 41, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280186.

Lang:1941:PSC

[Lan41] Richard O. Lang. Problems of statistical control–military aspects. Journal of the American Statistical Association, 36(213):11–26, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278808.

Lancaster:1942:MME

[Lan42] Otis E. Lancaster. Machine method for the extraction of cube root. Journal of the American Statistical Association, 37(217):112–115, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279437.

Lang:1944:PWP

[Lan44] Richard O. Lang. Post-War planning for British Official Statistics. Journal of the American Statistical Association, 39(227):271–280, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280031.

Lasken:1948:DDC

[Las48] Herman Lasken. Direct determination of compass settings for proportionate area pie-charts. Journal of the American Statistical Association, 43(241):94–95, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280069.

Latane:1942:SFD

[Lat42] Henry A. Latane. Seasonal factors determined by difference from average of adjacent months. Journal of the American Statistical Association, 37(220):517–522, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279035.

Latham:1946:OSW

[Lat46] Earl Latham. One statistical world. Journal of the American Statistical Association, 41(235):275–292, September 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280252.

Lavin:1946:IES

[Lav46] Marvin Lavin. Inspection efficiency and sampling inspection plans. Journal of the American Statistical Association, 41(236):432–438, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280567.

Lazarsfeld:1944:BRB

[Laz44] Paul F. Lazarsfeld. Book review: [untitled]. Journal of the American Statistical Association, 39(225):124–125, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280337.

Leavens:1944:BRB

[Lea44] Dickson H. Leavens. Book review: [untitled]. Journal of the American Statistical Association, 39(228):531–532, December 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280639.

Lebergott:1948:ENE

[Leb48] Stanley Lebergott. Earnings of nonfarm employees in the U.S., 1890– 1946. Journal of the American Statistical Association, 43(241):74–93, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280068.

LeClerg:1946:FRI

[LeC46] E. L. LeClerg. Forrest Rhinehart Immer, 1899–1946. Journal of the American Statistical Association, 41(234):242–243, June 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280493.

Lengyel:1944:IIB

[Len44] S. J. Lengyel. Insurance in the international balances of payments. Journal of the American Statistical Association, 39(228):428–438, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280620.

Leontief:1942:BRB

[Leo42] Wassily W. Leontief. Book review: [untitled]. Journal of the American Statistical Association, 37(217):149–150, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279442.

Leonard:1945:OSW

[Leo45] William R. Leonard. Opportunities for statistical work in the Federal Government. Journal of the American Statistical Association, 40(229):57–61, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280454.

Lester:1949:EMS

[Les49] A. M. Lester. The edge marking of statistical cards. Journal of the American Statistical Association, 44(246):293–294, June 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280601.

Levine:1942:IWL

[Lev42] Louis Levine. The impact of war on labor supply and labor utilization. Journal of the American Statistical Association, 37(217):54–62, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279429.

Levene:1949:BRB

[Lev49] Howard Levene. Book review: [untitled]. Journal of the American Statistical Association, 44(247):447–448, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280242.

Lewis:1940:BRB

[Lew40] Cleona Lewis. Book review: [untitled]. Journal of the American Statistical Association, 35(212):712–713, December 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279264.

Lewis:1941:NDS

[Lew41a] H. Gregg Lewis. The nature of the demand for steel. Journal of the American Statistical Association, 36(213):110–115, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278819.

Lewis:1941:DPE

[Lew41b] H. Gregg Lewis. On the distribution of the partial elasticity coefficient. Journal of the American Statistical Association, 36 (215):413–416, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279618.

Lew:1944:ISM

[Lew44] Edward A. Lew. Interpreting the statistics of medical examinations of selectees. Journal of the American Statistical Association, 39(227):345–356, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280040.

Lewis:1945:BRB

[Lew45] Edward E. Lewis. Book review: [untitled]. Journal of the American Statistical Association, 40(229):138–139, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280477.

Lew:1948:MSL

[Lew48] Edward A. Lew. Mortality statistics for life insurance underwritting. Journal of the American Statistical Association, 43(242):274– 289, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280373.

Li:1946:BRB

[Li46] Jerome C. R. Li. Book review: [untitled]. Journal of the American Statistical Association, 41(235):411–412, September 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280271.

Littauer:1946:BRB

[Lit46] Sebastian B. Littauer. Book review: [untitled]. Journal of the American Statistical Association, 41(233):111–112, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280160.

Littauer:1949:BRBb

[Lit49] Sebastian B. Littauer. Book review: [untitled]. Journal of the American Statistical Association, 44(248):577–579, December 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279915.

Lively:1945:BRB

[Liv45a] C. E. Lively. Book review: [untitled]. Journal of the American Statistical Association, 40(229):135–136, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280474.

Livingston:1945:CST

[Liv45b] J. A. Livingston. Charts should tell a story. Journal of the American Statistical Association, 40(231):342–350, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965628.

Livingston:1945:MPL

[Liv45c] S. Morris Livingston. The measurement of postwar labor supply and its capacity to produce. Journal of the American Statistical Association, 40(229):20–28, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280450.

Livingston:1947:BRB

[Liv47] S. Morris Livingston. Book review: [untitled]. Journal of the American Statistical Association, 42(240):641–643, December 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280023.

Lockley:1942:BRBa

[Loc42a] Lawrence C. Lockley. Book review: [untitled]. Journal of the American Statistical Association, 37(218):305–306, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279243.

Lockley:1942:BRBb

[Loc42b] Lawrence C. Lockley. Book review: [untitled]. Journal of the American Statistical Association, 37(220):555–556, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279050.

Longstreet:1942:BRB

[Lon42] Victor M. Longstreet. Book review: [untitled]. Journal of the American Statistical Association, 37(219):405–406, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279019.

Lorimer:1942:BRB

[Lor42] Frank Lorimer. Book review: [untitled]. Journal of the American Statistical Association, 37(217):155–157, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279448.

Lorimer:1943:BRB

[Lor43] Frank Lorimer. Book review: [untitled]. Journal of the American Statistical Association, 38(222):250–252, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279550.

Lotka:1943:SRS

[Lot43] Alfred J. Lotka. Some reflections — statistical and other — on a non-material universe. Journal of the American Statistical Association, 38(221):1–15, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279308.

Lotka:1944:GHV

[Lot44] Alfred J. Lotka. George H. Van Buren, 1874–1943. Journal of the American Statistical Association, 39(225):96, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280325.

Louden:1945:BRB

[Lou45] Mary V. Louden. Book review: [untitled]. Journal of the American Statistical Association, 40(230):274, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280148.

Love:1941:HNL

[Lov41] John W. Love. How a newspaperman likes statistics presented. Journal of the American Statistical Association, 36(214):233–239, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279743.

Lotka:1940:TBR

[LS40] Alfred J. Lotka and Mortimer Spiegelman. The trend of the birth rate by age of mother and order of birth. Journal of the American Statistical Association, 35(212):595–601, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279249.

Lukacs:1947:BRB

[LS47] Eugene Lukacs and Mortimer Spiegelman. Book review: [untitled]. Journal of the American Statistical Association, 42(238):335–340, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280668.

Levene:1948:BRB

[LS48] Howard Levene and Ronald W. Shephard. Book review: [untitled]. Journal of the American Statistical Association, 43(242):332–335, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280381.

Littauer:1949:BRBa

[LS49] Sebastian B. Littauer and Charles R. Scott, Jr. Book review: [untitled]. Journal of the American Statistical Association, 44(246):320– 322, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280610.

Littauer:1947:BRB

[LT47] Sebastian B. Littauer and L. H. C. Tippett. Book review: [untitled]. Journal of the American Statistical Association, 42(237):180– 184, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280201.

Lubin:1947:SEA

[Lub47a] Isador Lubin. Social and economic adjustments in a democratic world. Journal of the American Statistical Association, 42(237):11– 19, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280181.

Lubin:1947:TTP

[Lub47b] Isador Lubin. A tribute to two Past Presidents of the Association. Journal of the American Statistical Association, 42(237):1–2, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280176.

Lukacs:1948:BRB

[Luk48] Eugene Lukacs. Book review: [untitled]. Journal of the American Statistical Association, 43(241):135–136, March 1948. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280073.

Lumpkin:1944:BRB

[Lum44] Katharine Du Pre Lumpkin. Book review: The Movement of Factory Workers: A Study of a New England Industrial Community, 1937– 1939 and 1942, by Charles A. Myers; W. Rupert Maclaurin. Journal of the American Statistical Association, 39(227):397–398, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280051.

Lunden:1940:BRB

[Lun40] Walter A. Lunden. Book review: De Svenska Privatanstallda, en Sociologisk Studie, by Fritz Croner. Journal of the American Statistical Association, 35(210):444–446, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278749.

Lundberg:1941:BRB

[Lun41] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 36(215):440–442, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279630.

Lury:1946:ALS

[Lur46] D. B. De Lury. The analysis of Latin squares when some observations are missing. Journal of the American Statistical Association, 41(235):370–389, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280260.

Leighton:1947:BRB

[LW47] Walter Leighton and Charles P. Winsor. Book review: [untitled]. Journal of the American Statistical Association, 42(238):344–347, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280671.

MacLeod:1942:BRB

[Mac42] Alan MacLeod. Book review: [untitled]. Journal of the American Statistical Association, 37(220):550–551, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279046.

Madow:1940:NTD

[Mad40] William G. Madow. Note on tests of departure from normality. Journal of the American Statistical Association, 35(211):515–517, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279278. See

comments [BG41].

Madow:1946:SSR

[Mad46] Lillian H. Madow. Systematic sampling and its relation to other sampling designs. Journal of the American Statistical Association, 41(234):204–217, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280487.

Magnusson:1942:BRB

[Mag42] Leifur Magnusson. Book review: [untitled]. Journal of the American Statistical Association, 37(217):157–158, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279449.

Malenbaum:1941:BRB

[Mal41] Wilfred Malenbaum. Book review: [untitled]. Journal of the American Statistical Association, 36(215):455–456, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279642.

Manuele:1940:BRB

[Man40] Joseph Manuele. Book review: Statistical Method from the Viewpoint of Quality Control, by Walter A. Shewhart; W. Edwards Deming. Journal of the American Statistical Association, 35(210):426– 427, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278733.

Manuele:1942:BRB

[Man42] Joseph Manuele. Book review: [untitled]. Journal of the American Statistical Association, 37(220):559–560, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279054.

Mann:1945:BRB

[Man45] Henry B. Mann. Book review: [untitled]. Journal of the American Statistical Association, 40(231):396, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965646.

Marcuse:1945:ACC

[Mar45] Sophie Marcuse. An application of the control chart method to the testing and marketing of foods. Journal of the American Statistical Association, 40(230):214–222, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280127.

Marshall:1946:SAC

[Mar46] Herbert Marshall. Sedley Anthony Cudmore 1878–1945. Journal of the American Statistical Association, 41(233):75–76, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280158.

Martin:1949:BRB

[Mar49] Margaret Martin. Book review: [untitled]. Journal of the American Statistical Association, 44(245):146–147, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280364.

May:1940:BRB

[May40] George O. May. Book review: [untitled]. Journal of the American Statistical Association, 35(211):578–579, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279295.

Mayer:1942:BRB

[May42] Joseph Mayer. Book review: [untitled]. Journal of the American Statistical Association, 37(218):308–309, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279246.

May:1944:BRB

[May44] George O. May. Book review: Excess Profits Taxation, by Kenneth James Curran. Journal of the American Statistical Association, 39(227):391–392, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280046.

Merrell:1948:BRB

[MB48] Margaret Merrell and Joseph Berkson. Book review: [untitled]. Journal of the American Statistical Association, 43(241):145–153, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280078.

Mills:1943:AUB

[MBC^{+}43] Frederick C. Mills, E. Wight Bakke, Reavis Cox, Margaret G. Reid, Theodore W. Schultz, and Samuel S. Stratton. An appraisal of the U.S. Bureau of Labor Statistics Cost of Living Index. Journal of the American Statistical Association, 38(224):387–405, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279339.

McCormick:1943:AMF

[McC43] Thomas C. McCormick. An approach to the measurement of farm population pressure in Wisconsin. Journal of the American Statistical Association, 38(222):165–177, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279536.

McDiarmid:1940:BRB

[McD40] O. J. McDiarmid. Book review: Report of the Fifth Annual Research Conference on Economics and Statistics. Report of the Fifth Annual Research Conference on Economics and Statistics. Journal of the American Statistical Association, 35(209):187–189, March 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279672.

McDiarmid:1941:BRB

[McD41] O. J. McDiarmid. Book review: [untitled]. Journal of the American Statistical Association, 36(216):570–571, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278989.

McDiarmid:1942:BRB

[McD42] O. J. McDiarmid. Book review: [untitled]. Journal of the American Statistical Association, 37(218):306, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279244.

McDiarmid:1948:LCI

[McD48] Fergus J. McDiarmid. Life company investments and the capital markets. Journal of the American Statistical Association, 43(242):265– 273, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280372.

McGarry:1944:BRB

[McG44a] E. D. McGarry. Book review: [untitled]. Journal of the American Statistical Association, 39(228):534–535, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280643.

McGill:1944:SPS

[McG44b] Kenneth H. McGill. The statistical program of the selective service system. Journal of the American Statistical Association, 39(225):10– 20, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280318.

McIntyre:1941:BRB

[McI41a] Francis McIntyre. Book review: [untitled]. Journal of the American Statistical Association, 36(214):314–315, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279766.

McIntyre:1941:WPI

[McI41b] Francis McIntyre. Wholesale price indexes. Journal of the American Statistical Association, 36(214):185–190, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279737.

McIntyre:1942:BRB

[McI42] Francis McIntyre. Book review: [untitled]. Journal of the American Statistical Association, 37(219):400–401, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279015.

McPherson:1942:MOP

[McP42] J. C. McPherson. Mathematical operations with punched cards. Journal of the American Statistical Association, 37(218):275–281, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279225. Reprinted

in Annals of the History of Computing, 7(4), October 1985, 368–371.

McPeak:1945:PFM

[McP45] William McPeak. Problems of field management in Army opinion research. Journal of the American Statistical Association, 40(230):247– 248, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280136.

McVay:1947:SMA

[McV47] Francis E. McVay. Sampling methods applied to estimating numbers of commercial orchards in a commercial peach area. Journal of the American Statistical Association, 42(240):533–540, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280010.

Mautz:1943:PWL

[MD43] William H. Mautz and John D. Durand. Population and war labor supply. Journal of the American Statistical Association, 38(221):31– 42, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279310.

Mottley:1942:EFM

[ME42] C. McC. Mottley and Daniel R. Embody. The effect of the full Moon on trout fishing. Journal of the American Statistical Association, 37(217):41–47, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279427.

Meehan:1942:IDU

[Mee42] M. Joseph Meehan. Impact of defense upon industrial capacity and investment. Journal of the American Statistical Association, 37(217):

22–33, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279425.

Meehan:1945:IBS

[Mee45] M. Joseph Meehan. Industry and business statistics. Journal of the American Statistical Association, 40(230):229–231, June 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280129.

Mendershausen:1941:BRB

[Men41] Horst Mendershausen. Book review: [untitled]. Journal of the American Statistical Association, 36(213):154–155, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278833.

Merrell:1943:BRB

[Mer43] Margaret Merrell. Book review: [untitled]. Journal of the American Statistical Association, 38(221):133–134, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279332.

Merrell:1944:BRB

[Mer44] Margaret Merrell. Book review: Criminal Careers in Retrospect, by Sheldon Glueck; Eleanor Glueck. Journal of the American Statistical Association, 39(227):406–407, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280058.

Meyer:1949:BRB

[Mey49] Walter H. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 44(248):575–577, December 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279914.

Mighell:1944:BRB

[Mig44] Ronald L. Mighell. Book review: Food Rationing and Supply, 1943– 44 Food Rationing and Supply, 1943–44. Journal of the American Statistical Association, 39(227):403–405, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280056.

Miller:1943:BRB

[Mil43] J. P. Miller. Book review: [untitled]. Journal of the American Statistical Association, 38(221):128–129, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279327.

Milton:1945:BRB

[Mil45] Helen Slentz Milton. Book review: [untitled]. Journal of the American Statistical Association, 40(229):137–138, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280476.

Mills:1946:EPQ

[Mil46] Frederick C. Mills. Elasticity of physical quantities and flexibility of unit prices in the dimension of time. Journal of the American Statistical Association, 41(236):439–467, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280568.

Miller:1948:SPD

[Mil48] Morton D. Miller. Statistical problems in developing voluntary medical expense insurance plans. Journal of the American Statistical Association, 43(242):290–298, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280374.

Miner:1941:BRB

[Min41] John R. Miner. Book review: [untitled]. Journal of the American Statistical Association, 36(214):309, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279760.

Miner:1945:EF

[Min45] John R. Miner. Election of fellows. Journal of the American Statistical Association, 40(229):101, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280462.

Myers:1943:SWS

[MO43] Robert J. Myers and Harry Ober. Statistics for wage stabilization. Journal of the American Statistical Association, 38(224):425–437, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279342.

Modley:1948:BRB

[Mod48] Rudolf Modley. Book review: [untitled]. Journal of the American Statistical Association, 43(242):341–343, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280384.

Molina:1943:BRB

[Mol43] E. C. Molina. Book review: [untitled]. Journal of the American Statistical Association, 38(222):267–268, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279563.

Molina:1944:AF

[Mol44] Edward C. Molina. Arne Fisher, 1887–1944. Journal of the American Statistical Association, 39(226):251, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279958.

Montgomery:1940:BRB

[Mon40] Royal E. Montgomery. Book review: Labor in the United States, Basic Statistics for Social Security, by W. S. Woytinsky. Journal of the American Statistical Association, 35(210):446–447, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278750.

Moore:1944:IMP

[Moo44] Geoffrey H. Moore. Industrial materials production, World Wars I and II. Journal of the American Statistical Association, 39 (227):335–344, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280039.

Morehouse:1940:WCS

[Mor40a] E. W. Morehouse. The Wisconsin Committee on Statistics. Journal of the American Statistical Association, 35(211):483–488, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279274.

Morgenstern:1940:BRB

[Mor40b] Oskar Morgenstern. Book review: Business Cycles: A Theoretical, Historical, and Statistical Analysis of the Capitalist Process, by Joseph A. Schumpeter. Journal of the American Statistical Association, 35(210):423–424, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278731.

Morrison:1949:PDF

[Mor49] Nathan Morrison. By-product data and forecasting in unemployment insurance. Journal of the American Statistical Association, 44(247):397–405, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280238.

Mosak:1941:STI

[Mos41] Jacob L. Mosak. Some theoretical implications of the statistical analysis of demand and cost functions for steel. Journal of the American Statistical Association, 36(213):100–109, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278818.

Mosak:1945:BRB

[Mos45a] Jacob L. Mosak. Book review: [untitled]. Journal of the American Statistical Association, 40(229):123–125, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280464.

Mosteller:1945:BRB

[Mos45b] Frederick Mosteller. Book review: [untitled]. Journal of the American Statistical Association, 40(231):379–380, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965634.

Mosteller:1947:BRB

[Mos47] Frederick Mosteller. Book review: [untitled]. Journal of the American Statistical Association, 42(237):192–195, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280207.

Mosteller:1948:PD

[Mos48] Frederick Mosteller. On pooling data. Journal of the American Statistical Association, 43(242):231–242, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280369.

Moser:1949:USG

[Mos49a] C. A. Moser. The use of sampling in Great Britain. Journal of the American Statistical Association, 44(246):231–259, June 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280597.

Mosteller:1949:BRBa

[Mos49b] Frederick Mosteller. Book review: [untitled]. Journal of the American Statistical Association, 44(245):150–155, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280367.

Mosteller:1949:BRBb

[Mos49c] Frederick Mosteller. Book review: [untitled]. Journal of the American Statistical Association, 44(248):570–572, December 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279911.

Mottley:1949:BRB

[Mot49] Charles M. Mottley. Book review: [untitled]. Journal of the American Statistical Association, 44(247):456–457, September 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280248.

Mowbray:1941:OCA

[Mow41] A. H. Mowbray. Observation on correlation analysis. Journal of the American Statistical Association, 36(214):248–252, June 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279745.

McLaughlin:1943:BRBb

[MS43] Glenn E. McLaughlin and A. Smithies. Book review: [untitled]. Journal of the American Statistical Association, 38(223):367, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279791.

Mood:1946:BRB

[MS46] A. M. Mood and Milner B. Schaefer. Book review: [untitled]. Journal of the American Statistical Association, 41(233):128–130, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280165.

Morton:1946:BRB

[MT46] J. E. Morton and Gerhard Tintner. Book review: [untitled]. Journal of the American Statistical Association, 41(236):606–609, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280584.

Mosteller:1949:UUB

[MT49] Frederick Mosteller and John W. Tukey. The uses and usefulness of binomial probability paper. Journal of the American Statistical Association, 44(246):174–212, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280595.

Metropolis:1949:MCM

[MU49] Nicholas Metropolis and S. Ulam. The Monte Carlo method. Journal of the American Statistical Association, 44(247):335–341, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://links.jstor.org/sici?sici=0162-1459%
28194909%2944%3A247%3C335%3ATMCM%3E2.0.CO%3B2-3;   http:
//www.jstor.org/stable/2280232.

Mudgett:1946:BRB

[Mud46] Bruce D. Mudgett. Book review: [untitled]. Journal of the American Statistical Association, 41(235):395–397, September 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280263.

Muench:1948:BRB

[Mue48] Hugo Muench. Book review: [untitled]. Journal of the American Statistical Association, 43(241):153–155, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280079.

Murray:1945:BRB

[Mur45] William G. Murray. Book review: [untitled]. Journal of the American Statistical Association, 40(229):133–134, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280472.

Mushkin:1940:BRB

[Mus40] S. J. Mushkin. Book review: German Financial Policies 1932–1939, by Kenyon E. Poole. Journal of the American Statistical Association, 35(209):201–202, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279680.

Musgrave:1943:BRB

[Mus43] Richard A. Musgrave. Book review: [untitled]. Journal of the American Statistical Association, 38(224):476–477, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279350.

McLaughlin:1943:BRBa

[MW43a] Glenn E. McLaughlin and Georg Wolff. Book review: [untitled]. Journal of the American Statistical Association, 38(221):125–126, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279325.

Moore:1943:TSS

[MW43b] Geoffrey H. Moore and W. Allen Wallis. Time series significance tests based on signs of differences. Journal of the American Statistical Association, 38(222):153–164, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279535.

Myers:1941:DLS

[Mye41a] Howard B. Myers. Dynamics of labor supply. Journal of the American Statistical Association, 36(214):175–184, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279736.

Myers:1941:VSM

[Mye41b] Robert J. Myers. The validity and significance of male net reproduction rates. Journal of the American Statistical Association, 36(214):275–282, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279748.

Myers:1942:DML

[Mye42] Howard B. Myers. Defense migration and labor supply. Journal of the American Statistical Association, 37(217):69–76, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279431.

Myers:1943:CDR

[Mye43] Robert J. Myers. Comparison of demographic rates assumed by the national resources committee with actual experience. Journal of the American Statistical Association, 38(222):201–209, June 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279539.

Myers:1944:BRB

[Mye44] Robert J. Myers. Book review: [untitled]. Journal of the American Statistical Association, 39(226):263–264, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279968.

Myers:1945:BRB

[Mye45a] Robert J. Myers. Book review: [untitled]. Journal of the American Statistical Association, 40(231):384–385, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965637.

Myers:1945:WCU

[Mye45b] Robert J. Myers. Wartime changes in urban wage relationships. Journal of the American Statistical Association, 40(230):175–186, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280123.

Myers:1946:CCG

[Mye46] Howard B. Myers. Corrington Calhoun Gill, 1898–1946. Journal of the American Statistical Association, 41(235):393–394, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280262.

Myers:1947:CMM

[Mye47] R. J. Myers. A critique of “A Method of Making Actuarial Estimates for a Compulsory Health Insurance System”. Journal of the American Statistical Association, 42(237):123–127, March 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280192.

Myers:1949:BRB

[Mye49a] Robert Cobb Myers. Book review: [untitled]. Journal of the American Statistical Association, 44(246):315–317, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280608.

Myers:1949:BSU

[Mye49b] Robert J. Myers. Beneficiary statistics under the old-age and survivors insurance program and some possible demographic studies based on these data. Journal of the American Statistical Association, 44(247):388–396, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280237.

Nagelberg:1940:BRB

[Nag40] Morton S. Nagelberg. Book review: Graphic Presentation, by Willard C. Brinton. Journal of the American Statistical Association, 35(209): 185, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279669.

Nathan:1940:BRB

[Nat40] Otto Nathan. Book review: Die Schalthebel der Konjunktur, by Paul Binder. Journal of the American Statistical Association, 35(209):190– 191, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279674.

Neifeld:1941:BRB

[Nei41] M. R. Neifeld. Book review: [untitled]. Journal of the American Statistical Association, 36(213):160–161, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278838.

Nerlove:1940:BRB

[Ner40] S. H. Nerlove. Book review: [untitled]. Journal of the American Statistical Association, 35(211):564–568, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279288.

Nesti:1947:BRB

[Nes47] Anthony J. Nesti. Book review: [untitled]. Journal of the American Statistical Association, 42(240):644–651, December 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280025.

Newcomer:1942:BRB

[New42] Mabel Newcomer. Book review: [untitled]. Journal of the American Statistical Association, 37(220):553, December 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279048.

Neyman:1940:BRBa

[Ney40a] J. Neyman. Book review: [untitled]. Journal of the American Statistical Association, 35(211):558–559, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279284.

Neyman:1940:BRBb

[Ney40b] J. Neyman. Book review: [untitled]. Journal of the American Statistical Association, 35(211):562–564, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279287.

Neyman:1943:BRB

[Ney43] J. Neyman. Book review: [untitled]. Journal of the American Statistical Association, 38(221):134–135, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279333.

Nichols:1942:SWL

[Nic42] Charles K. Nichols. The statistical work of the league of nations in economic, financial and related fields. Journal of the American Statistical Association, 37(219):336–342, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279001.

Nichols:1943:BRB

[Nic43] Bervard Nichols. Book review: [untitled]. Journal of the American Statistical Association, 38(224):483, December 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279356.

Nickerson:1944:BRB

[Nic44] C. B. Nickerson. Book review: [untitled]. Journal of the American Statistical Association, 39(226):264–266, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279969.

Niessen:1945:SCP

[Nie45] A. M. Niessen. On the summation of certain progressions useful in time series analysis. Journal of the American Statistical Association, 40(229):98–100, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280461.

Niessen:1948:AEP

[Nie48] Abraham M. Niessen. Actuarial estimates for public sickness insurance plans. Journal of the American Statistical Association, 43(241): 61–73, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280067.

Noether:1949:CLN

[Noe49] Gottfried E. Noether. Confidence limits in the non-parametric case. Journal of the American Statistical Association, 44(245):89–100, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280352.

Norman:1940:BRB

[Nor40] Theodore Norman. Book review: Wheat Studies of the Food Research Institute, by Helen C. Farnsworth; Holbrook Working; Seasonal Aspects of the European Wheat Trade, by M. K. Bennett; Durum Wheats and their Utilization, by Carl L. Alsberg; World Wheat Survey and Outlook May 1939, by V. P. Timoshenko. Journal of the American Statistical Association, 35(209):196–197, March 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279678.

Nordin:1944:DSS

[Nor44] J. A. Nordin. Determining sample size. Journal of the American Statistical Association, 39(228):497–506, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280626.

Nortman:1945:BRB

[Nor45a] P. Bernard Nortman. Book review: [untitled]. Journal of the American Statistical Association, 40(229):141, March 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280479.

Norton:1945:CCS

[Nor45b] H. W. Norton. Calculation of chi-square for complex contingency tables. Journal of the American Statistical Association, 40(230):251– 258, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280138.

Nisselson:1948:BRB

[NP48] Harold Nisselson and Edgar Z. Palmer. Book review: [untitled]. Journal of the American Statistical Association, 43(241):155–157, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280080.

Nagel:1945:BRB

[NS45] Ernest Nagel and John H. Smith. Book review: Statistics, by L. H. C. Tippett. Journal of the American Statistical Association, 40(232):536–538, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280228.

Novick:1948:WPBa

[NS48a] David Novick and George A. Steiner. The war production Board’s statistical reporting experience. Journal of the American Statistical Association, 43(242):201–230, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280368.

Novick:1948:WPBb

[NS48b] David Novick and George A. Steiner. The war production Board’s statistical reporting experience, II and III. Journal of the American Statistical Association, 43(243):463–488, September 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280282.

Novick:1948:WPBc

[NS48c] David Novick and George A. Steiner. The war production Board’s statistical reporting experience: Part IV: Questionnaire units of measure and nomenclature. Journal of the American Statistical Association, 43(244):575–596, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280711.

Novick:1949:WPB

[NS49] David Novick and George A. Steiner. The War Production Board’s statistical reporting experience, V and VI. Journal of the American Statistical Association, 44(247):413–443, September 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280240.

Nugent:1940:BRB

[Nug40] Rolf Nugent. Book review: Personal Finance Comes of Age, by M. R. Neifeld. Journal of the American Statistical Association, 35(209):202– 204, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279681.

Orcutt:1949:SSM

[OC49] Guy H. Orcutt and Donald Cochrane. A sampling study of the merits of auto-regressive and reduced form transformations in regression analysis. Journal of the American Statistical Association, 44(247):356–372, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280234.

Ogburn:1940:ST

[Ogb40] William Fielding Ogburn. Statistical trends. Journal of the American Statistical Association, 35(209):252–260, March 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279470.

Ogle:1943:BRB

[Ogl43] Kenneth N. Ogle. Book review: [untitled]. Journal of the American Statistical Association, 38(222):268, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279564.

Olds:1949:TSQ

[OK49] Edwin G. Olds and Lloyd A. Knowler. Teaching statistical quality control for town and gown. Journal of the American Statistical Association, 44(246):213–230, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280596.

Olds:1946:BRB

[Old46] Edwin G. Olds. Book review: Publications Received. Journal of the American Statistical Association, 41(234):269, June 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280504.

Olds:1949:CBU

[Old49] Edward B. Olds. The city block as a unit for recording and analyzing urban data. Journal of the American Statistical Association, 44(248):485–500, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279901.

Olmstead:1942:BRB

[Olm42] P. S. Olmstead. Book review: A Significance Test for Time Series, by W. Allen Wallis; Geoffrey H. Moore. Journal of the American Statistical Association, 37(217):152–153, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279445.

Olmstead:1945:BRB

[Olm45] P. S. Olmstead. Book review: [untitled]. Journal of the American Statistical Association, 40(231):408–409, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965651.

Olds:1948:BRB

[OO48] Edwin G. Olds and Paul S. Olmstead. Book review: [untitled]. Journal of the American Statistical Association, 43(244):605–613, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280714.

Ore:1948:BRB

[Øre48] Øystein Øre. Book review: [untitled]. Journal of the American Statistical Association, 43(241):139, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280076.

Osborne:1942:SES

[Osb42] James G. Osborne. Sampling errors of systematic and random surveys of cover-type areas. Journal of the American Statistical Association, 37(218):256–264, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279222.

Ottman:1942:BRB

[Ott42] Frederick R. Ottman. Book review: [untitled]. Journal of the American Statistical Association, 37(217):150–151, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279443.

Palmer:1941:SCI

[Pal41a] Gladys L. Palmer. Some considerations involved in appraising the adequacy of occupational statistics. Journal of the American Statistical Association, 36(213):61–70, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278814.

Palyi:1941:BRB

[Pal41b] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 36(216):556–557, December 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278977.

Palyi:1942:BRB

[Pal42] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 37(219):403–405, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279018.

Palmer:1943:FVR

[Pal43a] Gladys L. Palmer. Factors in the variability of response in enumerative studies. Journal of the American Statistical Association, 38(222):143–152, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279534.

Palyi:1943:BRB

[Pal43b] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 38(221):126–128, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279326.

Palyi:1944:BRBa

[Pal44a] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 39(225):132–133, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280343.

Palyi:1944:BRBb

[Pal44b] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 39(226):267–268, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279971.

Paris:1940:BRB

[Par40] James D. Paris. Book review: International Monetary Organisation, by Michael A. Heilperin. Journal of the American Statistical Association, 35(210):434–435, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278740.

Paradiso:1942:BRB

[Par42] Louis J. Paradiso. Book review: [untitled]. Journal of the American Statistical Association, 37(218):294, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279233.

Paris:1943:BRB

[Par43] James D. Paris. Book review: [untitled]. Journal of the American Statistical Association, 38(223):373–374, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279796.

Parkin:1945:MSW

[Par45] Norman C. Parkin. Management statistics in war contract renegotiation. Journal of the American Statistical Association, 40 (232):504–519, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280219.

Paradiso:1948:SIC

[Par48] Louis J. Paradiso. Significance of inventories in the current economic situation. Journal of the American Statistical Association, 43(243):361–376, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280274.

Patterson:1946:UAF

[Pat46] R. E. Patterson. The use of adjusting factors in the analysis of data with disproportionate subclass numbers. Journal of the American Statistical Association, 41(235):334–346, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280257.

Pearl:1940:AP

[Pea40a] Raymond Pearl. The aging of populations. Journal of the American Statistical Association, 35(209):277–297, March 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279474.

Pearl:1940:RPO

[Pea40b] Raymond Pearl. Remarks of the presiding officer. Journal of the American Statistical Association, 35(209):217–219, March 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279463.

Peach:1947:BRB

[Pea47] Paul Peach. Book review: [untitled]. Journal of the American Statistical Association, 42(239):475–477, September 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280555.

Peach:1949:BRB

[Pea49] Paul Peach. Book review: [untitled]. Journal of the American Statistical Association, 44(248):569–570, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279910.

Peck:1942:BRB

[Pec42] Gustav Peck. Book review: [untitled]. Journal of the American Statistical Association, 37(219):394–395, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279010.

Perlman:1940:ECD

[Per40] Jacob Perlman. Extent and causes of differences in hourly earnings. Journal of the American Statistical Association, 35(209):1–12, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279650.

Perkins:1942:CAB

[Per42] John S. Perkins. A critical appraisal of business statistics. Journal of the American Statistical Association, 37(218):220–224, June 1942.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279216.

Peterson:1945:BRB

[Pet45] A. I. Peterson. Book review: [untitled]. Journal of the American Statistical Association, 40(231):395–396, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965645.

Peterson:1947:API

[Pet47] Arthur G. Peterson. Agricultural price index numbers. Journal of the American Statistical Association, 42(240):597–604, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280016.

Phelps:1940:BRB

[Phe40] Harold A. Phelps. Book review: Research and Statistical Methodology: Books and Reviews, 1933–1938, by Oscar Krisen. Journal of the American Statistical Association, 35(209):185–186, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279670.

Phelps:1941:SBP

[Phe41] Roswell F. Phelps. Sumner B. Pearmain, 1859–1941. Journal of the American Statistical Association, 36(216):545–546, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278966.

Phelps:1942:BRB

[Phe42] Harold A. Phelps. Book review: [untitled]. Journal of the American Statistical Association, 37(219):397–398, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279013.

Pierce:1943:CFM

[Pie43] Joseph A. Pierce. Correction formulas for moments of a groupeddistribution of discrete variates. Journal of the American Statistical Association, 38(221):57–62, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279312.

Pierce:1944:NSP

[Pie44a] Joseph A. Pierce. A note to “On the Summation of Progressions Useful in Time Series Analysis”. Journal of the American Statistical Association, 39(228):521, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280631.

Pierce:1944:SPU

[Pie44b] Joseph A. Pierce. On the summation of progressions useful in time series analysis. Journal of the American Statistical Association, 39 (227):387–389, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280044.

Pitman:1949:BRB

[Pit49] E. J. G. Pitman. Book review: [untitled]. Journal of the American Statistical Association, 44(247):454–455, September 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280246.

Platt:1943:MDC

[Pla43] John R. Platt. A mechanical determination of correlation coefficients and standard deviations. Journal of the American Statistical Association, 38(223):311–318, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279781.

Platt:1944:NMD

[Pla44] John R. Platt. Note to “A Mechanical Determination of Correlation Coefficients and Standard Deviations”. Journal of the American Statistical Association, 39(226):254, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279960.

Politz:1943:FVI

[Pol43] Alfred Politz. Family versus individual in measurement of audiences. Journal of the American Statistical Association, 38(222):233–237, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279544.

Pritchett:1943:BRB

[Pri43] C. Herman Pritchett. Book review: [untitled]. Journal of the American Statistical Association, 38(224):487–488, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279360.

Proudfoot:1940:BRBa

[Pro40a] Malcolm J. Proudfoot. Book review: The Structure and Growth of Residential Neighborhoods in American Cities, by Homer Hoyt. Journal of the American Statistical Association, 35(209):205–207, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279683.

Proudfoot:1940:BRBb

[Pro40b] Malcolm J. Proudfoot. Book review: [untitled]. Journal of the American Statistical Association, 35(211):587–588, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279301.

Proudfoot:1942:STT

[Pro42] Malcolm J. Proudfoot. Sampling with transverse traverse lines. Journal of the American Statistical Association, 37(218):265–270, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279223.

Proudfoot:1943:BRB

[Pro43] M. J. Proudfoot. Book review: [untitled]. Journal of the American Statistical Association, 38(221):135–136, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279334.

Politz:1949:AGH

[PS49] Alfred Politz and Willard Simmons. An attempt to get the “not at homes” into the sample without callbacks. Journal of the American Statistical Association, 44(245):9–16, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280347.

Pyle:1940:BRB

[Pyl40] J. F. Pyle. Book review: Marketing Burned-Clay Products, by A. Hamilton Chute. Journal of the American Statistical Association, 35(210):436, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278742.

Raiford:1942:SCD

[Rai42] Theodore E. Raiford. Skewness of combined distributions. Journal of the American Statistical Association, 37(219):391–393, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279009.

Rao:1947:GMA

[Rao47] C. Radhakrishna Rao. General methods of analysis for incomplete block designs. Journal of the American Statistical Association, 42 (240):541–561, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280011.

Rasor:1949:FLC

[Ras49] Eugene A. Rasor. The fitting of logistic curves by means of a nomograph. Journal of the American Statistical Association, 44 (248):548–553, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279905.

Ruggles:1947:EAE

[RB47] Richard Ruggles and Henry Brodie. An empirical approach to economic intelligence in world war II. Journal of the American Statistical Association, 42(237):72–91, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280189.

Reed:1941:BRB

[Ree41a] Harold L. Reed. Book review: [untitled]. Journal of the American Statistical Association, 36(215):459, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279645.

Reede:1941:AES

[Ree41b] Arthur H. Reede. Adequacy of employment statistics. Journal of the American Statistical Association, 36(213):71–80, March 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278815.

Reed:1942:BUD

[Ree42] Vergil D. Reed. Business uses of data by census tracts and blocks. Journal of the American Statistical Association, 37(218):238–244, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279219.

Reed:1944:BRBa

[Ree44a] Lowell J. Reed. Book review: [untitled]. Journal of the American Statistical Association, 39(228):522–523, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280632.

Reed:1944:BRBb

[Ree44b] Lowell J. Reed. Book review: [untitled]. Journal of the American Statistical Association, 39(228):523–524, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280633.

Reed:1944:WTF

[Ree44c] Vergil D. Reed. War time facts and peace time needs. Journal of the American Statistical Association, 39(226):144–154, June 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279945.

Renne:1943:BRB

[Ren43] R. R. Renne. Book review: [untitled]. Journal of the American Statistical Association, 38(222):262–264, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279559.

Reynolds:1941:BRB

[Rey41] Lloyd G. Reynolds. Book review: [untitled]. Journal of the American Statistical Association, 36(213):146–148, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278825.

Reynolds:1942:BRBa

[Rey42a] Lloyd G. Reynolds. Book review: [untitled]. Journal of the American Statistical Association, 37(217):158–159, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279450.

Reynolds:1942:BRBb

[Rey42b] Lloyd G. Reynolds. Book review: [untitled]. Journal of the American Statistical Association, 37(217):159–160, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279451.

Reynolds:1944:BRB

[Rey44] Lloyd G. Reynolds. Book review: [untitled]. Journal of the American Statistical Association, 39(226):256–257, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279962.

Rice:1945:PIF

[RHTH45] Stuart A. Rice, A. Ford Hinrichs, Howard R. Tolley, and Philip M. Hauser. Problems of integrating Federal statistics: a round table. Journal of the American Statistical Association, 40(230):237–244, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280134.

Rice:1943:QCA

[Ric43] Wm. B. Rice. Quality control applied to business administration. Journal of the American Statistical Association, 38(222):228–232, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279543.

Rice:1944:IAS

[Ric44] Stuart A. Rice. The inter American statistical institute: Four years old. Journal of the American Statistical Association, 39(226):135– 143, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279944.

Richardson:1948:VFF

[Ric48] Lewis F. Richardson. Variation of the frequency of fatal quarrels with magnitude. Journal of the American Statistical Association, 43(244):523–546, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280704.

Rice:1949:WLA

[Ric49] Stuart A. Rice. William Lane Austin (1871–1949): James Clyde Capt (1888–1949). Journal of the American Statistical Association, 44(248):565–566, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279908.

Rider:1944:BRB

[Rid44] Paul R. Rider. Book review: [untitled]. Journal of the American Statistical Association, 39(228):527, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280636.

Rider:1947:BRB

[Rid47] Paul R. Rider. Book review: [untitled]. Journal of the American Statistical Association, 42(238):331–332, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280666.

Riefler:1942:GS

[Rie42] Winfield W. Riefler. Government and the statistician. Journal of the American Statistical Association, 37(217):1–11, March 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279422.

Riggleman:1941:BRB

[Rig41] John R. Riggleman. Book review: [untitled]. Journal of the American Statistical Association, 36(214):316–317, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279768.

Riggleman:1944:BRB

[Rig44] John R. Riggleman. Book review: [untitled]. Journal of the American Statistical Association, 39(226):261–262, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279966.

Robinson:1940:BRB

[Rob40] Louis N. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 35(211):585–587, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279300.

Robinson:1941:CES

[Rob41a] Leland Rex Robinson. Corporate earnings on share and borrowed capital in percentages of gross income (1918–1940). Journal of the American Statistical Association, 36(214):253–264, June 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279746.

Robinson:1941:BRBa

[Rob41b] Louis N. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 36(214):322–324, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279774.

Robinson:1941:BRBb

[Rob41c] Louis N. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 36(215):457–458, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279644.

Roberts:1944:DPN

[Rob44] David R. Roberts. The distribution of private, non-agricultural employees in the United States by straight-time hourly wage rates. Journal of the American Statistical Association, 39(228):469–478, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280623.

Robbins:1949:BRB

[Rob49] Herbert Robbins. Book review: [untitled]. Journal of the American Statistical Association, 44(247):453, September 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280245.

Romig:1945:BRB

[Rom45] H. G. Romig. Book review: [untitled]. Journal of the American Statistical Association, 40(231):413–415, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965654.

Romig:1949:BRB

[Rom49] Harry G. Romig. Book review: [untitled]. Journal of the American Statistical Association, 44(245):142–144, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280362.

Roper:1940:SPO

[Rop40] Elmo Roper. Sampling public opinion. Journal of the American Statistical Association, 35(210):325–334, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278721.

Ross:1941:GT

[Ros41] Frank A. Ross. Gusztav Thirring, 1861–1941. Journal of the American Statistical Association, 36(216):547, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278968.

Rosander:1942:UIT

[Ros42] A. C. Rosander. The use of inversions as a test of random order. Journal of the American Statistical Association, 37(219):352–358, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279003.

Rosander:1946:BRB

[Ros46] A. C. Rosander. Book review: A Note on Statistical Books. Journal of the American Statistical Association, 41(236):621–623, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280590.

Rogers:1943:DRD

[RT43] Beatrice A. Rogers and Donald S. Tucker. Davis Rich Dewey, 1858– 1942. Journal of the American Statistical Association, 38(221):107– 109, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279321.

Roos:1948:BRB

[RW48] Charles F. Roos and H. D. Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 43(241):163–167, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280083.

Spurr:1948:SCM

[SA48] William A. Spurr and David R. Arnold. A short-cut method of fitting a logistic curve. Journal of the American Statistical Association, 43(241):127–134, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280072.

Salter:1942:CDC

[Sal42] Leonard A. Salter, Jr. A comment on Deming’s classification of problems of inference. Journal of the American Statistical Association, 37(220):540–542, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279039.

Salera:1945:BRB

[Sal45] Virgil Salera. Book review: [untitled]. Journal of the American Statistical Association, 40(229):129–130, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280467.

Samuelson:1943:BRB

[Sam43] Paul A. Samuelson. Book review: [untitled]. Journal of the American Statistical Association, 38(223):369–370, September 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279793.

Sanders:1941:BRB

[San41a] Barkev S. Sanders. Book review: [untitled]. Journal of the American Statistical Association, 36(215):439–440, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279629.

Sandomire:1941:ACP

[San41b] Marion M. Sandomire. Accumulating cubes with punch cards. Journal of the American Statistical Association, 36(216):507–514, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278958.

Sanderson:1942:NWC

[San42a] Fred H. Sanderson. The new wheat condition figures, based on weather factors, for the prairie provinces. Journal of the American Statistical Association, 37(220):473–488, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279031.

Sandomire:1942:CSC

[San42b] Marion M. Sandomire. A computational short cut for regressions based on unequal frequencies. Journal of the American Statistical Association, 37(218):282–284, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279226.

Sanderson:1944:CR

[San44a] F. H. Sanderson. Concluding remarks. Journal of the American Statistical Association, 39(225):100–101, March 1944. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280328.

Sanderson:1944:NWC

[San44b] F. H. Sanderson. The new wheat condition figures, based on weather factors, for the prairie provinces rejoinder. Journal of the American Statistical Association, 39(225):97–99, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280326.

Sasuly:1940:BRBa

[Sas40a] Max Sasuly. Book review: Initiation à l’analyse Statistique, by Leo Duge de Bernoville. Journal of the American Statistical Association, 35(209):183–184, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279667.

Sasuly:1940:BRBb

[Sas40b] Max Sasuly. Book review: [untitled]. Journal of the American Statistical Association, 35(212):711–712, December 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279263.

Satterthwaite:1941:BRBa

[Sat41a] Franklin E. Satterthwaite. Book review: [untitled]. Journal of the American Statistical Association, 36(216):575–576, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278994.

Satterthwaite:1941:BRBb

[Sat41b] Franklin E. Satterthwaite. Book review: [untitled]. Journal of the American Statistical Association, 36(216):576, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278995.

Saulnier:1940:BRB

[Sau40] R. J. Saulnier. Book review: Golden Avalanche, by Frank D. Graham; Charles R. Whittlesey. Journal of the American Statistical Association, 35(210):435–436, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278741.

Sayre:1941:CLI

[Say41] Robert A. Sayre. Cost of living indexes. Journal of the American Statistical Association, 36(214):191–200, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279738.

Scates:1941:BRB

[Sca41] Douglas E. Scates. Book review: [untitled]. Journal of the American Statistical Association, 36(216):573–574, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278992.

Scates:1942:BRB

[Sca42] Douglas E. Scates. Book review: [untitled]. Journal of the American Statistical Association, 37(217):154–155, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279447.

Schwartz:1940:BRB

[Sch40] Harry Schwartz. Book review: Tests of Significance, What They Mean and How to Use Them, by John H. Smith. Journal of the American Statistical Association, 35(209):182–183, March 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279666.

Schenker:1941:BRB

[Sch41a] David Schenker. Book review: [untitled]. Journal of the American Statistical Association, 36(213):149, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278827.

Schmalz:1941:BRB

[Sch41b] Carl N. Schmalz. Book review: [untitled]. Journal of the American Statistical Association, 36(216):553–554, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278974.

Schumacher:1942:BRB

[Sch42a] F. X. Schumacher. Book review: [untitled]. Journal of the American Statistical Association, 37(220):549–550, December 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279045.

Schwartz:1942:BRB

[Sch42b] Harry Schwartz. Book review: [untitled]. Journal of the American Statistical Association, 37(218):296–297, June 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279235.

Schneider:1943:BRB

[Sch43] John B. Schneider. Book review: [untitled]. Journal of the American Statistical Association, 38(222):257–258, June 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279555.

Schilling:1944:ADP

[Sch44a] Walter Schilling. Analysis of the data of a public health organization by the control chart method. Journal of the American Statistical Association, 39(227):311–324, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280037.

Scholz:1944:BRB

[Sch44b] Karl Scholz. Book review: [untitled]. Journal of the American Statistical Association, 39(225):131–132, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280342.

Schrock:1944:MMC

[Sch44c] Edward M. Schrock. Matters of misconception concerning the quality control chart. Journal of the American Statistical Association, 39(227):325–334, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280038.

Schneider:1945:OSW

[Sch45] David M. Schneider. Opportunities for statistical work in state and local governments. Journal of the American Statistical Association, 40(229):62–70, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280455.

Scheffe:1946:BRBb

[Sch46] Henry Scheffé. Book review: [untitled]. Journal of the American Statistical Association, 41(234):252–254, June 1946. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280498.

Scheffe:1947:CRC

[Sch47a] Henry Scheffé. [correction: “the relation of control charts to analysis of variance and chi-square tests]. Journal of the American Statistical Association, 42(240):634, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280019.

Scheffe:1947:RCC

[Sch47b] Henry Scheffé. The relation of control charts to analysis of variance and chi-square tests. Journal of the American Statistical Association, 42(239):425–431, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280545.

Schilling:1947:FDR

[Sch47c] Walter Schilling. A frequency distribution represented as the sum of two Poisson distributions. Journal of the American Statistical Association, 42(239):407–424, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280544.

Schwartz:1947:USS

[Sch47d] Harry Schwartz. On the use of Soviet statistics. Journal of the American Statistical Association, 42(239):401–406, September 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280543.

Scheffe:1948:BRB

[Sch48a] Henry Scheffé. Book review: [untitled]. Journal of the American Statistical Association, 43(242):321–323, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280377.

Schuyler:1948:OIA

[Sch48b] Garret L. Schuyler. The ordering of n items assigned to k rank categories by votes of m individuals. Journal of the American Statistical Association, 43(244):559–563, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280706.

Sekar:1949:MEB

[SD49] C. Chandra Sekar and W. Edwards Deming. On a method of estimating birth and death rates and the extent of registration. Journal of the American Statistical Association, 44(245):101–115, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280353.

Stephan:1940:SPP

[SDH40] Frederick F. Stephan, W. Edwards Deming, and Morris H. Hansen. The sampling procedure of the 1940 population census. Journal of the American Statistical Association, 35(212):615–630, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279251.

Sebba:1949:BRB

[Seb49] Gregor Sebba. Book review: [untitled]. Journal of the American Statistical Association, 44(246):332–333, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280617.

Seltzer:1940:POE

[Sel40] Lawrence H. Seltzer. The problem of our excessive banking reserves. Journal of the American Statistical Association, 35(209):24– 36, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279652.

Stock:1948:CUS

[SH48] J. Stevens Stock and Joseph R. Hochstim. Commercial uses of sampling. Journal of the American Statistical Association, 43 (244):509–522, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280703.

Shapiro:1943:DDC

[Sha43] Solomon Shapiro. The distribution of deposits and currency in the United States, 1929–1939. Journal of the American Statistical Association, 38(224):438–444, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279343.

Shaw:1944:BRB

[Sha44] Clifford R. Shaw. Book review: [untitled]. Journal of the American Statistical Association, 39(225):127–128, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280339.

Shewart:1946:ASF

[She46] Walter A. Shewart. The advancing statistical front. Journal of the American Statistical Association, 41(233):1–15, March 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280152.

Shields:1940:MPP

[Shi40] Murray Shields. A measure of purchasing power inflation and deflation. Journal of the American Statistical Association, 35 (211):461–470, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279272.

Shiskin:1942:NMS

[Shi42] Julius Shiskin. A new multiplicative seasonal index. Journal of the American Statistical Association, 37(220):507–516, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279034.

Shiskin:1945:BRB

[Shi45] Julius Shiskin. Book review: [untitled]. Journal of the American Statistical Association, 40(230):270–272, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280146.

Shoup:1943:BRB

[Sho43] Carl S. Shoup. Book review: [untitled]. Journal of the American Statistical Association, 38(223):372–373, September 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279795.

Shryock:1941:GPS

[Shr41] Henry S. Shryock, Jr. General population statistics. Journal of the American Statistical Association, 36(215):376–380, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279612.

Shryock:1943:IMW

[Shr43] Henry S. Shryock, Jr. Internal migration and the war. Journal of the American Statistical Association, 38(221):16–30, March 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279309.

Sibley:1942:BRB

[Sib42] Elbridge Sibley. Book review: [untitled]. Journal of the American Statistical Association, 37(217):148, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279441.

Siegel:1940:HEU

[Sie40] Irving H. Siegel. Hourly earnings and unit labor cost in manufacturing. Journal of the American Statistical Association, 35 (211):455–460, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279271.

Siegel:1941:DBP

[Sie41a] Irving H. Siegel. The difference between the Paasche and Laspeyres index-number formulas. Journal of the American Statistical Association, 36(215):343–350, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279608.

Siegel:1941:FND

[Sie41b] Irving H. Siegel. Further notes on the difference between indexnumber formulas. Journal of the American Statistical Association, 36(216):519–524, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278960.

Siegel:1942:IND

[Sie42a] Irving H. Siegel. Index-number differences: Geometric means. Journal of the American Statistical Association, 37(218):271–274, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279224.

Siegel:1942:MCU

[Sie42b] Irving H. Siegel. The measurement of capacity utilization. Journal of the American Statistical Association, 37(220):430–436, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279027.

Siegel:1943:NCS

[Sie43] Irving H. Siegel. Note on a common statistical inequality. Journal of the American Statistical Association, 38(222):217–222, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279541.

Siegel:1944:CPA

[Sie44] Irving H. Siegel. The concept of productive activity. Journal of the American Statistical Association, 39(226):218–228, June 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279951.

Siegel:1945:GII

[Sie45] Irving H. Siegel. The generalized “ideal” index-number formula. Journal of the American Statistical Association, 40(232):520–523, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280220.

Siegel:1946:BRB

[Sie46] Irving H. Siegel. Book review: JASA Review Policy. Journal of the American Statistical Association, 41(236):623–624, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280591.

Simon:1941:ISM

[Sim41] Leslie E. Simon. On the initiation of statistical methods for quality control in industry. Journal of the American Statistical Association, 36(213):53–60, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278813.

Simon:1942:ASM

[Sim42] Leslie E. Simon. Application of statistical methods to ordnance engineering. Journal of the American Statistical Association, 37 (219):313–324, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278999.

Simmons:1943:ELD

[Sim43] C. D. Simmons. Edward Lewis Dodd, 1875–1943. Journal of the American Statistical Association, 38(222):247–248, June 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279547.

Simon:1945:STB

[Sim45] Herbert A. Simon. Statistical tests as a basis for “yes-no” choices. Journal of the American Statistical Association, 40(229):80–84, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280458.

Snedecor:1942:RDS

[SK42] George W. Snedecor and Arnold J. King. Recent developments in sampling for agricultural statistics. Journal of the American Statistical Association, 37(217):95–102, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279435.

Shryock:1949:CSS

[SL49] Henry S. Shryock, Jr. and Norman Lawrence. The current status of state and local population estimates in the Census Bureau. Journal of the American Statistical Association, 44(246):157–173, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280594.

Smart:1940:BRB

[Sma40] L. Edwin Smart. Book review: [untitled]. Journal of the American Statistical Association, 35(211):561–562, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279286.

Smit:1940:BRB

[Smi40] Carel Jan Smit. Book review: The Influence of Prices on Exports. Journal of the American Statistical Association, 35(209):189–190, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279673.

Smith:1941:BRB

[Smi41a] Dan Throop Smith. Book review: [untitled]. Journal of the American Statistical Association, 36(214):321–322, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279773.

Smith:1941:SPU

[Smi41b] Herbert G. Smith. Some practical uses for hyperbolic grids. Journal of the American Statistical Association, 36(214):293–298, June 1941.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279750.

Smith:1942:BRB

[Smi42] John H. Smith. Book review: [untitled]. Journal of the American Statistical Association, 37(219):398–399, September 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279014.

Smith:1943:USM

[Smi43] Herbert G. Smith. Utility of statistical method in aerodynamics. Journal of the American Statistical Association, 38(223):341–345, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279785.

Smith:1945:BRB

[Smi45a] John H. Smith. Book review: [untitled]. Journal of the American Statistical Association, 40(231):380–382, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965635.

Smithies:1945:BRB

[Smi45b] Arthur Smithies. Book review: General-Equilibrium Theory in International Trade, by Jacob L. Mosak. Journal of the American Statistical Association, 40(232):533, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280226.

Smith:1946:BRB

[Smi46] John H. Smith. Book review: [untitled]. Journal of the American Statistical Association, 41(235):404–405, September 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280268.

Smith:1947:LCA

[Smi47] George C. Smith, Jr. Lorenz curve analysis of industrial decentralization. Journal of the American Statistical Association, 42 (240):591–596, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280015.

Snavely:1943:BRB

[Sna43] Tipton R. Snavely. Book review: [untitled]. Journal of the American Statistical Association, 38(224):478–479, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279352.

Snedecor:1942:UTS

[Sne42] George W. Snedecor. The use of tests of significance in an agricultural experiment station. Journal of the American Statistical Association, 37(219):383–386, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279007.

Snedecor:1947:BRB

[Sne47] George W. Snedecor. Book review: [untitled]. Journal of the American Statistical Association, 42(240):635–638, December 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280020.

Snedecor:1948:PBC

[Sne48] George W. Snedecor. A proposed basic course in statistics. Journal of the American Statistical Association, 43(241):53–60, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280066.

Snedecor:1949:UFS

[Sne49] George W. Snedecor. On a unique feature of statistics. Journal of the American Statistical Association, 44(245):1–8, March 1949. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280346.

Snider:1940:BRB

[Sni40] Joseph L. Snider. Book review: A Statistical Study of Credit Unions in New York, by Francis E. Wilcox. Journal of the American Statistical Association, 35(210):443–444, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278748.

Sogge:1944:ICU

[Sog44] Tillman M. Sogge. Industrial classes in the United States in 1940. Journal of the American Statistical Association, 39(228):516–518, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280629.

Solterer:1941:SHR

[Sol41] J. Solterer. A sequence of historical random events: Do Jesuits die in three’s? Journal of the American Statistical Association, 36(216):477–484, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278954.

Solomon:1949:BRB

[Sol49] Herbert Solomon. Book review: [untitled]. Journal of the American Statistical Association, 44(245):133–134, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280358.

Somers:1941:ADF

[Som41] Herman M. Somers. Adequacy of data in the field of public aid. Journal of the American Statistical Association, 36(213):81–90, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278816.

Spengler:1941:BRB

[Spe41] J. J. Spengler. Book review: [untitled]. Journal of the American Statistical Association, 36(214):305–306, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279756.

Spengler:1944:BRB

[Spe44] Joseph J. Spengler. Book review: [untitled]. Journal of the American Statistical Association, 39(226):255–256, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279961.

Spielmans:1948:PG

[Spi48] John V. Spielmans. Profile graphs. Journal of the American Statistical Association, 43(241):96–108, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280070.

Spurr:1940:GSC

[Spu40] William A. Spurr. A graphic short cut to the moving average method of measuring seasonality. Journal of the American Statistical Association, 35(212):667–670, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279256.

Spurr:1941:BRB

[Spu41] William A. Spurr. Book review: [untitled]. Journal of the American Statistical Association, 36(216):578–579, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278997.

Spurr:1946:BRB

[Spu46] William A. Spurr. Book review: [untitled]. Journal of the American Statistical Association, 41(236):610–612, December 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280586.

Spurr:1949:BRB

[Spu49] William A. Spurr. Book review: [untitled]. Journal of the American Statistical Association, 44(245):134–136, March 1949. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280359.

Stephan:1940:CMD

[Ste40a] Frederick F. Stephan. The Centenary Membership Directory. Journal of the American Statistical Association, 35(210):1, June 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278756.

Stephan:1940:CAS

[Ste40b] Frederick F. Stephan. The centenary of the American statistical association. Journal of the American Statistical Association, 35(209):213– 215, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279461.

Stephan:1940:NC

[Ste40c] Frederick F. Stephan. Nominating committee. Journal of the American Statistical Association, 35(211):554, September 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279281.

Stephan:1941:BRB

[Ste41] Frederick F. Stephan. Book review: [untitled]. Journal of the American Statistical Association, 36(215):437–438, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279627.

Stein:1945:BRB

[Ste45] Herbert Stein. Book review: [untitled]. Journal of the American Statistical Association, 40(231):391–393, September 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965643.

Stephan:1948:HUM

[Ste48] Frederick F. Stephan. History of the uses of modern sampling procedures. Journal of the American Statistical Association, 43(241): 12–39, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280064.

Stigler:1948:BRB

[Sti48] George J. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 43(244):603–605, December 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280713.

Stoner:1941:FEF

[Sto41] Paul Matthew Stoner. Fitting the exponential function and the Gompertz function by the method of least squares. Journal of the American Statistical Association, 36(216):515–518, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278959.

Strohkarck:1946:BRB

[Str46] Frank Strohkarck. Book review: [untitled]. Journal of the American Statistical Association, 41(234):252, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280497.

Sukhatme:1947:CPP

[Suk47a] P. V. Sukhatme. Correction: The problem of plot size in large scale yield surveys. Journal of the American Statistical Association, 42(239):460, September 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280549.

Sukhatme:1947:PPS

[Suk47b] P. V. Sukhatme. The problem of plot size in large-scale yield surveys. Journal of the American Statistical Association, 42(238):297– 310, June 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280658.

Spurr:1948:BRB

[SvS48] William A. Spurr and Victor von Szeliski. Book review: [untitled]. Journal of the American Statistical Association, 43(241):167–171, March 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280084.

Scheffe:1946:BRBa

[SW46a] Henry Scheffé and B. L. Welch. Book review: [untitled]. Journal of the American Statistical Association, 41(233):137–140, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280171.

Stewart:1946:ESP

[SW46b] Charles Stewart and Loring Wood. Employment statistics in the planning of a full-employment program. Journal of the American Statistical Association, 41(235):313–321, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280255.

Sweezy:1941:BRB

[Swe41] Maxine Yaple Sweezy. Book review: [untitled]. Journal of the American Statistical Association, 36(216):569–570, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278988.

Sweezy:1944:BRB

[Swe44] Maxine Sweezy. Book review: How Nazi Germany has Controlled Business, by L. Hamburger. Journal of the American Statistical Association, 39(227):407–408, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280059.

Szatrowski:1945:RBP

[Sza45] Zenon Szatrowski. The relationship between price change and price level for common stocks. Journal of the American Statistical Association, 40(232):467–483, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280216.

Szatrowski:1946:CGM

[Sza46] Zenon Szatrowski. Calculating the geometric mean from a large amount of data. Journal of the American Statistical Association, 41(234):218–220, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280488.

Taeuber:1944:BRB

[Tae44] Conrad Taeuber. Book review: The Displacement of Population in Europe, by Eugene M. Kulischer. Journal of the American Statistical Association, 39(227):395, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280049.

Taeuber:1945:FWS

[Tae45a] Conrad Taeuber. Farm wage statistics for a full employment program. Journal of the American Statistical Association, 40(230):234– 235, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280132.

Taeuber:1945:DSS

[Tae45b] Irene B. Taeuber. The demographic statistics of Southern and Eastern Asia. Journal of the American Statistical Association, 40(229):29–37, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280451.

Taeuber:1947:SAS

[Tae47] Conrad Taeuber. Some aspects of the statistics program in the Department of Agriculture. Journal of the American Statistical Association, 42(237):41–45, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280187.

Taft:1944:BRB

[Taf44] Philip Taft. Book review: [untitled]. Journal of the American Statistical Association, 39(228):542–543, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280647.

Taitel:1941:PIV

[Tai41] Martin Taitel. Price indexes as viewed from the stand-point of the national defense program. Journal of the American Statistical Association, 36(214):201–209, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279739.

Teper:1943:OCL

[Tep43] Lazare Teper. Observations on the cost of living index of the Bureau of Labor Statistics. Journal of the American Statistical Association, 38(223):271–284, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279777.

Tepping:1943:EDS

[THD43] Benjamin J. Tepping, William N. Hurwitz, and W. Edwards Deming. On the efficiency of deep stratification in block sampling. Journal of the American Statistical Association, 38(221):93–100, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279317.

Thorp:1943:PC

[Tho43] Willard L. Thorp. Paul T. Cherington, 1876–1943. Journal of the American Statistical Association, 38(224):471–472, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279347.

Thomas:1945:BRB

[Tho45] Dorothy S. Thomas. Book review: [untitled]. Journal of the American Statistical Association, 40(230):265–267, June 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280143.

Thom:1946:BRB

[Tho46] H. C. S. Thom. Book review: [untitled]. Journal of the American Statistical Association, 41(234):249–252, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280496.

Thorp:1947:IF

[Tho47] Willard L. Thorp. Irving Fisher, 1867–1947. Journal of the American Statistical Association, 42(238):311, June 1947. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280659.

Thorp:1948:SFP

[Tho48] Willard L. Thorp. Statistics and foreign policy. Journal of the American Statistical Association, 43(241):1–11, March 1948. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280063.

Thurstone:1948:EMM

[Thu48] L. L. Thurstone. The edge-marking method of analyzing data. Journal of the American Statistical Association, 43(243):451–462, September 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280281.

Timoshenko:1943:BRB

[Tim43] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 38(222):258–260, June 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279556.

Tintner:1940:AET

[Tin40a] Gerhard Tintner. The analysis of economic time series. Journal of the American Statistical Association, 35(209):93–100, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279659.

Tintner:1940:BRB

[Tin40b] Gerhard Tintner. Book review: [untitled]. Journal of the American Statistical Association, 35(212):716, December 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279267.

Tintner:1946:SAM

[Tin46] Gerhard Tintner. Some applications of multivariate analysis to economic data. Journal of the American Statistical Association, 41 (236):472–500, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280570.

Tintner:1947:BRB

[Tin47] Gerhard Tintner. Book review: [untitled]. Journal of the American Statistical Association, 42(238):312, June 1947. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280660.

Tintner:1948:BRB

[Tin48] Gerhard Tintner. Book review: [untitled]. Journal of the American Statistical Association, 43(243):497–499, September 1948. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280287.

Tippetts:1940:BRB

[Tip40] Charles S. Tippetts. Book review: World Economic Survey, Eighth Year, 1938–39; World Production and Prices, 1938–39. Journal of the American Statistical Association, 35(209):191–192, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279675.

Tolles:1945:LSF

[Tol45] N. Arnold Tolles. Labor statistics for a full employment program. Journal of the American Statistical Association, 40(230):235–236, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280133.

Torrens:1945:BRB

[Tor45] James G. Torrens. Book review: [untitled]. Journal of the American Statistical Association, 40(230):272–274, June 1945. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280147.

Tostlebe:1945:ESB

[Tos45] Alvin S. Tostlebe. Estimate of Series E bond purchases by farmers. Journal of the American Statistical Association, 40(231):317–329, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965626.

Tough:1940:BRB

[Tou40] Rosalind Tough. Book review: The Economics of Building, by Herbert W. Robinson. Journal of the American Statistical Association, 35(210):438–439, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278744.

Trefftzs:1941:BRBa

[Tre41a] Kenneth Lewis Trefftzs. Book review: [untitled]. Journal of the American Statistical Association, 36(213):156–157, March 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278835.

Trefftzs:1941:BRBb

[Tre41b] Kenneth Lewis Trefftzs. Book review: [untitled]. Journal of the American Statistical Association, 36(216):557–558, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278978.

Trefftzs:1941:BRBc

[Tre41c] Kenneth Lewis Trefftzs. Book review: [untitled]. Journal of the American Statistical Association, 36(216):558–559, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278979.

Truesdell:1941:NFP

[Tru41] Leon E. Truesdell. New features of the 1940 population census. Journal of the American Statistical Association, 36(215):361–368, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279610.

Truesdell:1944:BRB

[Tru44] Leon E. Truesdell. Book review: [untitled]. Journal of the American Statistical Association, 39(226):259–260, June 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279964.

Truesdell:1945:P

[Tru45] Leon E. Truesdell. Population. Journal of the American Statistical Association, 40(230):232–233, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280131.

Tolley:1944:WDA

[TT44] H. R. Tolley and Conrad Taeuber. Wartime developments in agricultural statistics. Journal of the American Statistical Association, 39(228):411–427, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280619.

Tucker:1942:DI

[Tuc42] Rufus S. Tucker. Distribution of income in 1935-36. Journal of the American Statistical Association, 37(220):489–495, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279032.

Tucker:1943:IRS

[Tuc43a] Donald S. Tucker. The interest rate and saving. Journal of the American Statistical Association, 38(221):101–102, March 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279318.

Tucker:1943:CIO

[Tuc43b] Rufus S. Tucker. The composition of income and ownership of capital by income classes in the United States in 1936. Journal of the American Statistical Association, 38(222):187–200, June 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279538.

Tucker:1945:IBS

[Tuc45] Rufus S. Tucker. Industry and business statistics: Discussion. Journal of the American Statistical Association, 40(230):231, June 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280130.

Tukey:1947:RBT

[Tuk47a] John W. Tukey. Review of A Teoria da Indução Estadistica, by J. Kingston. Journal of the American Statistical Association, 42(237):190–191, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280205.

Tukey:1947:RBF

[Tuk47b] John W. Tukey. Review of The Factorial Analysis of Human Ability, Second Edition, by Godfrey H. Thomson. Journal of the American Statistical Association, 42(237):199–200, March 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280209.

Tukey:1949:RBT

[Tuk49a] John W. Tukey. Review of Probability Theory for Statistical Methods, by F. N. David. Journal of the American Statistical Association, 44(248):567–568, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279909.

Tukey:1949:RBE

[Tuk49b] John W. Tukey. Review of Theory of Experimental Inference, by C. W. Churchman. Journal of the American Statistical Association, 44(245):136–139, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280360.

Thorndike:1943:CLC

[TW43] E. L. Thorndike and Ella Woodyard. The cost of living in cities in relation to their size and latitude. Journal of the American Statistical Association, 38(222):238–241, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279545.

Tukey:1946:BRB

[TW46a] John W. Tukey and Charles P. Winsor. Book review: Industrial Statistics. Journal of the American Statistical Association, 41 (235):406–411, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280270.

Tukey:1946:RBE

[TW46b] John W. Tukey and J. Wolfowitz. Review: Industrial Experimentation, by K. A. Brownlee. Journal of the American Statistical Association, 41(233):125–128, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280164.

Ullman:1945:ARC

[Ull45] Morris B. Ullman. Apportionment of representatives in congress on the basis of special populations. Journal of the American Statistical Association, 40(232):484–492, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280217.

Ulmer:1946:ETC

[Ulm46] Melville J. Ulmer. On the economic theory of cost of living index numbers. Journal of the American Statistical Association, 41 (236):530–542, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280574.

Vaile:1940:BRB

[Vai40] Roland S. Vaile. Book review: Trends in Size and Production of the Aggregate Farm Enterprise, 1909–36, by Raymond G. Bressler, Jr.; John A. Hopkins; Trends in Employment in Agriculture, 1909–36, by Eldon E. Shaw; John A. Hopkins. Journal of the American Statistical Association, 35(209):194–196, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279677.

Vaile:1941:BRB

[Vai41] Roland S. Vaile. Book review: [untitled]. Journal of the American Statistical Association, 36(214):320–321, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279772.

Veenstra:1940:BRB

[Vee40] Theodore A. Veenstra. Book review: [untitled]. Journal of the American Statistical Association, 35(211):584–585, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279299.

vonHuhn:1942:MSD

[vH42] R. von Huhn. Mechanization of statistical drafting. Journal of the American Statistical Association, 37(219):387–390, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279008.

Vickrey:1941:EFR

[Vic41] William Vickrey. The effects of Federal Revenue Acts of 1938, 1939, and 1940 on the realization of gains and losses on securities: Comment. Journal of the American Statistical Association, 36(215):431–433, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279623.

Villard:1941:BRB

[Vil41] Henry H. Villard. Book review: [untitled]. Journal of the American Statistical Association, 36(215):456–457, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279643.

Viner:1940:GAS

[Vin40] Jacob Viner. Greetings of the Allied Social Science Associations. Journal of the American Statistical Association, 35(209):273–274, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279472.

Vormelker:1941:HLW

[Vor41] Rose L. Vormelker. How the librarian would like to find statistics. Journal of the American Statistical Association, 36(214):225–232, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279742.

vonSzeliski:1940:CA

[vS40] Victor von Szeliski. Chapter activities. Journal of the American Statistical Association, 35(209):138–146, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279663.

Wadley:1945:APS

[Wad45] F. M. Wadley. An application of the Poisson series to some problems of enumerations. Journal of the American Statistical Association, 40(229):85–92, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280459.

Waite:1942:PFA

[Wai42] Warren C. Waite. Price fixing of agricultural products. Journal of the American Statistical Association, 37(217):13–21, March 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279424.

Waite:1944:IQI

[Wai44] Warren C. Waite. An interpretation of the quantity index. Journal of the American Statistical Association, 39(228):507–510, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280627.

Waite:1947:BRB

[Wai47] Warren C. Waite. Book review: [untitled]. Journal of the American Statistical Association, 42(239):486–487, September 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280560.

Walker:1940:BRB

[Wal40] Helen M. Walker. Book review: Mathematical Tables, Volume VII, The Probability Integral, by W. F. Sheppard. Journal of the American Statistical Association, 35(210):428–429, June 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278735.

Walker:1941:CEF

[Wal41] Q. Forrest Walker. A correction: The effects of Federal Revenue Acts of 1938, 1939, and 1940 on the realization of gains and losses on securities. Journal of the American Statistical Association, 36(213): 122, March 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278822.

Wallich:1942:BRB

[Wal42] Henry C. Wallich. Book review: [untitled]. Journal of the American Statistical Association, 37(218):301–302, June 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279240.

Wald:1943:BRB

[Wal43a] Abraham Wald. Book review: [untitled]. Journal of the American Statistical Association, 38(224):491–492, December 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279364.

Walker:1943:BRB

[Wal43b] Dilworth Walker. Book review: [untitled]. Journal of the American Statistical Association, 38(221):138–139, March 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279336.

Walker:1944:BRB

[Wal44] Helen M. Walker. Book review: [untitled]. Journal of the American Statistical Association, 39(225):125–127, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280338.

Wald:1945:SMS

[Wal45a] Abraham Wald. Sequential method of sampling for deciding between two courses of action. Journal of the American Statistical Association, 40(231):277–306, September 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965624.

Walker:1945:BRB

[Wal45b] Helen M. Walker. Book review: [untitled]. Journal of the American Statistical Association, 40(231):399–403, September 1945. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965648.

Walker:1945:RAS

[Wal45c] Helen M. Walker. The role of the American Statistical Association. Journal of the American Statistical Association, 40(229):1–10, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280448.

Wallace:1945:USF

[Wal45d] Henry A. Wallace. The use of statistics in the formulation of a national full employment policy. Journal of the American Statistical Association, 40(229):11–19, March 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280449.

Wald:1947:BRB

[Wal47a] A. Wald. Book review: [untitled]. Journal of the American Statistical Association, 42(237):185–187, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280203.

Walker:1947:BRB

[Wal47b] Helen M. Walker. Book review: [untitled]. Journal of the American Statistical Association, 42(238):342–344, June 1947. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280670.

Wallis:1949:SKR

[Wal49a] W. Allen Wallis. Statistics of the Kinsey Report. Journal of the American Statistical Association, 44(248):463–484, December 1949.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279900.

Walsh:1949:ASS

[Wal49b] John E. Walsh. Applications of some significance tests for the median which are valid under very general conditions. Journal of the American Statistical Association, 44(247):342–355, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280233.

Walsh:1949:BCS

[Wal49c] John E. Walsh. On the best choice of sample sizes for a t-test when the ratio of variances is known. Journal of the American Statistical Association, 44(248):554–558, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279906.

Walsh:1949:ILU

[Wal49d] John E. Walsh. On the “information” lost by using a t-test when the population variance is known. Journal of the American Statistical Association, 44(245):122–125, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280355.

Warne:1943:BRB

[War43] Colston Warne. Book review: [untitled]. Journal of the American Statistical Association, 38(223):379–381, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279801.

Wardwell:1945:BRB

[War45] Charles A. R. Wardwell. Book review: Theory of Economic Aggregates: A Manual of General Business Forecasting, by Montgomery D. Anderson. Journal of the American Statistical Association, 40 (232):524–526, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280221.

Warburton:1948:BRB

[War48] Clark Warburton. Bank reserves and business fluctuations. Journal of the American Statistical Association, 43(244):547–558, December 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280705.

Watts:1940:BRB

[Wat40] V. O. Watts. Book review: Philadelphia Labor Market Studies. Journal of the American Statistical Association, 35(209):192–193, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279676.

Watson:1941:BRBb

[Wat41a] Alfred N. Watson. Book review: [untitled]. Journal of the American Statistical Association, 36(215):454–455, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279641.

Watson:1941:BRBa

[Wat41b] J. P. Watson. Book review: [untitled]. Journal of the American Statistical Association, 36(215):450, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279637.

Watkins:1949:SRE

[Wat49] Ralph J. Watkins. Statistical requirements for economic mobilization. Journal of the American Statistical Association, 44(247):406–412, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280239.

Waugh:1943:CDV

[Wau43] Frederick V. Waugh. Choice of the dependent variable in regression analysis. Journal of the American Statistical Association, 38(222):210–216, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279540.

Waugh:1946:CPC

[Wau46] Frederick V. Waugh. The computation of partial correlation coefficients. Journal of the American Statistical Association, 41 (236):543–546, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280575.

Walker:1940:EFR

[WCH40] Q. Forrest Walker, W. T. Collins, and D. E. Higgins. The effects of Federal Revenue Acts of 1938, 1939, and 1940 on the realization of gains and losses on securities. Journal of the American Statistical Association, 35(212):602–614, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279250.

Walker:1941:EFR

[WCH41] Q. Forrest Walker, W. T. Collins, and D. E. Higgins. The effects of Federal Revenue Acts of 1938, 1939, and 1940 on the realization of gains and losses on securities: Rejoinder. Journal of the American Statistical Association, 36(215):433–435, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279624.

Weida:1944:HLR

[Wei44] Frank Mark Weida. Henry Lewis Rietz, 1875–1943. Journal of the American Statistical Association, 39(226):249–251, June 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279957.

Welsh:1941:BRB

[Wel41] Charles A. Welsh. Book review: [untitled]. Journal of the American Statistical Association, 36(216):561–562, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278981.

Welch:1944:BRB

[Wel44] Emmett H. Welch. Book review: Earnings and Social Security in the United States, by W. S. Woytinsky. Journal of the American Statistical Association, 39(227):398–399, September 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280052.

Wernette:1940:BRB

[Wer40] J. Philip Wernette. Book review: Turning Points in Business Cycles, by Leonard P. Ayres. Journal of the American Statistical Association, 35(210):424–426, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278732.

West:1942:BRB

[Wes42] Paul V. West. Book review: [untitled]. Journal of the American Statistical Association, 37(218):295–296, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279234.

Whelpton:1946:RRA

[Whe46] P. K. Whelpton. Reproduction rates adjusted for age, parity, fecundity, and marriage. Journal of the American Statistical Association, 41(236):501–516, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280571.

Whelpton:1947:BRB

[Whe47] P. K. Whelpton. Book review: [untitled]. Journal of the American Statistical Association, 42(240):643–644, December 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280024.

White:1943:SIM

[Whi43] R. Clyde White. The social insurance movement. Journal of the American Statistical Association, 38(223):358–364, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279788.

Wilks:1940:BRB

[Wil40a] S. S. Wilks. Book review: Methods of Statistical Analysis, by C. H. Goulden. Journal of the American Statistical Association, 35(209):181–182, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279665.

Willcox:1940:LSS

[Wil40b] Walter F. Willcox. Lemuel shattuck, statist, founder of the American statistical association. Journal of the American Statistical Association, 35(209):224–235, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279466.

Williams:1940:FCM

[Wil40c] Faith M. Williams. Factors to be considered in measuring intercity and interregional differences in living costs. Journal of the American Statistical Association, 35(211):471–482, September 1940. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279273.

Wilks:1941:BRB

[Wil41a] S. S. Wilks. Book review: [untitled]. Journal of the American Statistical Association, 36(213):148–149, March 1941. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278826.

Willcox:1941:JCS

[Wil41b] Walter F. Willcox. Josiah Charles Stamp, 1880–1941. Journal of the American Statistical Association, 36(216):546–547, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278967.

Williams:1941:BRB

[Wil41c] Alfred H. Williams. Book review: [untitled]. Journal of the American Statistical Association, 36(213):151–152, March 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278830.

Wilson:1941:BRB

[Wil41d] Edwin B. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 36(215):436–437, September 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279626.

Wilcox:1942:BRB

[Wil42a] Sidney W. Wilcox. Book review: [untitled]. Journal of the American Statistical Association, 37(219):401–402, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279016.

Wilcoxen:1942:MFS

[Wil42b] L. C. Wilcoxen. The market forecasting significance of market movements. Journal of the American Statistical Association, 37 (219):343–351, September 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279002.

Williams:1943:BRB

[Wil43] Ernest W. Williams, Jr. Book review: [untitled]. Journal of the American Statistical Association, 38(224):485, December 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279358.

Williams:1944:BRB

[Wil44a] Ernest W. Williams, Jr. Book review: Railway Traffic Expansion and Use of Resources in World War II, by Thor Hultgren; Railway Wage-Rates, Employment and Pay, by John L. McDougall. Journal of the American Statistical Association, 39(227):401–402, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280054.

Wilson:1944:NWC

[Wil44b] C. F. Wilson. The new wheat condition figures, based on weather factors, for the prairie provinces: a further reply. Journal of the American Statistical Association, 39(225):99–100, March 1944. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280327.

Willcox:1946:FH

[Wil46a] Walter F. Willcox. Frederick L. Hoffman, 1865–1946. Journal of the American Statistical Association, 41(234):240–241, June 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280492.

Wilm:1946:DAM

[Wil46b] H. G. Wilm. The design and analysis of methods for sampling microclimatic factors. Journal of the American Statistical Association, 41(234):221–232, June 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280489.

Willcox:1949:OCA

[Wil49] Walter F. Willcox. Obituary: Conrad Alexander Verrijn Stuart (1865–1948). Journal of the American Statistical Association, 44(246):295–296, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280602.

Winslow:1944:BRB

[Win44] Harry J. Winslow. Book review: Unemployment Compensation Experiences of Beneficiaries in Columbus, Ohio: 1939–1940, by Rose L. Papier. Journal of the American Statistical Association, 39 (227):396–397, September 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280050.

Wishart:1946:BRB

[Wis46] John Wishart. Book review: [untitled]. Journal of the American Statistical Association, 41(234):259–262, June 1946. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280501.

Wallis:1941:STT

[WM41] W. Allen Wallis and Geoffrey H. Moore. A significance test for time series analysis. Journal of the American Statistical Association, 36(215):401–409, September 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279616.

Webb:1943:PAT

[WNP43] John N. Webb, M. Starr Northrop, and Stanley L. Payne. Practical applications of theoretical sampling methods. Journal of the American Statistical Association, 38(221):69–77, March 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279314.

Wolfe:1940:BRB

[Wol40] A. B. Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 35(211):588–589, September 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279302.

Wolff:1943:BRB

[Wol43] Georg Wolff. Book review: [untitled]. Journal of the American Statistical Association, 38(223):381–382, September 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279802.

Wolff:1944:BRB

[Wol44] George Wolff. Book review: [untitled]. Journal of the American Statistical Association, 39(226):257–259, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279963.

Wolfowitz:1946:BRB

[Wol46] J. Wolfowitz. Book review: [untitled]. Journal of the American Statistical Association, 41(235):405–406, September 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280269.

Wolfowitz:1949:BRB

[Wol49] J. Wolfowitz. Book review: [untitled]. Journal of the American Statistical Association, 44(246):330–331, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280615.

Woodbury:1941:MFL

[Woo41a] Robert M. Woodbury. Methods of family living studies, a reply to the review. Journal of the American Statistical Association, 36(214):299– 300, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279751.

Woodbury:1941:RR

[Woo41b] Robert M. Woodbury. Reply to the rejoinder. Journal of the American Statistical Association, 36(214):302, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279753.

Woods:1942:BRB

[Woo42a] Baldwin M. Woods. Book review: [untitled]. Journal of the American Statistical Association, 37(217):165–166, March 1942. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279455.

Woofter:1942:PPE

[Woo42b] T. J. Woofter, Jr. Preliminary population estimates based on ration book applications. Journal of the American Statistical Association, 37(220):437–440, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279028.

Woodbury:1944:ECS

[Woo44a] Robert Morse Woodbury. Economic consumption scales and their uses. Journal of the American Statistical Association, 39(228): 455–468, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280622.

Woofter:1944:MAF

[Woo44b] T. J. Woofter, Jr. A method of analysis of family composition and income. Journal of the American Statistical Association, 39 (228):488–496, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280625.

Woofter:1949:RNR

[Woo49] T. J. Woofter. The relation of the net reproduction rate to other fertility measures. Journal of the American Statistical Association, 44(248):501–517, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279902.

Working:1940:BRB

[Wor40a] E. J. Working. Book review: On Agricultural Policy, 1926–1938, by Joseph S. Davis. Journal of the American Statistical Association, 35(209):197–201, March 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279679.

Working:1940:WCP

[Wor40b] Holbrook Working. War and commodity prices. Journal of the American Statistical Association, 35(210):309–324, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278720.

Worthing:1941:BRB

[Wor41] Marion W. Worthing. Book review: [untitled]. Journal of the American Statistical Association, 36(216):566–568, December 1941. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278986.

Working:1943:SLF

[Wor43] Holbrook Working. Statistical laws of family expenditure. Journal of the American Statistical Association, 38(221):43–56, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279311.

Worthing:1944:RWE

[Wor44a] A. G. Worthing. A reply to W. Edwards Deming’s review of treatment of experimental data. Journal of the American Statistical Association, 39(226):252–254, June 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279959.

Worthing:1944:BRB

[Wor44b] Marion Worthing. Book review: [untitled]. Journal of the American Statistical Association, 39(228):533, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280641.

Working:1945:SQC

[Wor45] Holbrook Working. Statistical quality control in war production. Journal of the American Statistical Association, 40(232):425–447, December 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280214.

Working:1946:BRB

[Wor46a] Holbrook Working. Book review: [untitled]. Journal of the American Statistical Association, 41(233):112–115, March 1946. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280161.

Working:1946:NSP

[Wor46b] Holbrook Working. Note on sampling probabilities. Journal of the American Statistical Association, 41(234):238–239, June 1946. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280491.

Woytinsky:1941:BRB

[Woy41] W. S. Woytinsky. Book review: [untitled]. Journal of the American Statistical Association, 36(214):308–309, June 1941. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279759.

Williams:1942:CLI

[WRS42] Faith M. Williams, Frances R. Rice, and Emil D. Schell. Cost of living indexes in wartime. Journal of the American Statistical Association, 37(220):415–424, December 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279025.

Wickens:1943:NMT

[WW43] Aryness Joy Wickens and Faith M. Williams. Notes on Mr. Teper’s observations. Journal of the American Statistical Association, 38(223):284–286, September 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279778.

White:1944:SIM

[WW44] Eric Clyde White and R. Clyde White. The social insurance movement. Journal of the American Statistical Association, 39 (228):519–521, December 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280630.

Wynne:1944:BRBa

[Wyn44a] William H. Wynne. Book review: [untitled]. Journal of the American Statistical Association, 39(225):129–130, March 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280341.

Wynne:1944:BRBb

[Wyn44b] William H. Wynne. Book review: [untitled]. Journal of the American Statistical Association, 39(228):532, December 1944. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280640.

Yentis:1943:BRB

[Yen43] David Yentis. Book review: [untitled]. Journal of the American Statistical Association, 38(221):130–131, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279329.

Yerushalmy:1948:BRB

[Yer48] J. Yerushalmy. Book review: [untitled]. Journal of the American Statistical Association, 43(243):492–493, September 1948. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2280284.

Yntema:1940:BRBb

[Ynt40a] Dwight B. Yntema. Book review: Studies in Income and Wealth, Volume Three, by Conference on Research in National Income and Wealth. Journal of the American Statistical Association, 35 (210):432–433, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278738.

Yntema:1940:CSF

[Ynt40b] Dwight B. Yntema. Changing seasonal fluctuations in the amounts of public and private assistance and earnings on CWA and WPA projects in 116 urban areas, 1929-38. Journal of the American Statistical Association, 35(212):644–652, December 1940. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279253.

Yntema:1940:BRBa

[Ynt40c] Theodore O. Yntema. Book review: Mathematical Analysis for Economists, by R. G. D. Allen. Journal of the American Statistical Association, 35(210):427–428, June 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278734.

Yntema:1942:BRB

[Ynt42] Dwight B. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 37(219):402–403, September 1942. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279017.

Yoder:1942:BRB

[Yod42] Dale Yoder. Book review: [untitled]. Journal of the American Statistical Association, 37(217):160–161, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279452.

Yoder:1943:BRB

[Yod43] Dale Yoder. Book review: [untitled]. Journal of the American Statistical Association, 38(221):131, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279330.

Yoder:1944:BRB

[Yod44] Dale Yoder. Book review: [untitled]. Journal of the American Statistical Association, 39(225):128–129, March 1944. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280340.

Young:1943:BRB

[You43] Charles E. Young. Book review: [untitled]. Journal of the American Statistical Association, 38(222):261–262, June 1943. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279558.

Young:1946:APP

[You46] Charles E. Young. Applications and problems of productivity data. Journal of the American Statistical Association, 41(236):421–431, December 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280566.

Zeisel:1947:CMR

[Zei47] Hans Zeisel. Coordinating the measurements of radio listening. Journal of the American Statistical Association, 42(240):512–522, December 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280008.

Zimmerman:1940:BRB

[Zim40] Carle C. Zimmerman. Book review: [untitled]. Journal of the American Statistical Association, 35(212):709, December 1940. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279261.

Zimmerman:1941:R

[Zim41] Carle C. Zimmerman. Rejoinder. Journal of the American Statistical Association, 36(214):300–302, June 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279752.

Zimmermann:1943:BRB

[Zim43] Erich W. Zimmermann. Book review: [untitled]. Journal of the American Statistical Association, 38(222):254–256, June 1943. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279553.

Zubin:1949:BRB

[Zub49] Joseph Zubin. Book review: [untitled]. Journal of the American Statistical Association, 44(246):315, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280607.